A Bibliography of Publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer

(1980–1989)

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

28 October 2022 Version 3.08

Title word cross-reference

.COM [Wil83e].

100 [McK84, Wil85a]. 16-Bit [Dun83c, Dun87f, Cam86d, Dun82a, Dun82b, Dun82c, Dun82d, Dun83a, Dun83b, Dun83d, Dun83e, Dun83f, Dun83g, Dun83h, Dun83i, Dun83j, Dun83k, Dun83l, Dun84a, Dun84b, Dun84c, Dun84d, Dun84e, Dun84f, Dun84g, Dun84h, Dun84i, Dun84j, Dun85a, Dun85b, Dun85c, Dun85d, Dun85e, Dun85f, Dun85g, Dun85h, Dun85i, Dun85j, Dun86a, Dun86b, Dun86c, Dun86d, Dun86e, Dun86f, Dun86g, Dun86h, Dun86i, Dun86j, Dun87a, Dun87b, Dun87c, Dun87d, Dun87e, Dun87g, Gar84]. 16BST [Bar85, Bur83, Cat85, Coh85, Dun84l, Dun84m, Fer84, Gar84, Gol84, Hea85a, Hea85b, Kap84, Kaw84, Mit85b, Mor84, Pri84, Sea82, Smi85, Spe83, Wil83e, Wil83f].

2 [And85, And89a, Car85b, Cor87a, Dun88, Dun89a, Dun84m, Lad89c, MS84, MP89, Por88b, Por88p, Sha86a, Sha86f, Sha87h, Vos88, WA86]. 2.2 [MM84b]. 20 [Fis84]. 232 [Hug82]. 2x [Wat83].

3 [May87]. 3-D [May87]. 32000 [Rod86].

5th [Gri82].

64K [Wil83e]. 68000s [Kos82c]. 680xx [Tur87b, Tur87a]. 68K [Don85, Rea86].

8-bit [Bla82, Cam86d]. 80 [Ano83, Ash85a, Cla85, Cro84, Lou84, Ouv84, Sca83, Car82]. 8080/8086 [Mil82]. 8086/88 [Bur83]. 8088/8086 [Bla82]. 8088/8087 [Dun84l]. 83 [Dun84k, Sha83]. 86 [Ash85a, Cor82a]. ’89 [Fis89b]. 8in [Ham83].

9 [Cap87]. 9000 [Ano83].

A* [Nev89]. A/UX [Mea88a]. Abstraction [WA86]. Accent [Vas84]. Accurate [RB84]. Ackermann [New84]. acquisition [Bun89]. Act [Sti84, Flo89a]. action [Sti88]. ACTxx [New83]. Ada [Con86, Jon86, Kin88a, Mit83b, Mit83c, Mit83d, Mit83e, Ste88i]. Adapter [Bar86]. Add [Cha82a, Kan82a]. Adding [Ams88a, Car84, Carxxb, Day85]. Additions [Van82]. ADFGVX [Bur85]. Adler [Bow83b]. Advanced [Mar89]. Adventures [Ams89]. Adventurous [Ams88b]. Advice [Ham86b]. Advisor [Sch85a]. Age [Gen85, Har83c, Sut85e]. Aid [Tay82, Nut88]. AIM [FG82]. al [Mar87a]. Algorithm [Bro87b, Col83, Get89b, McI86, Mic82, Ruz84, Men89b, Nev89, Pet89a, Rus89]. Algorithms [Fre83, Gor82b, Gor83b, Ste87]. Alice [Pal85]. allocation [Moo88b, Pet89b]. allocators [And89c]. Alternative [How87, Vos88]. Amiga [Har87a]. AMPRO [Con85]. Analog [Mus87]. Analog-To-Digital [Mus87]. Analysis [Pir83]. Analytic [Lun83]. Analyzer [Lin88]. Analyzing [Sch85a]. Anatomy [Shi83]. Angell [Bow83a]. annealing [McL89]. ANSI [Lad89d, Rel87b]. API [Par88]. APL [Bur89]. APL*PLUS [Ung85]. APL*PLUS/PC [Ung85]. apologia [VW89]. Apple [Mea88a]. Apples [Pul83]. Appletalk [BL87]. application [Sch89b]. Applications [Bom84, Fre83, Genxx, Hol87d, Rel87c, Tur86j, Dun88, Fit89]. Applying [Leb88]. Approach [Hua84, RM88]. Approximation [Gor82b]. AppWare [Firxx]. ARC [Nel87]. Architecture [GPxx]. Archiving [Ash85a, Nel87]. Argc [Joh82]. Argum [Skj82a]. Arguments [Moo88a]. Argv [Joh82]. Arithmetic [Smi83a]. array [Ton88b]. Arrays [Tic88, And89d, Por88c]. Art [Mar87a, Por88b, Ste88c]. Artificial [Bro87a, Tel87j, Tel87a, Tel87b, Tel87c, Tel87d, Tel87e, Tel87f, Tel87g, Tel87h, Tel87i, Tel88d]. ASCII [Bla82, Gar84, Pas83a]. Assemble [Cam86c]. Assembler [And86a, And86b, And86c, Dun82e, Kos82a, Kos82b, Mil82, Per83, Rod86, Dil88]. Assemblers [New83, Sch89c]. Assembly [Dis87, Fox85, Gan83, Gin82, Rea86, Bla82, Dun84l, Moo88a, Til85]. Assembly-Language [Dis87]. Async [BL87]. Asynchronous [Car85a]. Attaching [Gol83]. Augusta [Mey83, Mit83a, Mit83b, Mit83c, Mit83d, Mit83e, Mit83f, Mit83g]. Automata [CM89]. Automated [Mil87]. Automatic [Sor86, Win89, Moo88a]. Automation [Hal84]. Automaton [Par86]. Autorouting [Nev89]. Avoiding [Ros89]. AWK [Mis89b]. AWK-like [Mis89b].

B [BD83, DB83, TT86]. B-Tree [BD83, DB83]. Background [Nic82]. Backtracking [Bow87a]. Backups [Gas89, Tur89]. Baggage [Ton89b]. Balancing [Sti84]. Banked [Woo85]. BANKSWAP [Woo85]. Barstow [Pro84]. Based [TT88, Kin89a]. BASIC [Car82, Beh83, Bro84b, Hum84, Irv83, Pir83, Rol83, Ros85, Sea82, Sey82, GT82, Mor83, Sha87h, Ton88b]. BASICFMT [Cro84]. BASICS [Ton88a]. Batch [Don85, Wil85b]. BBS [Har87b]. BDOS [Bol83, Jon83]. Becomes [Rea86]. Becoming [Cus89f].

Being [Cus89f]. Benchmark [Gol84, Kaw84, Mor84, Pri84, Spe83, Dun84l, Dun84m, Fer84, VW89]. Benchmarking [Fox89, Lad89a, RHV86, KP89]. Benchmarks [Col83, Gla83b]. Berry [Ord82]. Berthold [Gor82a]. best [MP89]. Better [Col83, Rea83a, Rea83b, TT86]. BetterBASIC [Tra85]. Between [Sor86, Joh89]. Beyond [Wal85]. Bibliography [Rog84]. BIGLOAD [Wil83e]. Binary [Bla82, Fre83, Gar84, Mat88, Pas83a, Wil84]. BIOS [Bol83, Gri84b, Smi84]. Bit [And85, Dun83c, Dun87f, Gay83, Mor86, Mus87, Wal85, Bla82, Cam86d, Dun82a, Dun82b, Dun82c, Dun82d, Dun83a, Dun83b, Dun83d, Dun83e, Dun83f, Dun83g, Dun83h, Dun83i, Dun83j, Dun83k, Dun83l, Dun84a, Dun84b, Dun84c, Dun84d, Dun84e, Dun84f, Dun84g, Dun84h, Dun84i, Dun84j, Dun85a, Dun85b, Dun85c, Dun85d, Dun85e, Dun85f, Dun85g, Dun85h, Dun85i, Dun85j, Dun86a, Dun86b, Dun86c, Dun86d, Dun86e, Dun86f, Dun86g, Dun86h, Dun86i, Dun86j, Dun87a, Dun87b, Dun87c, Dun87d, Dun87e, Dun87g, Gar84]. Bit-Mapped [Mor86]. Bit/Sec [Wal85]. Bjarne [Ste89o]. Board [Con85]. Boards [EK86, Ouv82]. Bob [Lac83, Lac83, Rod85]. Boca [Bro86]. Bolter [Sut85e]. Bookshelf [Con85]. Boot [Ros89]. Bootle [Rob82]. Bose [Cel85]. Bose-Nelson [Cel85]. Box [Wil83a]. Boyer [Men89b]. BRIE [Bro86]. Brief [Hym88b]. Bringing [Wil86]. Brothers [Ano88b, Ano88c, Ano88d]. Browser [Joh84]. Bryant [DeG84]. Bug [Hol87d, Rus85, Ame85, Mac83]. bugs [CV89]. Building [Bla88, Sch89a, Sim83]. Builds [Con82]. Bulletin [Ouv82]. Bursts [Bro85]. Bus [Firxx, Gol83, Wil85a]. Buttons [Har87a].

C [And89c, And89d, Ano85, Bar84b, Ber89a, Bol83, Car86a, Car89, Cha88b, Coh85, Dei87, Dlu89a, Dlu89b, Dlu89c, Dov84, Eck89, Edg89, Flo89c, Fou83, Fox89, Fre88a, Gay85, GR89a, GR89b, Gre89b, Hal82, Har86, Hen82a, HP84, Hen85, Hol84, Hol85a, Hol85b, Hol85c, Hol85d, Hol85e, Hol85f, Hol85g, Hol86a, Hol86b, Hol86c, Hol86d, Hol86e, Hol86f, Hol86g, Hol86h, Hol86i, Hol86j, Hol86k, Hol86m, Hol87e, Hol87f, Hol87g, Hol87h, Hol87i, Hol87j, Hol87k, Hol87l, Hol87m, Hol87n, Hol87o, Hol87p, Hol87a, Hol87d, Hol88a, Hol88b, Hol88c, Hol88d, Hol88e, Hol88f, How83, Joh82, Kle87, Kra85, Kva82, Lad89a, Lad89c, Lad89b, Lad89d, Lad89f, Mac83, McD84, McK83, Mer89, Mil87, Mis89b, Nut88, Pio85, Poo89, Por87a, Por87b, Rea83a, Rea83b, Rea86, RHV86, Rel87a, Rel87b, Rus89, Sch89a, Sch84a, Sch84b, Skj82a]. C [Skj83b, Skj83a, Skj83c, Skj84f, Skj84h, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84g, Sta83, Sta84, Sta89, Ste88b, Ste88c, Ste88d, Ste88e, Ste88f, Ste88g, Ste88h, Ste89a, Ste89b, Ste89c, Ste89d, Ste89e, Ste89f, Ste89g, Ste89h, Ste89i, Ste89j, Ste89k, Ste89l, Ste89m, Ste89n, Tay82, Tel88b, Tel88c, TB89, Tic88, Van82, War84, Wat89a, Wei89, Wil83b]. C-64 [Pio85]. C-INDEX [Fre88a]. C-to-Fortran [Flo89c]. C/UNIX [Skj83a, Skj83c, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84g, Skj83b, Skj84f, Skj84h]. Cache [Dei87]. Calculations [Cla84a]. Calisthenics [Wie84a, Wie84b, Wie84c]. Calls [Bol83]. Caltech [DeG84]. Can [Bli87, Jon83]. Card [Ano83, And89b]. Case [Mel89, KR88, Leb88]. Catalog [Ano88b, Ano88c, Ano88d]. cc [Sch84a]. CD [Ano88b]. CD-ROM [Ano88b]. Cellular [CM89, Par86]. Censorship [Gil82]. Centaur [Wig84e]. Chain [Cor87b, Har87c, Mit85b]. Challenges [Sha87h]. Character [Cla85]. Characters [Pal85]. charts [Sti88]. Check [Ing82]. Checkbook [Sti84]. checker [Bom82c]. CheckMate [Dav88]. Checksums [Blu82a, Blu82b]. Chest [Hol85a, Hol85b, Hol85c, Hol85d, Hol85e, Hol85f, Hol85g, Hol86a, Hol86b, Hol86c, Hol86d, Hol86e, Hol86f, Hol86g, Hol86h, Hol86i, Hol86j, Hol86k, Hol87e, Hol87f, Hol87g, Hol87h, Hol87i, Hol87j, Hol87k, Hol87l, Hol87m, Hol87n, Hol87o, Hol87p, Hol88a, Hol88b, Hol88c, Hol88d, Hol88e, Hol88f]. Chinese [Hua84]. Chiu [Wat83]. Christensen [Kra85]. Cinnabar [Dav88]. Cipher [Bur85]. Circle [Bli87, Lee83, McW83a]. Circles [Pul83]. Circuits [Mey83, Ste83]. class [Flo89a, Tel88b]. class-ier [Tel88b]. classes [Lad89b]. Classification [RW82]. Clinic [Cor82b, Cor82c, Cor82d, Cor82e, Cor82f, Cor82g, Cor82h, Cor82i, Cor82j, Cor82k, Cor83a, Cor83b, Cor83c, Cor83d, Cor83e, Cor83f, Cor83g, Cor83h, Cor83i, Cor83j, Cor83k, Cor84c, Cor84d, Cor84b, Cor84e, Cor84f, Cor84g, Cor84h, Cor84i, Cor84j, Cor84k, Cor84l, Cor85b, Cor85c, Cor85d, Cor85e, Cor85f, Cor85g, Cor85h, Cor85i, Cor85j, Cor85k]. CLINK [Ban84]. Clock [Kan82a, Wil85a]. Clothes [Gut89]. CMOS [Ste83]. Code [Bro85, Buv84, Fox85, Gor83a, Jam82, Nar87, Pir83, Ton89b, Tur86a, Joh89, Mit83a, Whi89]. Coding [Wis83]. Coffron [Ord84]. COHERENT [Gom84]. Collisions [Ros89]. Color [Der89, Mur83a, Sea82]. Column [Tra87a, Tra87b, Tra88b, Tra88c, Tra88d, Tra88e, Tra88f]. COM [Cat86b, Kap84, Zim87]. COM.ASM [Kap84]. Command [Day85, Ram83, Skj82a]. Comments [Doh83, Doh84, Mey83]. Commodore [Pio84, Ryl85]. Common [Irv83]. Common-Sense [Irv83]. CommonView [Ber89b]. Communications [Bro84a, Dun89a, Ser89]. COMMWHIZ [PCC82]. Compact [Ano88c]. compaction [Pet89b]. Comparator [Cor84m]. compared [Rel88b, Sch89c]. Comparing [Lad89c]. Comparison [Cor82a, Ste87]. Compatible [Hol87q]. compilation [Duv88]. Compiler [Buv84, Edg85, Feyx3, Fou83, Fra82a, Fra82b, Hen82a, Mac83, Mit83g, RMB85, Van82, Rel88b]. Compilers [Ano85, Dov84, Mit83f, RCS85, RHV86, Rel87a, Sha86a, Por88b]. Complex [Cla84a, Siy89]. Component [KWxx]. Compressing [Whi87]. Compression [Tur86a, Duv88, Nel89]. CompuServe [TT86]. Computer [Bow83a, Dun82f, Fle84, Lyo83, Ouv82, Roh84, PCC82, Sut85e, Wie84a, Wie84b, Wie84c, Ser89, Ups88, Lac83]. Computerized [Cro82]. Computers [Con85, Doh83, Doh84, Erd87, Gri82, Gri83, Gri84a, Har83c, Kan82b, Smi83a, Tur87a]. Computin [Lac83]. Computing [Har83c, Ode84, Mar87a]. CompuView [Elk89]. Concepts [BD83, DB83]. Concordance [Sta84]. Concurrency [Ber86a]. Concurrent [GR89a, GR89b]. Conditional [Don85, Ram83]. Conference [DeG84, Jam82]. Confessions [Lyo83]. Connection [CG84, Sho82b, Flo89c, Ord84]. connections [Tra88a]. Console [Mit85a]. Constant [Gra87]. Consulting [Gom84, Hin84, Cla84b]. Container [Hej89]. Continued [Mit83g]. Contour [May87]. Control [Ash83, Ode83, Sha86g, Wat89a, Hog85a, Ste88c, Win89]. Controller [Gay85]. Conversion [Gar84, Mus87, Tay82, Wil83d, Bla82]. Conversions [Pas83a]. Converting [Dun84k]. Coping [Siy89]. coprocessors [McM89]. COPY [Day85]. Copyright [Fre86]. CORBA [Betxx]. Coroutines [Mor83]. Corp [Kin85]. correct [Mey89]. Cosine [Ruz83a]. CP [Ano83, Ash85a, Ben82, Blu83a, Blu83b, Blu83c, Blu83d, Blu83e, Blu83f, Blu83g, Blu83h, Blu83i, Blu83j, Blu84a, Blu84b, Blu84c, Blu84d, Blu84e, Blu84f, Blu84g, Blu84h, Blu84i, Blu84j, Blu85a, Blu85b, Blu85c, Blu85d, Blu85e, Blu86, Bol83, Bom82b, Cor82a, Cor84m, Dae82, Don85, Gri84b, Ham83, Hea82a, Hea82b, Hea82c, Hea82d, Hea83a, Hea83b, Hea83c, Hea83d, Joh82, MM84a, MM84b, Mur83b, Pio84, Pio85, Rus85, Sil84, Smi84, Wei83, Wha83, Wil83d, Wil85b, Wil85c, Wil85d, Woo85]. CP/M [Ano83, Ash85a, Ben82, Blu83a, Blu83b, Blu83c, Blu83d, Blu83e, Blu83f, Blu83g, Blu83h, Blu83i, Blu83j, Blu84a, Blu84b, Blu84c, Blu84d, Blu84e, Blu84f, Blu84g, Blu84h, Blu84i, Blu84j, Blu85a, Blu85b, Blu85c, Blu85d, Blu85e, Blu86, Bol83, Bom82b, Cor82a, Cor84m, Dae82, Don85, Gri84b, Ham83, Hea82a, Hea82b, Hea82c, Hea82d, Hea83a, Hea83b, Hea83c, Hea83d, Joh82, MM84a, MM84b, Mur83b, Pio84, Pio85, Rus85, Sil84, Smi84, Wei83, Wha83, Wil83d, Wil85b, Wil85c, Wil85d, Woo85]. CP/M-68K [Don85]. CP/M-80 [Ano83, Ash85a]. CP/M-86 [Ash85a, Cor82a]. CPME [Blu83a, Blu83b, Gri84b, Mil84, Mit85a, Ram83, Rus85, Wil85b, Wil85d]. CRC [Rit86]. Creating [Ams88b, Pot89b, Pot89a]. Creative [Clu85]. Crescent [Ton89c]. Cross [And86a, And86b, And86c, Buv84, Kos82a, Kos82b, Mil82, New83, Rod86, Tay82, Hel88]. Cross-Reference [Tay82]. Crowded [Spe88a]. Crunchers [Wat83]. Crunchers-21 [Wat83]. Cryptographer [Sca86]. Cryptography [Bur84a, Bur84b, Mic88]. Cs [Rel88b]. Cubic [Ash86]. Culture [Sut85e]. CUPN [Gro84]. Curses [Hol87q]. Cursor [Ash83, Hog85a]. Curve [Ash86, Shi83]. customized [And89c].

D [May87]. D. [Tur89]. Data [Jon87, Kar89, Tho86, WA86, Whi87, Bun89, Nel89]. Data-flow [Kar89]. Database [ADL88]. Databases [Ste88a]. Datamost [Cri82]. Dates [Kin83]. David [Sut85e]. dBASE [Hea84a, Hea84b, Hea84c, Hea85c]. DDC [Beh83, Dav85, Ham83, Joh82, Pir83, Ros85, Smi83a, Til85]. DDJ [Ral89]. DDT [Mit85a]. Debugger [CV89]. Debugging [Bla83, Edg89, Gro88b, Men89a, Mit85a, Woo85]. Dec82 [Doh83]. Decision [Hog85b]. Decoded [Car82]. Decoding [Whi89]. Decompiler [May88]. Defense [Hol86m, Wha83]. deficiencies [Fer84]. Definition [Tin82]. Demand [Dah89]. demo [Rol83]. Dennis [Ams89, Rod85, Ste89o]. Descent [Mit83f]. Descriptive [Wil83b]. Design [DDJ86, Edl86, Feyx3, Ham87f, KR88, RMB85, Tak84, Win85]. Designer [Swa84i, Swa84j, Swa85n, Swa85o, Swa85p, Swa85q]. Designing [Gar87, Hec85, Wil85a]. DESQview [Str89]. Destroyed [Mur83b]. Developer [Cus89f, Par88]. Developing [Car88, Rel87c]. Development [Gea88, Sti88]. Device [Bas84, Kle87, Mit85b, Wat89b, Mis89a]. Device-driver [Mit85b]. Devil [Rob82]. Diagnostic [Ame83, Jen85]. Dialects [Sor86]. Diary [Ste88i]. Dictionary [Rob82]. Diff [Cor84m, Kra87]. Digital [Ame83, Hol87d, Mor86, Mus87, Shi83, Ste83]. Digits [Ing82]. Dimensional [Jon87, Gro88a]. DIR [Por88a]. Direct [Cha82a]. Directories [Mur83b]. Directory [Dae84, Mon83, Dlu89a, Ham83]. Disassembler [Wil83c]. Disc [Ano88c]. Discrete [GR89b]. Disk [Blu83a, Blu83b, Gol83, Rus85, Ste85, Wil83d]. DISKDOC [Kva82]. Diskette [Ame83, Dae84]. Disks [Tur89]. Display [Mit85b]. Display/Dump [Mit85b]. Dissolve [Mor86]. Distributed [Firxx, Genxx]. Divisibility [Gor83b]. Division [Mau82a, Smi83b]. DML [Har83d]. Do [Jon88, Kap84]. DO-COM.ASM [Kap84]. Dobb [Cor82b, Cor82c, Cor82d, Cor82e, Cor82f, Cor82g, Cor82h, Cor82i, Cor82j, Cor82k, Cor83a, Cor83b, Cor83c, Cor83d, Cor83e, Cor83f, Cor83g, Cor83h, Cor83i, Cor83j, Cor83k, Cor84c, Cor84d, Cor84b, Cor84e, Cor84f, Cor84g, Cor84h, Cor84i, Cor84j, Cor84k, Cor84l, Cor85b, Cor85c, Cor85d, Cor85e, Cor85f, Cor85g, Cor85h, Cor85i, Cor85j, Cor85k, Har83a, AS85a, Cor85a, Pit85, Rod85]. Doctor [AS85a, Pit85, Rod85]. documenting [Lad88, Nut88]. DOD [Con86]. DOL [Tho86]. Don’t [Mar87b]. Double [Hua84, Hel88]. Double-Headed [Hua84]. Down [Ton88a]. DP [Rob82]. Dr [Ams89]. Dr. [Cor82b, Cor82c, Cor82d, Cor82e, Cor82f, Cor82g, Cor82h, Cor82i, Cor82j, Cor82k, Cor83a, Cor83b, Cor83c, Cor83d, Cor83e, Cor83f, Cor83g, Cor83h, Cor83i, Cor83j, Cor83k, Cor84c, Cor84d, Cor84b, Cor84e, Cor84f, Cor84g, Cor84h, Cor84i, Cor84j, Cor84k, Cor84l, Cor85b, Cor85c, Cor85d, Cor85e, Cor85f, Cor85g, Cor85h, Cor85i, Cor85j, Cor85k, Cor85a]. Draw [Bli87]. Drawing [Ryl85, Smi85, And89b]. dream [Joh89]. Driver [McC83, Sch84a, Zim87, Mis89a, Mit85b]. Drivers [Bas84, Kle87, Wat89b]. Drives [Ano88b]. Dumb [Ash83]. Dump [Mit85b]. Dwo [Cri83]. Dynamic [Cor87a, JS89, Kin88b, McM87, Mer89, Ton89a]. Dysan [Ame83].

early [Swa89a]. Ecco [Ams89]. Echelon [Sim86]. Ed [Hal84, DeG84]. Editor [Bla83, Cla85, How83, McW83b, Rea82, Rea83a, Rea83b, Woo82, Cat86a, WH89]. Editorial [Eri88a, Eri88b, Eri88c, Eri88d, Eri88e, Eri88f, Eri88g, Eri88h, Eri89a, Eri89b, Eri89c, Eri89d, Eri89e, Eri89f, Eri89g, Eri89h, Eri89i, Eri89j, Eri89k, Eri89l, Flo89b, Har83b, Lad89e, Ouvxx, Rasxx, Rob88, Rud88, Spe87a, Spe87b, Spe87c, Spe88b, Spe88c, Spe88d, Sut88, Swa84a, Swa84b, Swa84c, Swa84d, Swa84e, Swa84f, Swa84g, Swa84h, Swa85a, Swa85b, Swa85c, Swa85d, Swa85e, Swa85f, Swa85g, Swa85h, Swa85i, Swa85j, Swa85k, Swa85l, Swa86a, Swa86b, Swa86c, Swa86d, Swa86e, Swa86f, Swa86g, Swa86h, Swa86i, Swa87a, Swa87b, Swa87c, Swa87d, Swa87e, Swa87f, Swa87g, Swa87h, Swa87i, Swa88b, Tur86b, Tur86c, Tur86d, Wig83a, Wig83b, Wig83c, Wig83d, Wig83e, Wig83f, Wig83g, Wig83h, Wig83i, Wig84a, Wig84b, Wig84c, Wig84d, Swa88a]. Editors [Edw85, Nic82, TT87]. Education [Hol87r]. Efficient [Bro87b]. EGA [Bre87, Cha88a, Cha88b]. Eight [EK86]. elements [Ton88b]. Ellipse [Enr83]. Embedded [KR88]. Emperor [Gut89]. Emulate [Fis84]. encoding [Dal89, Zig89]. Encryption [TT84a, TT84b, Tho86]. Engine [Bro86]. Engineering [Boy88, Rob88, War85]. Enhanced [Bar86, Bur85]. Enhancing [How83]. Envelope [Lev87]. environment [Bau89, Laz88]. Environments [Boy88, Swa85r, Joh89, Pro84]. Epson [Sho82b, Cla85, Gay83, Pal85]. erious [Wei89]. Errors [Rus85]. Essay [Feyx3, Gri82]. Ethics [Hol86l]. Eugene [Ste83]. Evaluation [Kan82b]. Event [IB89, GR89b]. every [Jon88]. Ex [Har83a]. Excalibur [Kin85]. Exchange [Blu83c, Blu83d, Blu83e, Blu83f, Blu83g, Blu83h, Blu83i, Blu83j, Blu84a, Blu84b, Blu84c, Blu84d, Blu84e, Blu84f, Blu84g, Blu84h, Blu84i, Blu84j, Blu85a, Blu85b, Blu85c, Blu85d, Blu85e, Blu86, Hea82a, Hea82b, Hea82c, Hea83a, Hea83b, Hea83c, Hea83d, Sch85c, Sch85d, Wil85c]. Excuse [Bom82a]. Exec [Hol87d, Kok89]. Executable [Laz89]. execute [Kap84]. Execution [Wei86]. Expand [Skj82b]. Expansion [Ano83, Gat83, Ham83]. Experiment [Bro87a]. Experimental [Tur87b]. Expert [Gri87b, Gri87a, Par84, Par86, Ano88a]. Expert-2 [Par86]. Expression [Hol84]. Expressions [Hol87b]. Extended [Dlu89a, Sha86g, Zim87, Tho89]. Extensible [Fra82a, Hel89]. Extension [DT84]. Extensions [Sil84, Ams88a, Mis89b].

Facility [Sta83]. Factoring [Ham86a]. Factors [War85]. Fantasies [AS85b, AS85c]. Fantasy [AS85a]. Farvour [Car82]. Fast [Cha87, Gor83b, Ham86e, Lee83, McW83a, Ruz83b, Ruz83c, Ruz83d, Ruz83e, Spe88a, Ull84, Gas89, Tur89, Bar84a]. Faster [Irv83, Men89b, Pul83, Ton89b]. Fatten [LR85]. Feb86 [Sch85b]. Festschrift [AS85a, Pit85, Rod85]. Few [Fit88]. Fgrep [Ash85b]. Fifth [Doh83, Gri83]. Fig [Dun84k, Mon83]. Fig-Forth [Dun84k, Mon83]. File [Cad84, Coh85, Cor84m, Hug82, Joh84, Mon83, Mur83b, Ste87, Wil83d, Ham83, Men89c, Wil83e, Lun83]. Files [Ash85a, BMO88, Mes84, TT86, Qui89]. Filter [Hea85a, Hea85b, Mot84]. filtering [Kli89]. Filters [Fra89, Hol87d]. Find [Hym88a, Hym88b]. Finder [Vas84, Hym88a]. Finite [Bow87b, Smi89]. First [Ber89b, Hua84, McI86, Wil86]. First-Order [McI86]. fit [MP89]. Fitting [Ash86]. Five [Skj82b, Rel88b]. Fix [Mar86a, Rus85, Mac83]. Flames [Eri88i, Eri88j, Swa86j, Swa86k, Swa86l, Swa86m, Swa86n, Swa86o, Swa86p, Swa87j, Swa87k, Swa87l, Swa87m, Swa87n, Swa87o, Swa87p, Swa87q, Swa87r, Swa87s, Swa87t, Swa87u, Swa88m, Swa88n, Swa88o, Swa88p, Swa88q, Swa88r, Swa88s, Swa88t, Swa88u, Swa88v, Swa89n, Swa89o, Swa89p, Swa89q, Swa89r, Swa89s, Swa89t, Swa89u, Swa89v, Swa89w, Swa89x, Swa89y]. Flexible [Wil83a]. FLIP [Gar83a]. Float [Mar87b]. Floating [DT84, Gol84, Kaw84, Mon82, Mor84, Pri84, Smi83a, Spe83, Dun84l, Dun84m, Fer84]. Floating-Point [DT84, Mor84, Dun84l, Dun84m, Fer84]. Floppy [Wil83d]. flow [Kar89]. Foehn [Hin84]. Fok [Cri83]. Font [Cha88b]. Force [Kin89a]. Force-based [Kin89a]. Formal [Tin82]. FORTH [Dun82e, Tin82, Bar82, Bar83, Bar84a, Ber83, Ber84b, Bla83, Bro85, Buv84, Cad84, Car86b, Cha87, Col83, Cor84a, Dil88, Dun84k, Eme82, Fle87, Gat83, Gla83a, Ham86a, Ham86b, Ham87b, Ham87a, Hua84, IB89, Jam82, Lin86, Man83, McC83, Mon82, Mon83, Mor83, Ode82, Ode83, Per83, RMB85, Ruz83b, Ruz83c, Ruz83d, Ruz83e, Sha83, Sha86g, Sol82, Sol83, Tay83, Tin83, Tra87a, Tra87b, Tra87c, Tra88b, Tra88c, Tra88d, Tra88e, Tra88f, Tra88a, Wil85e, Wis83, Woo82]. Forth-83 [Dun84k, Sha83]. Forth-Like [Ber83, Ber84b]. Fortran [Flo89c, Kaw84]. Four [Pie87, Por88b]. Fourier [Bar84a]. FPCA [Fis89b]. Frequency [Sim83]. Front [Sho82a]. Full [Bla83]. Full-Screen [Bla83]. Function [Bur83, New84, Smi84, Bar84b, Hym88a, Hym88b]. Functional [Fis89b]. Functions [Car84, Carxxb, Coh85, Hol87q, Wei83]. Fundamental [Feyx3]. Future [Gre83]. FVG [DT84]. FX [Cla85]. FX-80 [Cla85]. fx80char [Cla85].

G [Gor82c]. Gadgets [Har87a]. game [Wig84e, Gar83b]. Games [Wat83]. Gaspar [Tur89]. GEM [Fit89, Swa85r]. General [McI86]. Generalized [Hol84]. GENERATE.BAS [Hum84]. Generating [Lan89]. Generation [Doh83, Doh84, Gri82, Gri83, Gri84a, Gri86, Mic82]. Generator [Rol83, Shi83, Sta84, Tay82]. generic [Rus89]. Gestalt [RM88]. Getting [CV89, Gas88, Ton88a]. Glitches [Mar86a]. global [Mis89a]. GNU [Sta85]. GO [Tin83]. Goes [Car86b, MM84a, MM84b]. Going [Lad89d, Tur87a]. Graham [Hal84]. graphical [Dah88]. Graphics [Bad83, Bar86, Bre87, Cal87, Dae83, Dun82f, Gay83, Hog85b, McN86, Mor86, Mur83a, Por87a, Por87b, Por89a, Por89b, Por89c, Por89d, Por89e, Por89f, Por89g, Sea82, McM89, Ups88, Bow83a]. Graphs [Kie89]. Great [Rit86]. Greek [FT84, TF84]. Grep [Hol84]. Guest [Feyx3, Gri82, Lad89e, Rasxx]. Guide [Irv83, Wal84, War84, Ste89b]. Guidebook [Ste83]. Guides [Por88d].

H [Sho82a, Woo82]. H-19 [Woo82]. H-8 [Sho82a]. Hack [Pit85]. Hackers [AS85a]. Hacks [Ack87]. Hamming [Whi89]. Hamming-Code [Whi89]. Handling [BMO88, Mil87]. Hard [Gol83, Tur89]. Hartley [Ull84]. Hashing [Flo87, Hol87b, Hel89]. Hayes [HG82]. Head [Gas88]. Headed [Hua84]. Heapsort [Rus89]. Help [Sta83]. Henry [Gor82a, Wat83]. Hex [Gar84, Pas83a]. hexadecimal [Bla82]. High [And87, Cla84b, EK86, Flo87, Men89c, Sum87, Wei83]. High-Level [And87, Sum87, Wei83]. High-Performance [Flo87]. High-Speed [EK86, Men89c]. High-Tech [Cla84b]. Highfalutin [Lac83]. History [Mes84, New84]. Hnatek [Ste83]. Homebrew [Dlu89b]. Horn [Gor82a]. HP [Bar85]. huge [Por88c]. Human [DDJ86, Edl86, War85]. Hyper [And89b]. HyperAnimation [Gas88]. HyperCard [All88a, Kru88d]. Hypertext [Gre83].

I/O [Blu83a, Blu83b, Car84, Carxxb, Dun88]. Ian [Bow83a]. IBM [Ord84, Camxx, Dae83, Dae84, Dun82f, EK86, Fun82, FM85, Gar83a, Gar83b, Hum84, Lin86, Mur83a, PP85, Pir83, Sea82, Sha86a, Smi85, Zim87]. IBM-PC [Gar83b]. ICE [Rel88a]. icon [WH89]. Ideas [Gro84]. ier [Tel88b]. II [Bur89, Bur84b, Hea84a, Hea84b, Hea84c, Mil82, Pot89a]. III [Mit83f, Ouv84, Ruz83e]. Image [BMO88, Duv88, Duv89, Whi87]. Images [May87]. ImageWriter [Mayxx]. Implementation [Gin82, Hol87a, Tay83]. Implementations [FM85, Swa89a]. Implemented [Ber83]. Implementing [Get89b, Mil84, Por88o, Ser89, Valxxa]. Impossibilities [Erd87]. Improved [GT82, Ban84, Wil85b]. Improvement [Mau82a]. Improving [Por88a]. Inc. [Hor85, Sol85, PCC82]. Income [Sch85a]. independent [Ton89a]. Index [Ral89, Fre88a]. Inefficient [Har86]. Inference [Bro86]. Infinite [TT84a, TT84b]. Information [Gen85, Nic82]. INIT [Ros89]. initialization [Mar88]. INP [Mil84]. Input [Ham87b]. Inserting [Ton88b]. Instructions [Tan83]. Instrument [Bom84]. Integer [Cat85, Cat86a, Gra87, Ham86e, Smi83b, Ton88b]. Integers [Mar87b]. Integrated [Ste83]. Integration [Broxx, DeG84]. Intelligence [Tel87j, Tel87a, Tel87b, Tel87c, Tel87d, Tel87e, Tel87f, Tel87g, Tel87h, Tel87i, Tel88d]. Interactive [Fra82a, Pro84]. Interest [Col86a, Col86b, Col86c, Col86d, Col86e, Col86f, Cor82l, Cor82m, Cus89a, Cus89b, Cus89c, Cus89d, Cus89e, Rag85a, Rag85b, Rag85c, Sut84a, Sut84b, Sut84c, Sut85a, Sut85b, Sut85c, Sut85d, Swa85m, Wie82a, Wie82b, Wie82c, Wie82d, Wie82e, Wie82f, Wie82g, Wie83a, Wie83b, Wie83c, Wie83d, Wie83e, Wie83f, Wie83g, Wie83h, Wie83i, Wie83j, Wie83k, Wie83l]. Interface [Cor84a, DDJ86, Edl86, Wil83f]. interfaces [Dah88]. Interfacing [FG82]. Internals [Rog84]. Interoperable [Valxxb, Valxxa]. Interpreted [Tur87b]. Interpreter [Mit83a, Eck89, Sch89a]. Interpreters [Tur86a, AI89]. Interpretive [Gor82c]. Interprocess [Dun89a]. Interrupt [Mil87]. Interrupts [Bom82b, Bom84]. Interview [DDJ86, Ste89o]. Introducing [Valxxb]. Introduction [All88a, Bow83a, Dov84, MS84, Sey84, Wil85e]. Introductory [Wal84]. ISAM [BD83, DB83]. Issues [Gen85, McN86, Sch85a, War85]. Italic [Cha88b]. IV [Mit83g]. IZE [Ano88d].

J. [Sut85e]. Jacque [Lyo83]. James [Car82, Ord84]. Jef [DDJ86]. John [Car83, Cla84b]. Jorg [Hal84]. JRT [Hun83]. Julian [Kin83].

Keep [App86, Cor84b]. Kelly [Rob82]. Kelly-Bootle [Rob82]. Kermit [And89a]. Kernel [Ber86b, Pas83b, Gre89b]. Key [Bur84a, Bur84b, TT84a, TT84b, Dav85, Mic88]. key-rollover [Dav85]. Kit [Mea88b]. Klaus [Gor82a]. Knowledge [But87a, TT88]. Knowledge-Based [TT88]. Kriya [Hor85].

Lance [Gan83]. Lane [Spe88a]. Language [Ams88b, Ber83, Bro84b, Car83, Dis87, Dov84, Fox85, Fra89, Gan83, Ord82, Rea86, Ton89a, Tur87b, Bla82, Dun84l, Hum84, Moo88a, Tel88c, Til85]. Language-independent [Ton89a]. Languages [And87, Ber84a, Ber84b, Cor84b, Sum87, TT88, Gor82c]. Large [Bro87b, CFH88, DeG84]. Large-Scale [CFH88]. larger [Wil83e]. LaserJet [Bar85]. Learning [Jon86]. length [Dal89, Zig89]. Less [Ton89b]. Let [PP85, GT82]. Letter [Cat86a, Ame85, Bar84b, Bla83, Col83, Eve84, Fou83, Gla83b, Gro84, Jen85, Mac83, Mey83, New84, Ruz83a, Smi83a, Ull84, Van82, Wha83, Wil83b]. Level [And87, Clu85, Sum87, Wei83, Fis89a]. Leventhal [Gan83, Mau82a]. Liberating [AS85b, AS85c]. Libraries [JS89]. Library [HP84]. Life [Gar83b]. Lifetree [Fun82]. Light [Cop88a, Cop88b, Hol87s, Spe87d, Spe87e, Spe87f, Spe88e, Spe88f, Spe88g, Swa88k, Swa88l, Tho87, Tur86e, Tur86f, Tur86g, Tur86h, Tur86i, Tur87c, Tur87d, Tur87e, Tur87f]. like [Mis89b, Ber83, Gri87b, Gri87a, Hol84, Ber84b]. Line [Mar86a, MP89, Rol83, Skj82a, Smi85, Jon88]. Line-of-best-fit [MP89]. LineDrawing [Por88o]. Link [JS89, Hug82]. Linkage [Mor83]. linked [Ste89a]. Linking [Cor87a, Sch89d, Wei83, Hum84]. links [Ser89]. Lipinski [Bow83b]. Lisp [Tel87j, Gol84, Ams88a, Tel87k, Gor82a]. list [Sta89]. Listing [Dae84]. lists [Ste89a]. Little [Con85]. LLL [GT82]. Load [Wil83e, Kap84]. Loader [Bar82]. Local [Sha86f]. Loeliger [Gor82c]. Logic [But87a, But87b, Mal85, Aye89]. Logitech [Sch85b]. Logo [Pri84]. Lok [Cri83]. Look [Hun83, Ber89b]. Low [Hol87d, Mur83a, Sea82]. Low-Pass [Hol87d]. Low-Res [Mur83a, Sea82]. LRU [Get89b]. Ltd [Wei89]. LZW [Nel89].

M [Ben82, Blu83a, Blu83b, Blu83c, Blu83d, Blu83e, Blu83f, Blu83g, Blu83h, Blu83i, Blu83j, Blu84a, Blu84b, Blu84c, Blu84d, Blu84e, Blu84f, Blu84g, Blu84h, Blu84i, Blu84j, Blu85a, Blu85b, Blu85c, Blu85d, Blu85e, Blu86, Bol83, Bom82b, Cor84m, Dae82, Gri84b, Ham83, Hea82a, Hea82b, Hea82c, Hea82d, Hea83a, Hea83b, Hea83c, Hea83d, Joh82, MM84a, MM84b, Mur83b, Pio84, Pio85, Rus85, Sil84, Smi84, Wei83, Wha83, Wil83d, Wil85b, Wil85c, Wil85d, Woo85]. M-68K [Don85]. M-80 [Ano83, Ash85a]. M-86 [Ash85a, Cor82a]. Mac [AS85b, AS85c, Bas85a, Bas85b, Bas85c, LR85, Mea88b, Pit85]. MacApp [Bia89]. MacFORTH [Clu85]. Machina [Har83a]. machine [Hum84, Tel87j]. Machines [Bow87b, Smi89]. Macintosh [ADL88, All88b, Coh85, Cus89f, Har87a, Kat86, Pie87, Wat89a, Wat89b]. Macro [Gin82]. Macro-Assembly [Gin82]. Macros [Bur83]. Macs [Kru88a, Kru88b, Kru88c]. Magic [Fre83, Pal85, Wil84]. Mainframe [CG84]. Maintaining [Moi89]. Maintenance [Cad84, Kva82]. Make [Ing82, Str89]. Making [Flo89c]. Man [Sut85e]. Management [Bia89, Lun83, Mar89, Sey82]. Manager [Dei87, Sta89, MP89, Sch89b]. Mandelbrot [Kat86]. Manifesto [Sta85]. Manipulation [And85]. Many [Bli87]. Mapped [Mor86]. Mapping [Moo88b].

Maps [May87]. MASM [Hel88]. Matching [Ash85b, Bow87b, RM88]. Math [TF84, FT84, Hin84]. mathematics [Duv89]. Matrix [Ruz83b, Ruz83c, Ruz83d, Ruz83e]. Matters [TT87]. May83 [Ban84]. May84 [Doh84]. MBASIC [Sha84]. MBOOT [Pio85]. MC68000 [Buv84, Cat86b, Pas83b]. McWORDER [McW83b]. Me [Bom82a]. Measurement [Cro82]. Meets [Mayxx, And89a]. Mega [Rel88a]. Memory [Bia89, McM87, Sey82, Ste86, Woo85, And89c, Dah89, Kok89, Mar89, Mer89, Moo88b, Pet89b, Tho89]. Merge [Ash84]. Message [Hil88]. Message-Passing [Hil88]. Metaphor [Hec85]. Metric [Mar86b]. Micro [CG84, Jon86]. Microcomputer [Wal84]. Microcomputers [Car85b, Kin83]. Microprocessor [Bro85]. Micros [Mit83b, Mit83c, Mit83d, Mit83e]. Microsoft [Car82, Hol87d, JS89, PP85]. microSUB [Hin84]. Minds [Gri86]. Mini [Fle87, Tur86a, AI89]. Mini-Interpreters [Tur86a, AI89]. minilanguages [AI89]. Minimax [Ruz84]. Minimizer [Mar86b]. Mistake [Feyx3]. Mnemonics [Mau82b]. mode [Gre89a, Mar88]. Model [Camxx, KWxx]. Modeling [Sof86, Bow89]. Models [Ouv84]. Modem [Blu85f, Jam82, McK84]. MODEM7 [Pio85]. Modems [Wal85]. Modifications [Van82]. Modifying [Tur89]. mods [Ban84]. Modula [And85, And89a, Car85b, Dun84m, Lad89c, MS84, Por88b, Por88p, Sch85b, Sha86a, Sha86f, Sha87h, WA86]. Modula-2 [And85, And89a, Car85b, Dun84m, Lad89c, MS84, Por88b, Por88p, Sha86a, Sha86f, Sha87h, WA86]. Modula-Feb86 [Sch85b]. Module [Por88o, Win89]. Modules [Sha86f, Sol83]. Monitor [Bom84]. Monty [War85]. Moore [Men89b]. Most [Ing82]. Motorola [Bar83]. Mouse [PP85, Wil83f]. Move [Cor84b]. MS [Bow89, Cor82a, Hea85a, Hea85b, Kle87, Lad88, Mis89a, Mit85b, Nel87, RCS85, Sch89c, Wei86]. MS-DOS [Bow89, Cor82a, Hea85a, Hea85b, Kle87, Lad88, Mis89a, Mit85b, Nel87, RCS85, Sch89c]. MSDOS [Ano85]. Mulisp [Car84, Carxxb]. Mullish [Wat83]. Multi [Kos82c]. Multi-68000s [Kos82c]. multidimensional [And89d]. MultiFinder [Der88]. multiple [Ser89]. Multiplications [Gra87]. Multipliers [Gra87]. Multitasking [Ber86b, Lin87, McM89, Tur86j, Gre89b, Gre89a, Kar89]. Mushroom [Pal85]. Music [Gar87]. MYCIN [Gri87a, Gri87b]. MYCIN-Like [Gri87a, Gri87b]. Mysteries [Car82]. Mystery [Rit86].

Names [Ham87a]. Naming [Ham87a]. Native [Buv84]. Native-Code [Buv84]. NBASIC [Sha84]. NCI [Gom84]. Nelson [Cel85]. Neon [Hor85]. net [Swa89a]. NetBIOS [Men89c]. nets [Kli89]. Network [Bro87a, Col84, Get89a, IB89, Kie89, Lyo83, Gom84]. Networks [Mel89, Kin89b]. Neural [Bro87a, Kli89, Mel89, Kin89b, Swa89a]. NeXT [Genxx]. noise [Kli89]. Nondeterministic [Ode83]. North [Sho82b]. North* [Hen82b]. Norton [Por88d]. Notebook [Skj83b, Skj83a, Skj83c, Skj84f, Skj84h, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84g]. Notes [Win85, Swaxx]. Novell [Firxx]. Novice [Con82]. NR [Hol87a]. Nroff [Hol87a, Hol87b]. NS32000 [Cam86a]. Nu [Rel88a]. Nu-Mega [Rel88a]. Numbers [Cla84a, Fre83, Wil84]. Numerals [Hol87b]. Numeric [Beh83, Ham87b, Smi83a]. Numerical [Gor82b, Mar87a]. Nyan [Bow83b].

O [Bow83a, Blu83a, Blu83b, Car84, Carxxb, Dun88]. Object [Camxx, Dye89, Fra89, Gro88a, Gut89, Kie89, KWxx, Ste88a, Tel87k, Tel87l, Tel87m, Wat89a, Aye89, Dun89b, Flo89a, Hej89, Tel88c]. Object-Oriented [Dye89, Fra89, Ste88a, Tel87k, Tel87l, Tel87m, Gro88a, Aye89, Flo89a, Tel88c]. Objective [Tel88c]. Objective-C [Tel88c]. Objects [Rov89, Valxxb, Valxxa, Flo89d]. off [Ste88b]. OLE [Broxx]. OMG [Betxx]. Open [Coh85]. OpenDoc [Rusxx]. Operating [Cap87, Hil88, McK83, Swa85r, Tur86j]. Operations [Ruz83b, Ruz83c, Ruz83d, Ruz83e]. Operators [Wil83b]. Optical [Ano88c, Wil83f]. Optimization [Row89]. Optimizer [Fox85]. Optimizing [Bau89, Gra87, McD84, Rel87a]. Order [McI86]. Oriented [Ber83, Dye89, Fra89, Gut89, Rea82, Ste88a, Tel87k, Tel87l, Tel87m, Aye89, Flo89a, Gro88a, Tel88c]. Orrfelt [Lac83, Lac83]. OS-9 [Cap87]. OS/2 [Cor87a, Dun88, Dun89a, MP89, Vos88]. OSF [Hin89]. Other [Car82]. Oubliette [Wig84e]. Output [Cha82a, Hol87q]. Overlays [McM87]. Overview [All88b, Con86]. own [AI89, Sch89a].

P [Mit83a, Sch84b]. P-Code [Mit83a]. PABX [Roh84]. Package [Bom84, Mit84, Mon82, Ber89b]. Page [Sha83]. paged [Dah89]. Panel [Sho82a, Wat89a]. panning [Cha88a]. Paradigms [Swa88c, Swa88d, Swa88e, Swa88f, Swa88g, Swa88h, Swa88i, Swa88j, Swa89b, Swa89c, Swa89d, Swa89e, Swa89f, Swa89g, Swa89h, Swa89i, Swa89j, Swa89k, Swa89l, Swa89m]. Parallel [Ash85b, Jon88, Rin85, Wil83a, Bau89, Str89]. Parameters [Sha86f]. Parentheses [Ber84a, Ber84b]. Parser [Dlu89b, Har83d, Hol84]. parsers [Lan89]. Parsing [Mot84, Sey84]. Part [Bur84a, Bur84b, Hen82b, Mil82, Mit83f, Mit83g, Ruz83e]. PASCAL [Spe83, Ame85, Ber86a, Cal87, Car85b, Flo89d, Ham86e, Hej89, Hun83, Jen85, Kie89, Kin88a, Lin87, MS84, McM87, PP85, Pot89b, Sha86f, Sor86]. Pascal/Microsoft [PP85]. Pass [Hol87d]. Passing [Hil88]. Patch [Wil85d]. Patrick [Gor82a]. Pattern [Ash85b, Bow87b, RW82, RM88, Kin89b]. Paul [Gor82a]. pay [Jon88]. PC [Ord84, Sea82, Zim87, Cor84a, Dae83, Dae84, EK86, Fun82, FM85, Gar83a, Gar83b, Gla83b, Kin85, Lin86, McN86, Mur83a, Sha86a, Smi85, Sol85, Tho89, Ung85]. PC-DOS [Cor84a]. PC/PCjr [Smi85]. PCjr [Smi85]. PCs [PP85, Tel87k]. PCX [Qui89]. PCYACC [Lan89]. PDO [Genxx]. PDP [Ham82a, Ham82b]. PDP-1802 [Ham82a, Ham82b]. PDQ [Ton89b]. Peephole [Fox85]. Performance [Flo87, Kan82b]. Persistent [Rov89]. person [Bom82c]. Personal [Dun82f, Fle84, Kos82c, Roh84]. perspective [Dah88]. Philosophy [Ham86b]. Phonetic [Pal85]. Photorealism [Ups88]. Pidgin [Ben82, Gal82, Gin82]. PIP [Wil85d]. PIP.COM [Mil84]. PISTOL [Ber83]. PL [Dov84, Rea86]. PL/68K [Rea86]. PL/C [Dov84]. Place [Nar87]. Platform [Par88]. Plays [War85]. Plot [Gay83]. Plots [Bar86]. Plotter [McW83a, Enr83]. Plotting [Smi85]. Plus [Elk89, Rus85, Sil84, Wha83, Woo85, Bur89, Cor84m]. Point [DT84, Gol84, Kaw84, Mon82, Mor84, Pri84, Smi83a, Smi85, Spe83, Dun84l, Dun84m, Fer84]. Poly [Dun89b]. Polyforth [Ruz83a]. poor [Bom82c]. Port [Zim87]. Portability [Fit88, Bla88, Fou83, Joh89]. Portable [Gin82, Rea82, Fit89, Joh89]. Portably [Ber83]. Porting [Sor86]. power [TB89]. Powers [Ham86e]. Practical [Bow83a, Ham86b, Pal85]. Practice [Bro84a]. precedence [Pet89a]. Precision [Beh83]. Precompiled [Sol83]. Prelude [Tra87c]. Preparing [Rel87b]. Preprocessor [Sch84b, Sha84]. Presentation [MP89, Sch89b]. Press [Mar87a]. Primitive [Car84, Carxxb]. Primitives [Hog85b]. Print [Dae82, Rin85]. Printer [Bad83, Poo89]. Printers [Cla85]. Printing [Bar85, Dae83, Pal85]. Prior [Gil82]. Priority [Bro87b, Hol87c]. privileges [Dun88]. Problems [Beh83]. Procedural [Sha86f]. Procedure [Ruz83a, Ber89a]. procedures [Wil85b]. Processing [Don85, Mel89]. Processor [GT82, Skj82a]. processors [Hog85a]. ProDOS [Day85, Ste85]. Products [Elk89, Dlu89c]. Professional [Ton89c]. Program [Bom84, Fit88, Gro84, Jam82, Joh84, Kat86, Lin88, Mau82a, Mor83, Pal85, PP85, RW82, Sch85a, Sti84, Tak84, Tay82, Wei83, Mak89, Sch89d, Sea82]. Programmable [Edw85, Sim83]. Programmer [ADL88, All88b, Lew88, Mea88b, Skj83b, Skj83a, Skj83c, Skj84f, Skj84h, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84g, War84, Ste89b]. Programmers [MS84, Kin88a]. Programming [All88a, Bro84b, CFH88, Der89, Dis87, Dov84, Dun89c, Dun89d, Dun89e, Dun89f, Dun89g, Dun89h, Dun89i, Dun89j, Dun89k, Dun89l, Dun89m, Dye89, Edw85, Fis89b, Ham86c, Ham86d, Ham87f, Ham87c, Ham87d, Ham87e, Hea84a, Hea84b, Hol86l, Hol87r, Mal85, Nel86, Ord82, Por88o, Por88e, Por88f, Por88g, Por88h, Por88i, Por88j, Por88k, Por88l, Por88m, Por88n, Por89a, Por89b, Por89c, Por89d, Por89e, Por89f, Por89g, Por89h, Pro84, Sch84a, Sha86b, Sha86c, Sha86d, Sha86e, Sha87a, Sha87b, Sha87c, Sha87d, Sha87e, Sha87f, Sha87g, Sha88, Sol82, Ste88a, Ste88d, Ste88e, Ste88f, Ste88g, Ste88h, Ste89c, Ste89d, Ste89e, Ste89f, Ste89g, Ste89h, Ste89i, Ste89j, Ste89k, Ste89l, Ste89m, Ste89n, Swa88c, Swa88d, Swa88e, Swa88f, Swa88g, Swa88h, Swa88i, Swa88j, Swa89b, Swa89c, Swa89d, Swa89e, Swa89f]. Programming [Swa89g, Swa89h, Swa89i, Swa89j, Swa89k, Swa89l, Swa89m, Tel87l, Tel87m, Wil85e, Dun89b, Fis89a, Flo89a, GR89a, Jon88, Por88a, Por88c, Ste88b, Ste88c]. Programs [Bre87, Cra88, Der88, Edg89, Kap84, Men89a, Pot89b, Pot89a, Siy89]. Projects [CFH88]. Prolog [Edg85, Laz89, But87b, Cor85l, Sof86, But87a, Laz88, Sch85a, Swa88a, Tel88a]. Prolog/V [Laz88]. PROLOGs [Pie87]. Proposal [Sha86g, Tur87b, KP89]. protected [Gre89a, Mar88]. Protection [Fre86]. Protocol [Lin88, Tay83, TT86]. Protocols [Bro84a, Car85a, Kra85]. Proving [Rei88]. Prowriter [Nel85]. pseudo [Ton89a]. pseudo-structures [Ton89a]. Pseudocode [Tak84, Ash84]. Ptree [Pet89a]. Public [Bur84a, Bur84b, MM84b, MM84a]. Pushing [Lev87]. Put [Sho82a]. Putting [Dah88, Nar87]. Puzzling [Ams89].

Quadtrees [Whi87]. Queues [Bro87b, Hol87c]. Quick [Lad89a, And89b, Lad89f]. QuickDraw [Der89, Mayxx]. QuickPak [Ton89c]. QuickPascal [MV89]. Quiet [App86]. Quirks [Hea84c]. Quong [Cri83].

R [Ste83, Lad89d]. R. [Ord82]. Radio [Ouv84]. RAM [Dei87, Ste85]. RAM-Cache [Dei87]. Randal [DeG84]. Random [Rus85]. Raskin [DDJ86]. RATFOR [Eve84]. Raton [Bro86]. Ray [Cor84b]. RCV [Tay83]. Read [Mes84, Rus85]. Real [Bow89, Bun89, Cra88, Tur86j, Vos88, Wil85a, GR89a, KP89, Tra88a]. Real-Time [Cra88, Tur86j, Vos88, Wil85a, Bow89, Bun89, GR89a, KP89, Tra88a]. Realizable [AS85b, AS85c, AS85a]. really [Jon88]. Reasoning [Hal84]. Receive [Hug82].

Recipes [Mar87a]. RECLAIM [Mur83b]. Reclamation [Mur83b]. recognition [Kin89b]. Recorder [Gar87]. Recursion [Eme82]. Recursive [Mit83f]. RED [Rea83a, Rea83b]. Reference [Tay82]. Regular [Hol84]. relation [Fou83]. Relocating [Bar82]. Repair [Kva82]. Repository [Con86]. Representation [But87a, Smi83a]. Res [Mur83a, Sea82]. Resident [Sil84, Ano88a]. Resolution [Rei88]. RESORT [Kra84]. Resource [Mea88b]. Resources [War84]. Rest [Mea88a]. REVAS [Wil83c]. REVASCO [Wil83c]. Review [Ams89, Blu85f, Bow83a, Bow83b, Car82, Car83, Cha82b, Cla84b, Clu85, Con85, Cri82, Cri83, Dav88, DeG84, Dlu89c, EK86, Elk89, Fre88a, Fun82, Gan83, Gom84, Gor82a, Gor82c, Hal84, HG82, Hin84, Hor85, Kin85, Lou84, Lun83, Lyo83, Mar87a, McK84, New83, Ord82, Ord84, Ouv84, Pro84, Rob82, Sch85b, Sim86, Sol85, PCC82, Ste83, Sut85e, Tra85, Ung85, Wat83, Wei89, Wig84e, Wil83c, Edw85, Por88b, Ton89b, Ton89c]. Reviews [Ashxxa, Ashxxb, Balxx, Carxxa, Casxxa, Casxxb, Cauxx, Claxxa, Claxxb, Corxx, Craxxa, Craxxb, Detxx, Dunxx, Erixx, Favxx, Gahxx, Galxx, Grixx, Harxx, Heaxx, Hutxx, Irvxx, Johxx, Jolxxa, Jolxxb, Jolxxc, Kacxx, Kapxx, Kroxx, Lanxxa, Lanxxb, Lozxx, Mitxx, Morxx, Nelxx, Parxx, Pfexx, Porxx, Reixx, Relxx, Rotxx, Rowxx, Salxx, Schxx, Shoxx, Staxx, Stexx, Swaxx, Telxx, Valxxc, Varxx, Rel88b]. Revisited [Har87c, Dal89, Por88c, Win89]. Revolution [Lyo83]. RHEALSTONE [KP89]. Riddle [Ros85, Til85]. Right [And87, Cam86c, Car89]. Ritchie [Ste89o]. Robotic [McK84]. Robotics [Blu85f]. Roll [AI89]. rollover [Dav85]. ROM [Ano88b, Nar87]. Roman [Hol87b]. Ronald [Gor82c]. Root [Cat86a, Cat85]. Roots [Cam86a]. Rotations [Goo83]. Rough [Gor82b]. Routine [Lee83, Smi85, Cat85]. routines [Hum84]. RS [Hug82]. RS-232 [Hug82]. RSA [Bur84a, Bur84b]. RSX [Gri84b, Rus85, Sil84]. Rules [Ham87f]. Run [Dal89, Zig89, And89d, Kin88b]. run-time [Kin88b]. Runic [Fra82a, Fra82b]. Running [Cop88a, Cop88b, Hol87s, Spe87d, Spe87e, Spe87f, Spe88e, Spe88f, Spe88g, Swa88k, Swa88l, Tho87, Tur86e, Tur86f, Tur86g, Tur86h, Tur86i, Tur87c, Tur87d, Tur87e, Tur87f, Sch89d, Ste88b]. Rupp [Ano88b, Ano88c, Ano88d].

S [Blu85f, Gol83, McK84, Wil85a]. S-100 [McK84, Wil85a, Gol83]. Sandewall [Pro84]. Savage [Dun84l, Dun84m, Fer84, Gol84, Kaw84, Mor84, Pri84, Spe83]. Save [Coh85]. Saville [Gan83]. SAVVY [Kin85]. Say [McC83]. SBC [Tan83]. Scale [CFH88, DeG84]. Scheduler [IB89]. Science [Har83c]. Scientific [Mar87a]. Scientist [Lyo83]. SCISTAR [FT84, Nel85, TF84]. SCOOPS [Tel87m]. Scrabble [War85]. Screen [Bla83, How83, Rea82, Rea83a, Rea83b, Woo82, Ste88c]. Screen-Oriented [Rea82]. Screens [Bre87, Mor86]. Scripts [Har87b]. scrolling [Cha88a]. SCSI [Bas85a, Bas85b, Bas85c]. SD [Ham83, Swa88a]. Sea [Car86b]. Search [Cam86b, Bar84b, Pet89a]. searches [Dlu89a, Men89b]. Searching [Flo87]. Sec [Wal85]. Secret [Fre86, Mic88]. security [Moi89]. SELDSK [Smi84]. Semantic [Rei88]. Send [Hug82, TT86, Tay83]. Send/Receive [Hug82]. Sense [Irv83]. Serial [Lin88, Wil83a]. Serial-To-Parallel [Wil83a]. Series [Gat83, Rod86]. Setting [Pet89a]. setup [Mur83a, Sea82]. Shack [Ouv84]. Shape [Ryl85]. Shasha [Ams89]. Shell [Har87b, Hen82b, Skj82b]. Shifts [Goo83]. Short [Mey83]. Shortcut [Nel85]. Show [Rod85]. Shrobe [Pro84]. SID [Mit85a]. sidebar [Tur89]. Siekmann [Hal84]. Sieve [Col83]. Signal [Mel89]. Signed [Smi83b]. Simple [Bad83, Bar86, Cla84a, May88, Mic82, Mit84, Mot84, Ruz84, Tur86j, Wat83, Eck89]. Simplified [Mau82b]. Simulated [McL89]. Simulation [CM89, GR89b]. simulations [Kin89a]. Simulator [Cat86b, Aye89]. Sinclair [Lac83, Wat83, Cha82a, Cha82b, Con82]. Sine [Cam86b]. Single [Lew88]. Sixth [Doh84, Gri84a, Gri86]. Sketching [PP85]. SKIPIF [Ram83]. Skipping [Ram83]. slots [Kru88d]. SMAL [Lou84]. SMAL/80 [Lou84]. Small [Hen82b, Kan82b, Van82, Fou83, Hal82, Hen82a, HP84, Hen85, Joh82, Kva82, Mac83, McK83, Sch84a, Sch84b, Sta83, Sta84]. Small-C [Fou83, Hal82, Hen82a, HP84, Hen85, Joh82, Kva82, Mac83, McK83, Sch84a, Sch84b, Sta83, Sta84]. Smaller [Irv83]. Smalltalk [Laz88, TB89]. Smartmodem [HG82]. smooth [Cha88a]. Soft [Rel88a]. Soft-ICE [Rel88a]. Software [Boy88, Con86, Dun82a, Dun82b, Dun82c, Dun82d, Dun83a, Dun83b, Dun83d, Dun83e, Dun83f, Dun83g, Dun83h, Dun83i, Dun83j, Dun83k, Dun83l, Dun83c, Dun84a, Dun84b, Dun84c, Dun84d, Dun84e, Dun84f, Dun84g, Dun84h, Dun84i, Dun84j, Dun85a, Dun85b, Dun85c, Dun85d, Dun85e, Dun85f, Dun85g, Dun85h, Dun85i, Dun85j, Dun86a, Dun86b, Dun86c, Dun86d, Dun86e, Dun86f, Dun86g, Dun86h, Dun86i, Dun86j, Dun87a, Dun87b, Dun87c, Dun87d, Dun87e, Dun87g, Dun87f, Ham87f, Hec85, Lew88, Rob88, Swa84i, Swa84j, Swa85n, Swa85o, Swa85p, Swa85q, Ton89c, Wig84e, Win85, Ber89b, Bla88, Kru88d, Mey89, Sti88, Dav88, Fun82, Tra85]. Solid [Ryl85]. Solutions [Hea84c, Clu85]. Some [Fre83, Gro84, Mey83, Mor84, Win85, Wis83]. Son [Fra82b]. SORCIM [New83]. Sort [Cel85, Man83]. Sorted [Dae84]. Sorting [McI86]. Sound [Lev87]. Soundex [How87]. Source [Fox85]. Sow [Cri83]. SP [Ham86a, Ham86b, Ham87b, Ham87a].

Special [KR88]. specifications [Laz89]. Speed [EK86, Rel88b, Men89c]. Speeding [Wei86]. spelling [Bom82c]. Spelunking [Lad88]. Spiecher [Cri82]. Splines [Ash86]. Spooler [Rin85]. Spreadsheet [Mor84]. SPSS [Sol85]. SPSS/PC [Sol85]. SQL [Gea88]. Square [Cam86a, Cat86a, Cat85]. SS [Ham83]. SS-SD [Ham83]. Stack [Ber83, HG82]. Stan [Rob82]. Standard [DT84, Tra87c]. Standards [Cro82, Ham86b, Sha86g, Wis83, Fou83]. Star [Sho82b]. Star-to-Epson [Sho82b]. State [Bow87b, Por88b, Smi89]. statements [Fox89]. States [Bow83b]. Statistical [Hol87d]. Step [Wil86]. Still [Gri83]. Stock [Mes84]. Stone [Har83c]. Storage [Ano88c]. Streams [GS89, Ode84]. String [Bar84b, Lad89b, Men89b]. Strings [McD84]. Stroustrup [Ste89o]. structure [Gro84]. Structured [Dun89c, Dun89d, Dun89e, Dun89f, Dun89g, Dun89h, Dun89i, Dun89j, Dun89k, Dun89l, Dun89m, Ham86c, Ham86d, Ham87f, Ham87c, Ham87d, Ham87e, Por88o, Por88e, Por88f, Por88g, Por88h, Por88i, Por88j, Por88k, Por88l, Por88c, Por88m, Por88n, Por89h, Sha84, Sha86b, Sha86c, Sha86d, Sha86e, Sha87a, Sha87b, Sha87c, Sha87d, Sha87e, Sha87f, Sha87g, Sha88, Dun89b, Por88a]. Structures [Sha86g, Kin88b, Ton89a]. Study [Mel89]. Style [Sol82]. Submit [Ram83]. Subroutine [Mes84]. Subroutines [Skj82b, Gan83]. Subset [Mit83b, Mit83c, Mit83d, Mit83e]. Subversive [Cor85a]. Suggestion [Ber84b]. Summit [Tra85]. SuperScripsit [Ouv84]. Survey [Tho86]. Swaine [Eri88i, Eri88j, Swa86j, Swa86k, Swa86l, Swa86m, Swa86n, Swa86o, Swa86p, Swa87j, Swa87k, Swa87l, Swa87m, Swa87n, Swa87o, Swa87p, Swa87q, Swa87r, Swa87s, Swa87t, Swa87u, Swa88m, Swa88n, Swa88o, Swa88p, Swa88q, Swa88r, Swa88s, Swa88t, Swa88u, Swa88v, Swa89n, Swa89o, Swa89p, Swa89q, Swa89r, Swa89s, Swa89t, Swa89u, Swa89v, Swa89w, Swa89x, Swa89y]. SWAP [Mak89]. Symbols [TF84, FT84]. Syntax [Gla83a]. Synthesizer [Sim83]. System [Ano88a, Bur84a, Bur84b, Bur85, Camxx, Cap87, Cor84b, GPxx, Kan82a, Kos82c, McK83, Mon83, Sch84a, Sil84, Sof86, TT84a, TT84b, Tur86j, Bur89, Hin89, Moi89, Pet89b, Get89a, Sim86]. Systems [Cri82, Gri87b, Gri87a, Hil88, Hor85, KR88, Par84, PCC82, Wei89, Fre88a, Lun83].

table [Dav85]. Tables [Pir83, Ber89a]. Tackling [CFH88]. Taligent [GPxx]. Talking [Cor84b]. Target [RMB85]. Taste [TT87]. TAWK [Eck89]. Tax [Sch85a]. Tech [Cla84b]. Technical [Cor82a]. Techniques [Dis87, Hea84a, Hea84b, Hea85c]. Technologies [Broxx, Kin85, Rel88a]. Technology [Leb88, Row89]. TEE [Hea85a]. Telecommunications [Bro84b]. Teletext [Bow83b]. template [Sch89b]. ten [Edw85]. Terminal [Dav85]. Terminals [Ash83]. Test [Gal82, Ste86]. Tests [Mor84, Dav85]. Text [McW83b, Nic82, TT87]. Theorem [Rei88]. Theory [Bro84a]. Third [DeG84]. Thoughts [Fit88]. Threaded [Bro85, Gor82c, Gor83a, Mat88]. Threaded-Code [Bro85]. Thrills [EK86]. TIFF [BMO88]. Time [Bom82a, Cra88, Ros89, Tur86j, Vos88, Wil85a, And89d, Bow89, Bun89, GR89a, KP89, Kin88b, Tra88a]. Timed [IB89]. Timex [Lac83, Wat83]. Timex/Sinclair [Wat83]. Tiny [AS85a, McW83b]. Tips [Hea85c]. TNZ [Cam86d]. Today [Rel87c]. Token [Mot84]. Tool [Woo85, Mic88]. Toolbox [ADL88, Dun82a, Dun82b, Dun82c, Dun82d, Dun83a, Dun83b, Dun83d, Dun83e, Dun83f, Dun83g, Dun83h, Dun83i, Dun83j, Dun83k, Dun83l, Dun83c, Dun84a, Dun84b, Dun84c, Dun84d, Dun84e, Dun84f, Dun84g, Dun84h, Dun84i, Dun84j, Dun85a, Dun85b, Dun85c, Dun85d, Dun85e, Dun85f, Dun85g, Dun85h, Dun85i, Dun85j, Dun86a, Dun86b, Dun86c, Dun86d, Dun86e, Dun86f, Dun86g, Dun86h, Dun86i, Dun86j, Dun87a, Dun87b, Dun87c, Dun87d, Dun87e, Dun87g, Dun87f, Por87a, Por87b, Sca86, Bas85a, Bas85b, Bas85c]. Toolkit [Cal87, Wei89]. Tools [Gea88, Jen85, Swa88a]. Topics [TT88]. TOPS [Fis84]. TOPS-20 [Fis84]. TopView [Swa85r]. tour [Cor85l]. Trace [Jon83]. Trade [Fre86]. Transfer [Hug82]. transfers [Men89c]. Transform [Ull84, Bar84a]. Translating [Qui89]. Translation [Ord82]. Translations [Car83]. Translator [Cam86d, Sca83]. transputers [Swa88a]. Tree [Jam82, BD83, DB83]. Trees [Mat88]. trials [Rel88b]. TRIM [Hea85b]. Trio [Fre88a]. TRS [Cro84, Ouv84, Sca83, Car82]. TRS-80 [Cro84, Ouv84, Sca83, Car82]. TRTA [Ste86, Tur86a, Tur87b]. True [Sha87h]. TSD [Hec85, Hog85a, Win85]. TSR [Men89a, Pot89b, Pot89a, Wei89]. TSX [Tan83]. Turbo [PP85, Ame85, Ber86a, Cal87, CV89, EK86, Flo89d, Hej89, Jen85, Lad89a, Lad89f, Lin87, McM87, Por87a, Por87b, Pot89b, Sha86f, Tel88a]. Turing [Sut85e]. Two [FM85, Mus87, Swa89a, TB89]. Two-Bit [Mus87]. TXS [Tan83]. Tydeman [Bow83b]. Types [Jon87, Hej89].

U. [Blu85f]. U.S. [McK84]. Ultimate [Rin85, Sti84]. Undocumented [Jam89, Lad88]. United [Bow83b]. units [Gro88a]. UNIX [Cra88, GS89, Hol87q, Mea88a, Sch85c, Sch85d, Skj82b, Skj83a, Skj83c, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84g, Bas84, Col84, Fre88b, Har87b, Hol84, Poo89, Rog84, Skj83b, Skj84f, Skj84h, Wal84, Wil83d]. UNIX-Compatible [Hol87q]. Unix-Like [Hol84]. Unix-To-Unix [Col84]. Unixes [Wal84]. Unstructured [Wil85e]. Update [Car85b, Hen85]. Upon [Mau82a]. USART [Gay85]. use [Mer89]. User [Car88, Ste83, Hea82d]. Users [Nel85]. Using [Bom84, Bow87b, Bre87, Dae83, Gay83, Har87b, Hog85b, Kin89b, Ouv82, Par88, Rei88, Ruz83a, Sti88, Tak84, Tho89, Dlu89a, Get89a]. Utilities [Col84, Hug82, Nel87, Hym88a]. Utility [Bar85, Dae82, Gar83a, Kva82, Mur83b, Ram83, Wil83a, Mak89]. UX [Mea88a].

V [Clu85, Laz88]. v.2 [Fou83, Hen82a, Mac83]. V1.3a [Cri83]. V2.2 [Wil85d]. Validation [Ste88i]. Vallee [Lyo83]. Variable [Fis89a, Mar86b, Mis89a]. Variable-level [Fis89a]. Variables [Hog85b, Moo88a]. Varieties [Wal84]. Vector [Mic82, Shi83]. VEdit [Elk89]. Vendor [Ste88i]. Ver [Dlu89c, Elk89, Fre88a, Rel88a, Wil85b]. VERSA [Lun83]. VERSA-FILE [Lun83]. Version [Ame85, Gol84, Hor85, Kin85, Sch85b, Tra85, Ung85, Ban84]. Versus [Car85b, Lad89f]. Very [DeG84]. VGA [Cha88a, Cha88b]. via [Duv88, Hug82]. Victor [Ano83]. Video [Cha82a]. Videotex [Bow83b]. Virtual [Fle84, Tic88, Dah89]. Visual [Dye89]. VMS [Poo89]. Vocabularies [Sha83]. VolksMicro [PCC82]. Volkswriter [Fun82]. VOS [Hen82b]. Votrax [McC83]. VP [App86, But87b, Erd87, Hol86m, Hol86l, Hol87r, Sum87]. vs [Cor82a, Fre88b, Spe88a].

W [Ord84]. want [Dun89b]. Wars [Nel87]. Watcom [Dlu89c]. Way [CM89, TT86]. Ways [Bli87]. Weather [Par84]. Western [Sut85e]. Westico [Lou84]. Where [Tur87a]. while [Sch89d]. Wildcards [Skj82b]. Wilson [Hal84]. Winchester [Gol83]. Window [Get89a, Mit84]. Windowing [Hol87q, Swa85r, Get89a, Hin89, Joh89]. Windows [JS89, Lin86, Swa85r]. Winston [Gor82a]. Winthrop [Gan83]. Wirth [Por88o]. WizardCopy [Gas89, Tur89]. word [Hog85a]. Words [Ode83]. Wordstar [Fis84, FT84, TF84]. Work [Sho82a]. Workshop [All88b]. Workstation [Leb88]. Worm [Ste86]. WP6502 [Cri83]. Write [Cri82]. Write-On [Cri82]. Writeable [Gla83a]. Writing [Cra88, Der88, Dun88, Fit89, Fra82b, Fra89, Kle87, Mey89, Mis89b, Wat89b]. Written [Edg85, Par86]. Wrong [Car86a, Sum87].

X [Fit89, Get89a]. X/GEM [Fit89]. Xanadu [Gre83]. XCMD [Kru88d]. XCMDs [And89b]. Xerox [Tel87j]. XFCN [Kru88d]. XMODEM [Tay83, Smi89]. Xor [Cor87b, Har87c].

Years [Har83c]. Yes [Jon83]. Yin [Wat83].

Z [Sim86].  Z-System [Sim86].  Z80

[Gan83, Gin82, Goo83, Sca83, Til85]. Z8000 [Ode82]. Zarella [Car83, Cla84b]. Zoomracks [Hec85]. ZSID [Mit85a]. Zwimpfer [Bow83b]. ZX81 [Cha82a, Cha82b, Con82].

References

Ackerman:1987:H

[Ack87] Mark S. Ackerman. 6502 hacks. Dr. Dobb’s Journal of Software Tools, 12(2):24–26, 29–32, February 1987. CODEN DDJOEB. ISSN 0888-3076.

Al-Dhelaan:1988:PDM

[ADL88] Abdullah Al-Dhelaan and Ted G. Lewis. A programmer’s database for the Macintosh toolbox. Dr. Dobb’s Journal of Software Tools, 13(1):24–26, 28–31, 33, 35, 36, 41, 48–52, 54, 57–65, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1989:RYO

[AI89] Michael Abrash and Dan Illowsky.

Roll your own minilanguages with mini-interpreters. Dr. Dobb’s Journal of Software Tools, 14(9): 52–58, 61–62, 102–103, September 1989. CODEN DDJOEB. ISSN 0888-3076.

Allen:1988:IHP

[All88a] Daniel Allen. An introduction to HyperCard programming. Dr. Dobb’s Journal of Software Tools, 13(??):56–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Allen:1988:OMP

[All88b] Daniel Allen. Overview: Macintosh programmer’s workshop. Dr. Dobb’s Journal of Software Tools, 13(??):20–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Amelang:1983:DDD

[Ame83] Loren Amelang. Dysan’s digital diagnostic diskette. Dr. Dobb’s Journal of Software Tools, 8(12): 40–??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Ames:1985:TPB

[Ame85] William Ames. Turbo Pascal bug in version 2.0 [letter]. Dr. Dobb’s Journal of Software Tools, 10(5):9– ??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Amsterdam:1988:AEL

[Ams88a] Jonathan Amsterdam. Adding extensions to Lisp. Dr. Dobb’s Journal of Software Tools, 13(10):19– 20, 22, 24, 27, 56, 58, October 1988. CODEN DDJOEB. ISSN 0888-3076.

Amsterdam:1988:CAL

[Ams88b] Jonathan Amsterdam. Creating an adventurous language. Dr. Dobb’s Journal of Software Tools, 13(4):18–??, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Amsterdam:1989:RPA

[Ams89] Jonathan Amsterdam. Review of The Puzzling Adventures of Dr. Ecco, by Dennis Shasha, in ER. Dr. Dobb’s Journal of Software Tools, 14(4):96–??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1985:BMM

[And85] Brian R. Anderson. Bit manipulation in Modula-2. Dr. Dobb’s Journal of Software Tools, 10(11): 38–??, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1986:CAa

[And86a] Brian R. Anderson. A 68000 cross assembler. Dr. Dobb’s Journal of Software Tools, 11(4):??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1986:CAb

[And86b] Brian R. Anderson. A 68000 cross assembler. Dr. Dobb’s Journal of Software Tools, 11(5):??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1986:CAc

[And86c] Brian R. Anderson. A 68000 cross assembler. Dr. Dobb’s Journal of Software Tools, 11(6):??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1987:WRH

[And87] Brian R. Anderson. What’s right with high-level languages? Dr. Dobb’s Journal of Software Tools, 12(6):14–??, June 1987. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1989:KMM

[And89a] Brian R. Anderson. Kermit meets Modula-2. Dr. Dobb’s Journal of Software Tools, 14(5):22–36, 83– 93, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Anderson:1989:QDX

[And89b] Jay Martin Anderson. Quick Drawing with XCMDs (Hyper Card). Dr. Dobb’s Journal of Software Tools, 14(5):63–66, 104, 106, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Anderson:1989:CCM

[And89c] Paul Anderson. C customized memory allocators. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):62– 66, 94, 1989. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1989:CMA

[And89d] Paul Anderson. C multidimensional arrays at run time. Dr. Dobb’s Journal of Software Tools, 14(8):50, 52, 54–57, 59–60, 124, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Anonymous:1983:CEC

[Ano83] Anonymous. CP/M-80 expansion card for the Victor 9000. Dr. Dobb’s Journal of Software Tools, 8(2):58–61, February 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Anonymous:1985:CCM

[Ano85] Anonymous. C compilers for MS-DOS. Dr. Dobb’s Journal of Software Tools, 10(8):30–??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1988:RES

[Ano88a] Anonymous. The RESIDENT EX-PERT system. Dr. Dobb’s Journal of Software Tools, 13(2):29, February 1988. CODEN DDJOEB. ISSN 1044-789X. Advertisement.

Anonymous:1988:RBCa

[Ano88b] Anonymous. Rupp brothers catalog: CD-ROM drives. Dr. Dobb’s Journal of Software Tools, 13(2):3, February 1988. CO-DEN DDJOEB. ISSN 1044-789X. Mailed with Dr. Dobb’s Journal #136 (February, 1988).

Anonymous:1988:RBCb

[Ano88c] Anonymous. Rupp brothers catalog: Compact disc optical storage. Dr. Dobb’s Journal of Software Tools, 13(2):4, February 1988. CO-DEN DDJOEB. ISSN 1044-789X. Mailed with Dr. Dobb’s Journal #136 (February, 1988).

Anonymous:1988:RBCc

[Ano88d] Anonymous. Rupp brothers catalog: IZE. Dr. Dobb’s Journal of Software Tools, 13(2):4, February 1988. CODEN DDJOEB. ISSN 1044-789X. Mailed with Dr. Dobb’s Journal #136 (February, 1988).

Appleman:1986:KIQ

[App86] Daniel Appleman. Keep it quiet in VP. Dr. Dobb’s Journal of Software Tools, 11(5):12–??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Albrecht:1985:THR

[AS85a] Bob Albrecht and M. Swaine. Tiny hackers: a realizable fantasy in Festschrift for Doctor Dobb. Dr. Dobb’s Journal of Software Tools, 10(2):30–??, February 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Albrecht:1985:RFLa

[AS85b] Bob Albrecht and Michael Swaine.

Realizable fantasies: Liberating the Mac. Dr. Dobb’s Journal of Software Tools, 10(4):12–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Albrecht:1985:RFLb

[AS85c] Bob Albrecht and Michael Swaine.

Realizable fantasies: Liberating the Mac. Dr. Dobb’s Journal of Software Tools, 10(7):92–??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Ashdown:1983:CCD

[Ash83] Ian E. Ashdown. Cursor control for dumb terminals. Dr. Dobb’s Journal of Software Tools, 8(12):32–34, 36, 33, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Ashdown:1984:MP

[Ash84] Ian E. Ashdown. Merge pseudocode for. Dr. Dobb’s Journal of Software Tools, 9(11):89– ??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Ashdown:1985:AFC

[Ash85a] Ian E. Ashdown. Archiving files with CP/M-80 and CP/M-86. Dr. Dobb’s Journal of Software Tools, 10(1):36–??, January 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Ashdown:1985:PPM

[Ash85b] Ian E. Ashdown. Parallel pattern matching and Fgrep. Dr. Dobb’s Journal of Software Tools, 10(9): 46–??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Ashdown:1986:CFC

[Ash86] Ian E. Ashdown. Curve fitting with cubic splines. Dr. Dobb’s Journal of Software Tools, 11(9): 24–??, September 1986. CODEN DDJOEB. ISSN 1044-789X.

Ashdown:19xx:R

[Ashxxa] Ian E. Ashdown. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Ashworth:19xx:R

[Ashxxb] Robert Ashworth. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Ayers:1989:OLS

[Aye89] Kenneth E. Ayers. An objectoriented logic simulator. Dr. Dobb’s Journal of Software Tools, 14(12):72, 75–76, 78, 130–132, 134, 136–138, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Baden:1983:SGP

[Bad83] Wil Baden. Simple graphics for printer. Dr. Dobb’s Journal of Software Tools, 8(11):86– 90, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Ballinger:19xx:R

[Balxx] Charles K. Ballinger. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Banchy:1984:CMM

[Ban84] Patrick Banchy. CLINK mods to May83 82 and improved version. Dr. Dobb’s Journal of Software Tools, 9(1):83–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Barnhart:1982:RLF

[Bar82] Joe Barnhart. A relocating loader in Forth. Dr. Dobb’s Journal of Software Tools, 7(9):34– ??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Barnhart:1983:FM

[Bar83] Joe Barnhart. Forth and the Motorola 68000. Dr. Dobb’s Journal of Software Tools, 8(9):18–20, 22, 24–26, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Barnhart:1984:FFF

[Bar84a] Joe Barnhart. Forth and the Fast Fourier Transform. Dr. Dobb’s Journal of Software Tools, 9(9): 34–??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Barton:1984:SSF

[Bar84b] Paul C. Barton. String search function in C [letter]. Dr. Dobb’s Journal of Software Tools, 9(8): 8–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Barnhart:1985:PUH

[Bar85] Joe Barnhart. A printing utility for the HP LaserJet in 16BST. Dr. Dobb’s Journal of Software Tools, 10(9):117–??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Barkakati:1986:SPE

[Bar86] Nabajyoti Barkakati. Simple plots with the enhanced graphics adapter. Dr. Dobb’s Journal of Software Tools, 11(5):42–??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Bass:1984:UDD

[Bas84] John L. Bass. Unix device drivers. Dr. Dobb’s Journal of Software Tools, 9(12):38–42, 44, 46–48, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Bass:1985:SYMa

[Bas85a] John Bass. SCSI your Mac, in MAC Toolbox. Dr. Dobb’s Journal of Software Tools, 10(9):94– ??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Bass:1985:SYMb

[Bas85b] John Bass. SCSI your Mac, in MAC Toolbox. Dr. Dobb’s Journal of Software Tools, 10(10):106–??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Bass:1985:SYMc

[Bas85c] John Bass. SCSI your Mac, in MAC Toolbox. Dr. Dobb’s Journal of Software Tools, 10(12):110– ??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Bauer:1989:OPE

[Bau89] Barr E. Bauer. Optimizing in a parallel environment. Dr. Dobb’s Journal of Software Tools, 14(11): 72, 74, 76, 79, 80, 82, 122– 123, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Bartholomew:1983:BTI

[BD83] Alan Bartholomew and Chris Deppe. B-Tree ISAM concepts. Dr. Dobb’s Journal of Software Tools, 8(6):18–??, June 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Behler:1983:PNP

[Beh83] Allan L. Behler, Jr. Problems with numeric precision in BASIC (in DDC). Dr. Dobb’s Journal of Software Tools, 8(11):10–??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Bender:1982:PUC

[Ben82] Andrew L. Bender. Pidgin for the 8080 under CP/M. Dr. Dobb’s Journal of Software Tools, 7(3): 24–??, March 1982. CODEN DDJOEB. ISSN 1044-789X.

Bergmann:1983:PFP

[Ber83] Ernest E. Bergmann. PISTOL — a Forth-Like portably implemented stack oriented language. Dr. Dobb’s Journal of Software Tools, 8(2):12, 14–15, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Bergman:1984:LP

[Ber84a] Ernest E. Bergman. Languages and parentheses. Dr. Dobb’s Journal of Software Tools, 9(7):102– 106, 108, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Bergmann:1984:LPS

[Ber84b] Ernest E. Bergmann. Languages and parentheses — A suggestion for Forth-like languages. Dr. Dobb’s Journal of Software Tools, 9(7):102–??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Bergmann:1986:CTP

[Ber86a] Ernest E. Bergmann. Concurrency and Turbo Pascal. Dr. Dobb’s Journal of Software Tools, 11(3):36–??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Berry:1986:MK

[Ber86b] Ken Berry. A multitasking kernel. Dr. Dobb’s Journal of Software Tools, 11(12):16–??, December 1986. CODEN DDJOEB. ISSN 0888-3076.

Berens:1989:CPT

[Ber89a] Timothy Berens. C procedure tables. Dr. Dobb’s Journal of Software Tools, 14(8):68, 70, 72–73, 128–132, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Bergman:1989:FLA

[Ber89b] Noel J. Bergman. First look at CommonView (software package). Dr. Dobb’s Journal of Software Tools, 14(10):74, 76, 78–80, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Betz:19xx:OC

[Betxx] Mark Betz. OMG’s CORBA. Dr. Dobb’s Journal of Software Tools, ??(??):8–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Bianchi:1989:MMM

[Bia89] Curt Bianchi. Memory management with MacApp. Dr. Dobb’s Journal of Software Tools, 14(??): 23–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Brown:1987:AA

[BL87] Richard E. Brown and Steve Ligett. Async Appletalk. Dr. Dobb’s Journal of Software Tools, 12(10):18–20, 22, 24–26, 28, 60, 62, 65–72, 74–76, 78–84, 86, 88–91, October 1987. CODEN DDJOEB. ISSN 0888-3076.

Blair:1982:BBA

[Bla82] Robert Blair. Binary (8-bit)

to ASCII hexadecimal conversion [8088/8086 assembly language]. Dr. Dobb’s Journal of Software Tools, 7(12):15–??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Blakeslee:1983:FDF

[Bla83] Thomas Blakeslee. Forth debugging from a full-screen editor [letter]. Dr. Dobb’s Journal of Software Tools, 8(9):10–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Blackham:1988:BSP

[Bla88] Greg Blackham. Building software for portability. Dr. Dobb’s Journal of Software Tools, 13(12): 18–26, December 1988. CODEN DDJOEB. ISSN 0888-3076.

Blinn:1987:HMW

[Bli87] James F. Blinn. How many ways can you draw a circle? Dr. Dobb’s Journal of Software Tools, 12(9): 18–20, 22, 24–26, September 1987. CODEN DDJOEB. ISSN 0888-3076.

Blum:1982:MCa

[Blu82a] Robert Blum. More on checksums. Dr. Dobb’s Journal of Software Tools, 7(4):36–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Blum:1982:MCb

[Blu82b] Robert Blum. More on checksums. Dr. Dobb’s Journal of Software Tools, 7(7):44–??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMDa

[Blu83a] Robert Blum. CP/M disk I/O in CPME. Dr. Dobb’s Journal of Software Tools, 8(3):??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMDb

[Blu83b] Robert Blum. CP/M disk I/O in CPME. Dr. Dobb’s Journal of Software Tools, 8(4):??, April 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEa

[Blu83c] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 8(5):??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEb

[Blu83d] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 8(6):??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEc

[Blu83e] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(7):??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEd

[Blu83f] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(8):??, August 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEe

[Blu83g] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(9):??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEf

[Blu83h] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(10):??, October 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEg

[Blu83i] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(11):??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1983:CMEh

[Blu83j] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(12):??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEa

[Blu84a] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 9(1):??, January 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEb

[Blu84b] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(2):??, February 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEc

[Blu84c] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(3):??, March 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEd

[Blu84d] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(4):??, April 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEe

[Blu84e] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(5):??, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEf

[Blu84f] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(7):??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEg

[Blu84g] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(9):??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEh

[Blu84h] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 9(10):??, October 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEi

[Blu84i] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Blum:1984:CMEj

[Blu84j] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 9(12):??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Blum:1985:CMEa

[Blu85a] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 10(2):??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Blum:1985:CMEb

[Blu85b] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 10(6):??, June 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1985:CMEc

[Blu85c] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 10(8):??, August 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Blum:1985:CMEd

[Blu85d] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 10(10):??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Blum:1985:CMEe

[Blu85e] Robert Blum. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 10(11):??, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Blum:1985:URM

[Blu85f] Robert Blum. U. S. Robotics modem review. Dr. Dobb’s Journal of Software Tools, 10(11):102– ??, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Blum:1986:CME

[Blu86] Robert Blum. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 11(1):??, January 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Budiansky:1988:HIF

[BMO88] Michael Budiansky, Anthony Meadow, and Rocky Offner. Handling image files with TIFF. Dr. Dobb’s Journal of Software Tools, 13(5):26–??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Bolstad:1983:CMB

[Bol83] Terje Bolstad. CP/M BDOS and BIOS calls for C. Dr. Dobb’s Journal of Software Tools, 8(6):22–??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Bomberger:1982:EMW

[Bom82a] Alan Bomberger. Excuse me what time is it? Dr. Dobb’s Journal of Software Tools, 7(2):50–??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Bomberger:1982:ICM

[Bom82b] Alan Bomberger. Interrupts and CP/M. Dr. Dobb’s Journal of Software Tools, 7(12):54–??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Bomberger:1982:PPS

[Bom82c] Alan Bomberger. A poor person’s spelling checker. Dr. Dobb’s Journal of Software Tools, 7(4):42–53, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Bomberger:1984:PMP

[Bom84] Alan Bomberger. Program monitor package — using interrupts to instrument applications. Dr. Dobb’s Journal of Software Tools, 9(11):26–34, 36–40, 42–46, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Bowers:1983:RPI

[Bow83a] Susan Bowers. Review of Practical Introduction to Computer Graphics, by Ian O. Angell. Dr. Dobb’s Journal of Software Tools, 8(8): 114–??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Bowers:1983:RTV

[Bow83b] Susan Bowers. Review of Teletext and Videotex in the United States, by Tydeman, Lipinski, Adler, Nyan, and Zwimpfer. Dr. Dobb’s Journal of Software Tools, 8(12):119–??, December 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Bowman:1987:B

[Bow87a] Charles F. Bowman. Backtracking. Dr. Dobb’s Journal of Software Tools, 12(8):24–??, August 1987. CODEN DDJOEB. ISSN 0888-3076.

Bowman:1987:PMU

[Bow87b] Charles F. Bowman. Pattern matching using finite state machines. Dr. Dobb’s Journal of Software Tools, 12(10):46–??, October 1987. CODEN DDJOEB. ISSN 0888-3076.

Bowling:1989:RMM

[Bow89] D. Bowling. Real-time modeling with MS-DOS. Dr. Dobb’s Journal of Software Tools, 14(2):26–28, 30, 32, 34, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Boyd:1988:SEE

[Boy88] Stowe Boyd. Software engineering environments. Dr. Dobb’s Journal of Software Tools, 13(??):8–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Breeden:1987:UEG

[Bre87] J. Brooks Breeden. Using EGA graphics screens in your programs. Dr. Dobb’s Journal of Software Tools, 12(11):46–??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Brooks:1984:CPT

[Bro84a] Leslie Brooks. Communications protocols — theory and practice. Dr. Dobb’s Journal of Software Tools, 9(2):26–29, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Broughton:1984:BLT

[Bro84b] R. S. Broughton. BASIC language telecommunications programming. Dr. Dobb’s Journal of Software Tools, 9(2):64–69, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Brodie:1985:TMB

[Bro85] Leo Brodie. A threaded-code microprocessor bursts Forth. Dr. Dobb’s Journal of Software Tools, 10(10):42–50, October 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Brown:1986:BBR

[Bro86] Robert Jay Brown, III. BRIE: The Boca Raton inference engine. Dr. Dobb’s Journal of Software Tools, 11(4):24–??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Brown:1987:ANN

[Bro87a] Robert Jay Brown. Artificial neural network experiment. Dr. Dobb’s Journal of Software Tools, 112(4):16–??, April 1987. CODEN DDJOEB. ISSN 0888-3076.

Brown:1987:EAL

[Bro87b] Robert Jay Brown, III. An efficient algorithm for large priority queues. Dr. Dobb’s Journal of Software Tools, 12(6):16–19, June 1987. CODEN DDJOEB. ISSN 0888-3076.

Brockschmidt:19xx:OIT

[Broxx] Kraig Brockschmidt. OLE integration technologies. Dr. Dobb’s Journal of Software Tools, ??(??): 42–??, October 19xx. CODEN DDJOEB. ISSN 1044-789X.

Bunnell:1989:RDA

[Bun89] Michael Bunnell. Real-time data acquisition. Dr. Dobb’s Journal of Software Tools, 14(6):36–38, 42– 44, 86, 88, 90, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Burk:1983:FM

[Bur83] Ronald L. Burk. Function macros for the 8086/88 in 16BST. Dr. Dobb’s Journal of Software Tools, 8(10):98–??, October 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Burton:1984:RPKa

[Bur84a] Charles E. Burton. RSA: A public key cryptography system part I. Dr. Dobb’s Journal of Software Tools, 9(3):16–??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Burton:1984:RPKb

[Bur84b] Charles E. Burton. RSA: A public key cryptography system part II. Dr. Dobb’s Journal of Software Tools, 9(4):32–??, April 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Burton:1985:EAC

[Bur85] Charles E. Burton. An enhanced ADFGVX cipher system. Dr. Dobb’s Journal of Software Tools, 10(2):48–??, February 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Burke:1989:APS

[Bur89] Chris Burke. APL PLUS system II. Dr. Dobb’s Journal of Software Tools, 14(2):72, 74, 76, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Butrick:1987:LKR

[But87a] Richard Butrick. Logic and knowledge representation in Prolog. Dr. Dobb’s Journal of Software Tools, 12(7):30–32, 34–35, 37–38, July 1987. CODEN DDJOEB. ISSN 0888-3076.

Butrick:1987:LPV

[But87b] Richard Butrick. Logic and PRO-LOG in VP. Dr. Dobb’s Journal of Software Tools, 12(1):14–??, January 1987. CODEN DDJOEB. ISSN 1044-789X.

Buvel:1984:FNC

[Buv84] Raymond Buvel. A Forth native-code cross compiler for the MC68000. Dr. Dobb’s Journal of Software Tools, 9(9):68– ??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Cadmus:1984:FMF

[Cad84] Ray Cadmus. File maintenance in Forth. Dr. Dobb’s Journal of Software Tools, 9(9):24–26, 28, 30– 33, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Callihan:1987:GTT

[Cal87] Hubert D. Callihan. A graphics toolkit for Turbo Pascal. Dr. Dobb’s Journal of Software Tools, 12(11):38–??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Campbell:1986:NSR

[Cam86a] R. A. Campbell. NS32000 square roots. Dr. Dobb’s Journal of Software Tools, 11(3):122–123, 106, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Campbell:1986:SS

[Cam86b] Richard A. Campbell. In search of a sine. Dr. Dobb’s Journal of Software Tools, 11(12):30–32, December 1986. CODEN DDJOEB. ISSN 1044-789X.

Campbell:1986:RA

[Cam86c] Richard A. Campbell. The right to assemble. Dr. Dobb’s Journal of Software Tools, 11(3):??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Campbell:1986:TBB

[Cam86d] Richard A. Campbell. TNZ: An 8bit to 16-bit translator. Dr. Dobb’s Journal of Software Tools, 11(10): 40–??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Campagnoni:19xx:ISO

[Camxx] F. R. Campagnoni. IBM’s system object model. Dr. Dobb’s Journal of Software Tools, ??(??):24– ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Capouch:1987:OOS

[Cap87] Brian Capouch. OS-9 operating system. Dr. Dobb’s Journal of Software Tools, 12(1):30– 32, 34–36, January 1987. CODEN DDJOEB. ISSN 0888-3076.

Cardell:1982:RMB

[Car82] James Cardell. Review of Microsoft BASIC Decoded and Other Mysteries for the TRS-80, by James Farvour. Dr. Dobb’s Journal of Software Tools, 7(4):60–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Carter:1983:RLT

[Car83] Michael Carter. Review of Language Translations, by John Zarella. Dr. Dobb’s Journal of Software Tools, 8(12):116–??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Carter:1984:API

[Car84] Michael Carter. Adding primitive I/O functions to Mulisp. Dr. Dobb’s Journal of Software Tools, 9(11):18–24, November 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Carroll:1985:AP

[Car85a] David W. Carroll. Asynchronous protocols. Dr. Dobb’s Journal of Software Tools, 10(8):92–95, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Carroll:1985:MVP

[Car85b] David W. Carroll. Modula-2 versus Pascal for microcomputers: an update. Dr. Dobb’s Journal of Software Tools, 10(11):28– 34, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Carew:1986:WWC

[Car86a] David Carew. What’s wrong with C. Dr. Dobb’s Journal of Software Tools, 11(6):12–??, June 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Carter:1986:FGS

[Car86b] Everett Carter. Forth goes to sea. Dr. Dobb’s Journal of Software Tools, 11(7):40–??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Carr:1988:DU

[Car88] Robert Carr. Developing for the user. Dr. Dobb’s Journal of Software Tools, 13(5):18–??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Carew:1989:WRC

[Car89] David Carew. What’s right with C? Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):96–??, Winter 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Carroll:19xx:R

[Carxxa] David W. Carroll. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Carter:19xx:APF

[Carxxb] Michael Carter. Adding primitive I/O functions to muLISP. Dr. Dobb’s Journal of Software Tools, ??(11):18–??, November 19xx. CODEN DDJOEB. ISSN 1044-789X.

Cashton:19xx:R

[Casxxa] Dennis Cashton. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Castle:19xx:R

[Casxxb] Tom Castle. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Cathey:1985:ISR

[Cat85] James Cathey. 68000 integer square root routine in 16BST. Dr. Dobb’s Journal of Software Tools, 10(5):118–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Cathey:1986:LEI

[Cat86a] J. Cathey. Letter to the editor [integer square root]. Dr. Dobb’s Journal of Software Tools, 11(8): 14, 82–85, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Cathey:1986:CSM

[Cat86b] James Cathey. COM: An 8080 simulator for the MC68000. Dr. Dobb’s Journal of Software Tools, 11(1):76–??, January 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Caulkins:19xx:R

[Cauxx] David Caulkins. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Celko:1985:BNS

[Cel85] Joseph Celko. Bose-Nelson sort.

Dr. Dobb’s Journal of Software Tools, 10(9):68–??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Courington:1988:TLS

[CFH88] William Courington, Jonathan Feiber, and Masahiro Honda. Tackling large-scale programming projects. Dr. Dobb’s Journal of Software Tools, 13(??):45–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Coye:1984:MMC

[CG84] Keith Coye and Alvin Grossman.

Micro to mainframe connection. Dr. Dobb’s Journal of Software Tools, 9(2):20–??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Chamkis:1982:ADV

[Cha82a] Jerry Chamkis. Add direct video output to [the] Sinclair ZX81. Dr.

Dobb’s Journal of Software Tools, 7(5):48–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Chamkis:1982:RSZ

[Cha82b] Jerry Chamkis. Review of the Sinclair ZX81. Dr. Dobb’s Journal of Software Tools, 7(4):59–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Chavez:1987:FF

[Cha87] Lori Chavez. Fast Forth for the 68000. Dr. Dobb’s Journal of Software Tools, 12(10):32–34, 36, October 1987. CODEN DDJOEB. ISSN 0888-3076.

Chalk:1988:EVS

[Cha88a] Andrew J. Chalk. EGA and VGA smooth scrolling and panning. Dr. Dobb’s Journal of Software Tools, 13(12):62–69, 101–110, December 1988. CODEN DDJOEB. ISSN 0888-3076.

Chalk:1988:IFC

[Cha88b] Andrew J. Chalk. An Italic font in C for the EGA and VGA. Dr. Dobb’s Journal of Software Tools, 13(3):36–??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Clark:1984:SCC

[Cla84a] David D. Clark. Simple calculations with complex numbers. Dr. Dobb’s Journal of Software Tools, 9(10):30–??, October 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Clark:1984:RHT

[Cla84b] Robert Clark. Review of High-Tech Consulting, by John Zarella. Dr. Dobb’s Journal of Software Tools, 9(6):80–??, June 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Clark:1985:FCE

[Cla85] David D. Clark. fx80char: A character editor for Epson FX-80 printers. Dr. Dobb’s Journal of Software Tools, 10(4):28–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Clark:19xx:Ra

[Claxxa] David D. Clark. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Clark:19xx:Rb

[Claxxb] Dean Clark. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Clute:1985:RML

[Clu85] Alan Clute. Review of Mac-FORTH level 1 V. 1.2; level 2 V. 2.1; level 3 V. 3.0 from Creative Solutions. Dr. Dobb’s Journal of Software Tools, 10(10):100–??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Coppola:1989:CAN

[CM89] Jean F. Coppola and Francis T.

Marchese. Cellular automata: A new way of simulation. Dr. Dobb’s Journal of Software Tools, 14(9): 34, September 1989. CODEN DDJOEB. ISSN 1044-789X.

Cohen:1985:FOS

[Coh85] Michael Cohen. File “open” and “save” functions in C for the Macintosh in 16BST. Dr. Dobb’s Journal of Software Tools, 10(3):96–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Colburn:1983:BAS

[Col83] Donald Colburn. A better algorithm for sieve benchmarks in Forth [letter]. Dr. Dobb’s Journal of Software Tools, 8(9):9– ??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Coleman:1984:UNU

[Col84] Ron Coleman. Unix-To-Unix network utilities. Dr. Dobb’s Journal of Software Tools, 9(2):30– 31, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Colby:1986:Ia

[Col86a] Wendelin Colby. Of interest. Dr. Dobb’s Journal of Software Tools, 11(2):??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Colby:1986:Ib

[Col86b] Wendelin Colby. Of interest. Dr. Dobb’s Journal of Software Tools, 11(4):??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Colby:1986:Ic

[Col86c] Wendelin Colby. Of interest. Dr. Dobb’s Journal of Software Tools, 11(5):??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Colby:1986:Id

[Col86d] Wendelin Colby. Of interest. Dr. Dobb’s Journal of Software Tools, 11(6):??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Colby:1986:Ie

[Col86e] Wendelin Colby. Of interest. Dr. Dobb’s Journal of Software Tools, 11(7):??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Colby:1986:If

[Col86f] Wendelin Colby. Of interest. Dr.

Dobb’s Journal of Software Tools, 11(8):??, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Connell:1982:NBS

[Con82] John Connell, Jr. Novice builds the Sinclair ZX81. Dr. Dobb’s Journal of Software Tools, 7(11): 28–??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Conn:1985:RAL

[Con85] Richard Conn. Review of AM-PRO Little Board and Bookshelf computers. Dr. Dobb’s Journal of Software Tools, 10(5):106–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Conn:1986:ODA

[Con86] Richard Conn. Overview of DOD Ada software repository. Dr. Dobb’s Journal of Software Tools, 11(2):60–??, February 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Copeland:1988:RLa

[Cop88a] Ron Copeland. Running light. Dr.

Dobb’s Journal of Software Tools, 13(9):??, September 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Copeland:1988:RLb

[Cop88b] Ron Copeland. Running light. Dr.

Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:CMV

[Cor82a] David E. Cortesi. CP/M-86 vs. MS-DOS: A technical comparison. Dr. Dobb’s Journal of Software Tools, 7(7):14–??, July 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCa

[Cor82b] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(1):??, January 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCb

[Cor82c] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(3):??, March 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCc

[Cor82d] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(4):??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCd

[Cor82e] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(6):??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCe

[Cor82f] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(7):??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCf

[Cor82g] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(8):??, August 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCg

[Cor82h] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(9):??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCh

[Cor82i] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(10):??, October 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCi

[Cor82j] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(11):??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:DDCj

[Cor82k] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 7(12):??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:Ia

[Cor82l] David E. Cortesi. Of interest. Dr.

Dobb’s Journal of Software Tools, 7(2):??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1982:Ib

[Cor82m] David E. Cortesi. Of interest.

Dr. Dobb’s Journal of Software Tools, 7(4):??, April 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCa

[Cor83a] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(2):??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCb

[Cor83b] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(3):??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCc

[Cor83c] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(4):??, April 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCd

[Cor83d] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(5):??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCe

[Cor83e] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(6):??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCf

[Cor83f] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(7):??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCg

[Cor83g] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(8):??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCh

[Cor83h] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(9):??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCi

[Cor83i] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(10):??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCj

[Cor83j] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(11):??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1983:DDCk

[Cor83k] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 8(12):??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Cornell:1984:FPI

[Cor84a] David Cornell. Forth to PC-DOS interface. Dr. Dobb’s Journal of Software Tools, 9(1):44–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCc

[Cor84b] D. E. Cortesi. Dr. Dobb’s Clinic: System languages; keep talking; move over, Ray. Dr. Dobb’s Journal of Software Tools, 9(2): 10–13, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCa

[Cor84c] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(1):??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCb

[Cor84d] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(2):??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCd

[Cor84e] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(3):??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCe

[Cor84f] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(6):??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCf

[Cor84g] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(7):??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCg

[Cor84h] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(8):??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCh

[Cor84i] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(9):??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCi

[Cor84j] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(10):??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCj

[Cor84k] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:DDCk

[Cor84l] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 9(12):??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1984:WDF

[Cor84m] David E. Cortesi. What’s the diff?:

A file comparator for CP/M Plus. Dr. Dobb’s Journal of Software Tools, 9(8):30–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDS

[Cor85a] David E. Cortesi. Dr. Dobb is a subversive. Dr. Dobb’s Journal of Software Tools, 10(5):22–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCa

[Cor85b] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(1):??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCb

[Cor85c] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(2):??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCc

[Cor85d] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(4):??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCd

[Cor85e] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(5):??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCe

[Cor85f] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(6):??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCf

[Cor85g] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(7):??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCg

[Cor85h] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(8):??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCh

[Cor85i] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(9):??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCi

[Cor85j] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(10):??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:DDCj

[Cor85k] David E. Cortesi. Dr. Dobb’s Clinic. Dr. Dobb’s Journal of Software Tools, 10(12):??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1985:TP

[Cor85l] David E. Cortesi. A tour of PRO-LOG. Dr. Dobb’s Journal of Software Tools, 10(3):44–63, March 1985. CODEN DDJOEB. ISSN 1044-789X. TK 7885 A1 D6.

Cortesi:1987:DL

[Cor87a] David E. Cortesi. Dynamic linking in OS/2. Dr. Dobb’s Journal of Software Tools, 12(12):18–??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Cortesi:1987:XC

[Cor87b] David E. Cortesi. Xor chain. Dr. Dobb’s Journal of Software Tools, 12(6):28–30, 34–35, 38, June 1987. CODEN DDJOEB. ISSN 0888-3076.

Cortesi:19xx:R

[Corxx] David E. Cortesi. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Cramer:1988:WRT

[Cra88] William D. Cramer. Writing realtime programs under UNIX. Dr. Dobb’s Journal of Software Tools, 13(6):18–??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Crayne:19xx:Ra

[Craxxa] Charles Crayne. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Crayne:19xx:Rb

[Craxxb] Diane Crayne. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Crimmins:1982:RWS

[Cri82] Jerry Crimmins. Review of writeon from Spiecher systems and datamost. Dr. Dobb’s Journal of Software Tools, 7(7):65–??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Criscione:1983:RWV

[Cri83]  James Criscione.  Review of

WP6502 V1.3a from Dwo Quong Fok Lok Sow. Dr. Dobb’s Journal of Software Tools, 8(11):92– ??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Crockford:1982:SCM

[Cro82] Douglas Crockford. Standards for computerized measurement. Dr. Dobb’s Journal of Software Tools, 7(2):57–??, February 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Crockett:1984:BT

[Cro84] Davy Crockett. BASICFMT for TRS-80. Dr. Dobb’s Journal of Software Tools, 9(4):60–75, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Custer:1989:Ia

[Cus89a] Janna Custer. Of interest. Dr.

Dobb’s Journal of Software Tools, 14(8):??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Custer:1989:Ib

[Cus89b] Janna Custer. Of interest. Dr.

Dobb’s Journal of Software Tools, 14(9):??, September 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1989:Ic

[Cus89c] Janna Custer. Of interest. Dr.

Dobb’s Journal of Software Tools, 14(10):??, October 1989. CODEN DDJOEB. ISSN 1044-789X.

Custer:1989:Id

[Cus89d] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 14(11):??, November 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1989:Ie

[Cus89e] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 14(12):??, December 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1989:BBM

[Cus89f] Janna Custer. On being or becoming a Macintosh developer. Dr. Dobb’s Journal of Software Tools, 14(??):54–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Catchings:1989:GBT

[CV89] William Catchings and Mark L. Van Name. Getting the bugs out with Turbo Debugger. Dr. Dobb’s Journal of Software Tools, 14(7): 64–70, 73, July 1989. CODEN DDJOEB. ISSN 0888-3076.

Daetwyler:1982:PUC

[Dae82] Dan Daetwyler. A print utility for CP/M. Dr. Dobb’s Journal of Software Tools, 7(6):36–??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Daetwyler:1983:PGU

[Dae83] Dan Daetwyler. Printing graphics using the IBM PC. Dr. Dobb’s Journal of Software Tools, 8(6):32– ??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Daetwyler:1984:SDD

[Dae84] Dan Daetwyler. Sorted diskette directory listing for the IBM PC. Dr.

Dobb’s Journal of Software Tools, 9(1):60–74, January 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Dahlgren:1988:PGI

[Dah88] Kent Dahlgren. Putting graphical interfaces into perspective. Dr. Dobb’s Journal of Software Tools, 13(11):32–41, November 1988. CODEN DDJOEB. ISSN 0888-3076.

Dahlgren:1989:DPV

[Dah89] Kent Dahlgren. Demand paged virtual memory. Dr. Dobb’s Journal of Software Tools, 14(4):32–37, 41, April 1989. CODEN DDJOEB. ISSN 0888-3076.

Daley:1989:RLE

[Dal89] Phil Daley. Run length encoding revisited. Dr. Dobb’s Journal of Software Tools, 14(5):130, 154, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Davis:1985:TKR

[Dav85] Charles Davis. Terminal keyrollover tests table of in DDC. Dr. Dobb’s Journal of Software Tools, 10(5):16–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Davis:1988:RCC

[Dav88] Randy Davis. Review of Check-Mate from Cinnabar Software. Dr. Dobb’s Journal of Software Tools, 13(2):134–??, February 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Day:1985:ACC

[Day85] Shawn Day. Adding a COPY command to ProDOS. Dr. Dobb’s Journal of Software Tools, 10(12):

54–??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Deppe:1983:BIC

[DB83] Chris Deppe and Alan Bartholomew. B-Tree ISAM concepts. Dr. Dobb’s Journal of Software Tools, 8(6):18– 21, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Editors:1986:HID

[DDJ86] DDJ Editors. Human interface design: Jef Raskin interview. Dr. Dobb’s Journal of Software Tools, 11(5):32–??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

DeGroot:1984:RTC

[DeG84] Douglas DeGroot. Review of third Caltech conference on very large scale integration, Ed. Randal E. Bryant. Dr. Dobb’s Journal of Software Tools, 9(1):89–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Deikman:1987:RCM

[Dei87] Alan Deikman. A RAM-Cache manager in C. Dr. Dobb’s Journal of Software Tools, 12(12):30– ??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Derossi:1988:WPM

[Der88] Chris Derossi. Writing programs for MultiFinder. Dr. Dobb’s Journal of Software Tools, 13(12):46– 54, 96–100, December 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Derossi:1989:PCQ

[Der89] Chris Derossi. Programming with color QuickDraw. Dr. Dobb’s Journal of Software Tools, 14(??):

8–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Dettmann:19xx:R

[Detxx] Terry R. Dettmann. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Dilworth:1988:AF

[Dil88] John B. Dilworth. An 80386 assembler in Forth. Dr. Dobb’s Journal of Software Tools, 13(10):28– 30, 33–35, 60, 63, 65–66, 68–70, 72, 81–82, October 1988. CODEN DDJOEB. ISSN 0888-3076.

Disque:1987:ALP

[Dis87] Tom Disque. 8088 assemblylanguage programming techniques. Dr. Dobb’s Journal of Software Tools, 12(7):24–26, 28, July 1987. CODEN DDJOEB. ISSN 0888-3076.

Dlugosz:1989:EDS

[Dlu89a] John M. Dlugosz. Extended directory searches using C++. Dr. Dobb’s Journal of Software Tools, 14(3):55–56, 58, 99–101, March 1989. CODEN DDJOEB. ISSN 0888-3076.

Dlugosz:1989:HCP

[Dlu89b] John M. Dlugosz. A homebrew C++ parser. Dr. Dobb’s Journal of Software Tools, 14(12):40, 43– 44, 46, 116, 118–119, 120, 122, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Dlugosz:1989:RWC

[Dlu89c] John M. Dlugosz. Review of Watcom C ver. 7.0 from Watcom Products. Dr. Dobb’s Journal of Software Tools, 14(9):74–80, September 1989. CODEN DDJOEB. ISSN 0888-3076.

Doherty:1983:CFG

[Doh83] Michael J. Doherty. Comments on “fifth generation computers” [Dec82 60]. Dr. Dobb’s Journal of Software Tools, 8(3):75–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Doherty:1984:CSG

[Doh84] Michael J. Doherty. Comments on “sixth generation computers” [May84 37]. Dr. Dobb’s Journal of Software Tools, 9(6):70–??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Donais:1985:CMC

[Don85] Roger E. Donais. CP/M-68K conditional batch processing. Dr. Dobb’s Journal of Software Tools, 10(12):70–??, December 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Dovey:1984:IPP

[Dov84] Morris Dovey. Introduction to PL/C: Programming language for compilers. Dr. Dobb’s Journal of Software Tools, 9(3):44–??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:FSF

[DT84] Ray Duncan and Martin Tracy. The FVG standard floating-point extension. Dr. Dobb’s Journal of Software Tools, 9(9):110– 115, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1982:BSTa

[Dun82a] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 7(9):??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1982:BSTb

[Dun82b] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 7(10):??, October 1982. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1982:BSTc

[Dun82c] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 7(11):??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1982:BSTd

[Dun82d] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 7(12):??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1982:FA

[Dun82e] Ray Duncan. FORTH 8086 assembler. Dr. Dobb’s Journal of Software Tools, 7(2):14–??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1982:GIP

[Dun82f] Ray Duncan. Graphics on IBM’s personal computer. Dr. Dobb’s Journal of Software Tools, 7(7):32– ??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTa

[Dun83a] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(1):??, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTb

[Dun83b] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(2):??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:ST

[Dun83c] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(2):18, 20–25, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTc

[Dun83d] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(3):??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTd

[Dun83e] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(4):??, April 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTe

[Dun83f] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(5):??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTf

[Dun83g] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(6):??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTg

[Dun83h] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(7):??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTh

[Dun83i] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(9):??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTi

[Dun83j] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(10):??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTj

[Dun83k] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(11):??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1983:BSTk

[Dun83l] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 8(12):??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTa

[Dun84a] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(1):??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTb

[Dun84b] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(2):??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTc

[Dun84c] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(3):??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTd

[Dun84d] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(4):??, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTe

[Dun84e] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(6):??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTf

[Dun84f] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(7):??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTg

[Dun84g] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(8):??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTh

[Dun84h] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(10):??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTi

[Dun84i] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:BSTj

[Dun84j] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 9(12):??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1984:CFF

[Dun84k] Ray Duncan. Converting Fig-Forth to Forth-83. Dr. Dobb’s Journal of Software Tools, 9(5):28– 32, 34–35, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Dunford:1984:SFPa

[Dun84l] Christopher J. Dunford. Savage floating-point benchmark in 8088/ 8087 assembly language in 16BST. Dr. Dobb’s Journal of Software Tools, 9(7):116–??, July 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Dunford:1984:SFPb

[Dun84m] Christopher J. Dunford. Savage floating-point benchmark in Modula-2 in 16BST. Dr. Dobb’s Journal of Software Tools, 9(8): 106–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTa

[Dun85a] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(1):??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTb

[Dun85b] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(3):??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTc

[Dun85c] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(4):??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTd

[Dun85d] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(5):??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTe

[Dun85e] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(6):??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTf

[Dun85f] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(7):??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTg

[Dun85g] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(8):??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTh

[Dun85h] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(9):??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTi

[Dun85i] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(11):??, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1985:BSTj

[Dun85j] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 10(12):??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTa

[Dun86a] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(1):??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTb

[Dun86b] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(2):??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTc

[Dun86c] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(4):??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTd

[Dun86d] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(5):??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTe

[Dun86e] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(6):??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTf

[Dun86f] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(7):??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTg

[Dun86g] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(8):??, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTh

[Dun86h] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(9):??, September 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTi

[Dun86i] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(10):??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1986:BSTj

[Dun86j] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 11(12):??, December 1986. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1987:BSTa

[Dun87a] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(2):??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1987:BSTb

[Dun87b] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(3):??, March 1987. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1987:BSTc

[Dun87c] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(5):??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1987:BSTd

[Dun87d] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(6):??, June 1987. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1987:BSTe

[Dun87e] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(7):??, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1987:ST

[Dun87f] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(7):106–108, 110–111, July 1987. CODEN DDJOEB. ISSN 0888-3076.

Duncan:1987:BSTf

[Dun87g] Ray Duncan. 16-bit software toolbox. Dr. Dobb’s Journal of Software Tools, 12(8):??, August 1987. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1988:WOA

[Dun88] Ray Duncan. Writing OS/2 applications with I/O privileges. Dr. Dobb’s Journal of Software Tools, 13(12):36–44, December 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Duncan:1989:ICO

[Dun89a] Ray Duncan. Interprocess communications in OS/2. Dr. Dobb’s Journal of Software Tools, 14(6): 14–15, 18, 20, 24–25, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Duntemann:1989:PWO

[Dun89b] J. Duntemann. Poly want an object? (structured programming). Dr. Dobb’s Journal of Software Tools, 14(11):142, 145–150, 154, 156–159, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Duntemann:1989:SPa

[Dun89c] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(2):??, February 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPb

[Dun89d] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(3):??, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPc

[Dun89e] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(4):??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPd

[Dun89f] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(5):??, May 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPe

[Dun89g] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(6):??, June 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPf

[Dun89h] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(7):??, July 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPg

[Dun89i] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(8):??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPh

[Dun89j] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(9):??, September 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPi

[Dun89k] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(10):??, October 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPj

[Dun89l] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(11):??, November 1989. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1989:SPk

[Dun89m] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(12):??, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Duncan:19xx:R

[Dunxx] Ray Duncan. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Duvanenko:1988:ICC

[Duv88] Victor J. Duvanenko. Image compression via compilation. Dr. Dobb’s Journal of Software Tools, 13(11):42–47, 82–90, November 1988. CODEN DDJOEB. ISSN 0888-3076.

Duvanenko:1989:IM

[Duv89] Victor J. Duvanenko. Image mathematics. Dr. Dobb’s Journal of Software Tools, 14(7):45–55, 92– 94, July 1989. CODEN DDJOEB. ISSN 0888-3076.

Dye:1989:VOO

[Dye89] Rob Dye. Visual object-oriented programming. Dr. Dobb’s Journal of Software Tools, 14(??):31– ??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Eckel:1989:TSI

[Eck89] Bruce Eckel. TAWK: a simple interpreter in C++. Dr. Dobb’s Journal of Software Tools, 14(5): 50–60, 98, 100–102, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Edgar:1985:CWP

[Edg85] G. A. Edgar. A compiler written in prolog. Dr. Dobb’s Journal of Software Tools, 10(5):84–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Edgar:1989:DCP

[Edg89] Robert Edgar. Debugging C programs. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):56–60, 93, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Edlin:1986:HID

[Edl86] James Edlin. Human interface design: From the outside in. Dr. Dobb’s Journal of Software Tools, 11(5):24–??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Edwards:1985:PEP

[Edw85] Mark U. Edwards. Programming editors programmable editors: a review of ten. Dr. Dobb’s Journal of Software Tools, 10(11):60– ??, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Elkins:1986:HST

[EK86] Michael Elkins and Stephen King. High-speed thrills: A review of eight Turbo boards for the IBM PC. Dr. Dobb’s Journal of Software Tools, 11(9):46–??, September 1986. CODEN DDJOEB. ISSN 1044-789X.

Elkins:1989:RVP

[Elk89] T. A. Elkins. Review of VEdit plus ver. 3 from CompuView products. Dr. Dobb’s Journal of Software Tools, 14(8):90–??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Emery:1982:RF

[Eme82] Glyn Emery. Recursion in Forth. Dr. Dobb’s Journal of Software Tools, 7(2):12–??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Enright:1983:EP

[Enr83] Michael T. Enright. Ellipse plotter. Dr. Dobb’s Journal of Software Tools, 8(12):19–??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Erdelsky:1987:ICV

[Erd87] Philip J. Erdelsky. Impossibilities in computers in VP. Dr. Dobb’s Journal of Software Tools, 12(5):14–??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Ea

[Eri88a] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 13(5):??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Eb

[Eri88b] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 13(6):??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Ec

[Eri88c] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 13(7):??, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Ed

[Eri88d] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 13(8):??, August 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Ee

[Eri88e] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 13(9):??, September 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Ef

[Eri88f] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Eg

[Eri88g] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 13(11):??, November 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1988:Eh

[Eri88h] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 13(12):??, December 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1988:SFa

[Eri88i] Jonathan Erickson. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1988:SFb

[Eri88j] Jonathan Erickson. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(11):??, November 1988. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ea

[Eri89a] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 14(1):??, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Eb

[Eri89b] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 14(2):??, February 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ec

[Eri89c] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(3):??, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ed

[Eri89d] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(4):??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ee

[Eri89e] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(5):??, May 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ef

[Eri89f] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(6):??, June 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Eg

[Eri89g] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(7):??, July 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Eh

[Eri89h] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(8):??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ei

[Eri89i] Jonathan Erickson. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(9):??, September 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ej

[Eri89j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 14(10):??, October 1989. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1989:Ek

[Eri89k] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 14(11):??, November 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1989:El

[Eri89l] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 14(12):??, December 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:19xx:R

[Erixx] Jonathan Erickson. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Evenden:1984:R

[Eve84] Gerald I. Evenden. RATFOR [letter]. Dr. Dobb’s Journal of Software Tools, 9(6):10–??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Favitta:19xx:R

[Favxx] Michael Favitta. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, December 19xx. CO-DEN DDJOEB. ISSN 1044-789X.

Ferguson:1984:SFP

[Fer84] Kenneth M. Ferguson. Savage floating-point benchmark deficiencies of in 16BST. Dr. Dobb’s Journal of Software Tools, 9(8):107–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Fey:19x3:CDA

[Feyx3] Edgar H. Fey, Jr. Compiler design, A fundamental mistake of — guest essay. Dr. Dobb’s Journal of Software Tools, ??(3):71–??, March 19x3. CODEN DDJOEB. ISSN 1044-789X.

Fusina:1982:IA

[FG82] Luca Fusina and Claudio Granuzzo. Interfacing the 68000 to an AIM 65. Dr. Dobb’s Journal of Software Tools, 7(1):12–??, January 1982. CODEN DDJOEB. ISSN 1044-789X.

Firmage:19xx:NAD

[Firxx] Joseph Firmage. Novell’s App-Ware distributed bus. Dr. Dobb’s Journal of Software Tools, ??(??): 50–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Fisher:1984:EWT

[Fis84] Steven Fisher. Emulate Wordstar on TOPS-20. Dr. Dobb’s Journal of Software Tools, 9(1):76–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Fischer:1989:VP

[Fis89a] R. Fischer. Variable-level programming. Dr. Dobb’s Journal of Software Tools, 14(6):46, 48, 50, 52, 54–55, 92, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Fischer:1989:FPF

[Fis89b] Ronald Fischer. Functional programming and FPCA ’89. Dr. Dobb’s Journal of Software Tools, 14(12):96–??, December 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Fitler:1988:FMT

[Fit88] William Fitler. A few more thoughts on program portability. Dr. Dobb’s Journal of Software Tools, 13(12):20–??, December 1988. CODEN DDJOEB. ISSN 1044-789X.

Fitler:1989:WPA

[Fit89] William Fitler. Writing portable applications with X/GEM. Dr. Dobb’s Journal of Software Tools, 14(3):38, 40–41, 92, March 1989. CODEN DDJOEB. ISSN 0888-3076.

Fleming:1984:VPC

[Fle84] Jim Fleming. Virtual personal computer. Dr. Dobb’s Journal of Software Tools, 9(2):32–??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Fletcher:1987:MF

[Fle87] G. Yates Fletcher. Mini Forth for the 68000. Dr. Dobb’s Journal of Software Tools, 12(1):22–24, 28, 29, January 1987. CODEN DDJOEB. ISSN 0888-3076.

Floyd:1987:HHS

[Flo87] Edwin T. Floyd. Hashing for highperformance searching. Dr. Dobb’s Journal of Software Tools, 12(2): 34–??, February 1987. CODEN DDJOEB. ISSN 0888-3076.

Floyd:1989:CAO

[Flo89a] Michael A. Floyd. A class act (object-oriented programming). Dr. Dobb’s Journal of Software Tools, 14(4):58–64, April 1989. CODEN DDJOEB. ISSN 0888-3076.

Floyd:1989:E

[Flo89b] Michael A. Floyd. Editorial. Dr.

Dobb’s Journal of Software Tools, 14(12):??, December 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1989:MCC

[Flo89c] Michael A. Floyd. Making the Cto-Fortran connection. Dr. Dobb’s Journal of Software Tools, 14(8): 22–23, 25–27, 102–104, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Floyd:1989:TPO

[Flo89d] Michael A. Floyd. Turbo Pascal with objects. Dr. Dobb’s Journal of Software Tools, 14(7):56–63, 95– 97, July 1989. CODEN DDJOEB. ISSN 0888-3076.

Furuta:1985:RTT

[FM85] Richard Furuta and Pierre A.

MacKay. Two TEX implementations for the IBM PC. Dr. Dobb’s Journal of Software Tools, 10(9): 80–89, 91, September 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Foulk:1983:SCC

[Fou83] Richard Foulk. Small-C compiler v.2 standards and portability in relation to [letter]. Dr. Dobb’s Journal of Software Tools, 8(3):11–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Fox:1985:POA

[Fox85] David L. Fox. A peephole optimizer for assembly language source code. Dr. Dobb’s Journal of Software Tools, 10(8):56–??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Fox:1989:BCS

[Fox89] David L. Fox. Benchmarking C statements. Dr. Dobb’s Journal of Software Tools, 14(2):60, 62, 65, 98–103, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Franz:1982:RIE

[Fra82a] Marty Franz. Runic: An interactive extensible compiler. Dr. Dobb’s Journal of Software Tools, 7(3):12–??, March 1982. CODEN DDJOEB. ISSN 1044-789X.

Franz:1982:SRW

[Fra82b] Marty Franz. Son of runic: Writing the compiler. Dr. Dobb’s Journal of Software Tools, 7(4):54–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Franz:1989:WFO

[Fra89] Marty Franz. Writing filters in an object-oriented language. Dr. Dobb’s Journal of Software Tools, 14(12):28, 30, 32, 34–36, 112, 114– 115, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Freed:1983:BMN

[Fre83] Edwin E. Freed. Binary magic numbers — some applications and algorithms. Dr. Dobb’s Journal of Software Tools, 8(4):24–??, April 1983. CODEN DDJOEB. ISSN 1044-789X.

Fressola:1986:CTS

[Fre86] Alfred A. Fressola. Copyright and trade secret protection. Dr. Dobb’s Journal of Software Tools, 11(3):12–??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Freeman:1988:RCI

[Fre88a] Neil Freeman. Review of C-INDEX+ ver. 3.1 from Trio Systems. Dr. Dobb’s Journal of Software Tools, 13(6):128–??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Frey:1988:UVU

[Fre88b] Donnalyn Frey. Unix vs. Unix. Dr.

Dobb’s Journal of Software Tools, 13(12):28–34, December 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Fletcher:1984:SGM

[FT84] Guy Fletcher and John Thomas.

SCISTAR: Greek and math symbols with Wordstar. Dr. Dobb’s Journal of Software Tools, 9(8): 26–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

FunkHouser:1982:RVI

[Fun82] Robert K. FunkHouser. Review of Volkswriter for the IBM PC from Lifetree Software. Dr. Dobb’s Journal of Software Tools, 7(7):61– ??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Gahlon:19xx:R

[Gahxx] Dean Gahlon. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Gale:1982:TP

[Gal82] William A. Gale. Test for pidgin. Dr. Dobb’s Journal of Software Tools, 7(11):18–??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Gallagher:19xx:R

[Galxx] Steve Gallagher. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Gans:1983:RZA

[Gan83] Paul J. Gans. Review of Z80 Assembly Language Subroutines, by Lance A. Leventhal and Winthrop Saville. Dr. Dobb’s Journal of Software Tools, 8(12):117–??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Garfinkel:1983:FUI

[Gar83a] Simson L. Garfinkel. FLIP utility for the IBM PC. Dr. Dobb’s Journal of Software Tools, 8(5):85– ??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Garfinkel:1983:GLI

[Gar83b] Simson L. Garfinkel. The Game of Life on the IBM-PC. Dr. Dobb’s Journal of Software Tools, 8(6):42– ??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Garinger:1984:BBA

[Gar84] James Garinger. Binary (16-bit) to ASCII hex conversion in 16BST. Dr. Dobb’s Journal of Software Tools, 9(10):87–??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Garvin:1987:DMR

[Gar87] Mark Garvin. Designing a music recorder. Dr. Dobb’s Journal of Software Tools, 12(5):22–23, 26, May 1987. CODEN DDJOEB. ISSN 0888-3076.

Gasper:1988:GHH

[Gas88] Elon Gasper. Getting A head with HyperAnimation. Dr. Dobb’s Journal of Software Tools, 13 (7):18–??, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Gaspar:1989:WFB

[Gas89] Don Gaspar. WizardCopy for fast backups. Dr. Dobb’s Journal of Software Tools, 14(??):45–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Gates:1983:SEF

[Gat83] Wendall C. Gates. Series expansion in Forth. Dr. Dobb’s Journal of Software Tools, 8(6):52–??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Gay:1983:UEB

[Gay83] Don Gay. Using the Epson bit plot graphics. Dr. Dobb’s Journal of Software Tools, 8(10):82–??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Gay:1985:CUC

[Gay85] Don Gay. C USART controller.

Dr. Dobb’s Journal of Software Tools, 10(6):60–??, June 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Geary:1988:SDT

[Gea88] Michael Geary. SQL development tools. Dr. Dobb’s Journal of Software Tools, 13(7):36–??, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Gengle:1985:IAI

[Gen85] Dean Gengle. Information age issues. Dr. Dobb’s Journal of Software Tools, 10(6):52–??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Gentry:19xx:DAN

[Genxx] Dennis Gentry. Distributed applications and NeXT’s PDO. Dr. Dobb’s Journal of Software Tools, ??(??):58–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Gettys:1989:NWU

[Get89a] James Gettys. Network windowing using the X Window System. Dr. Dobb’s Journal of Software Tools, 14(3):42, 44, 49–50, 53, March 1989. CODEN DDJOEB. ISSN 0888-3076.

Gettys:1989:ILA

[Get89b] Tom Gettys. Implementing the LRU algorithm. Dr. Dobb’s Journal of Software Tools, 14(4):34–??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Gilmore:1982:PC

[Gil82] John Gilmore. Prior censorship. Dr. Dobb’s Journal of Software Tools, 7(5):59–??, May 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Gintis:1982:PPZ

[Gin82] Herbert Gintis. Pidgin the portable — Z80 macro-assembly implementation. Dr. Dobb’s Journal of Software Tools, 7(10):25–??, October 1982. CODEN DDJOEB. ISSN 1044-789X.

Glass:1983:TMW

[Gla83a] Harvey Glass. Towards a more writeable Forth syntax. Dr. Dobb’s Journal of Software Tools, 8(11):

80–83, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Glass:1983:MBP

[Gla83b] James Glass. More benchmarks for the PC [letter]. Dr. Dobb’s Journal of Software Tools, 8(5):7–??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Goldman:1983:AWH

[Gol83] Oscar Goldman. Attaching a Winchester hard disk to the S-100 bus. Dr. Dobb’s Journal of Software Tools, 8(10):71–74, 76–80, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Goldberg:1984:LVS

[Gol84] Morton Goldberg. LISP version of the Savage floating point benchmark in 16BST. Dr. Dobb’s Journal of Software Tools, 9(6):82–??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Gomez:1984:RNC

[Gom84] A. Gomez. Review of NCI CO-HERENT from Network Consulting. Dr. Dobb’s Journal of Software Tools, 9(12):106–??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Goodman:1983:SRZ

[Goo83] Ron Goodman. Shifts and rotations on the Z80. Dr. Dobb’s Journal of Software Tools, 8(5):64–??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Gordon:1982:RLP

[Gor82a] Brian Gordon. Review of LISP, by Patrick Henry Winston and Berthold Klaus Paul Horn. Dr. Dobb’s Journal of Software Tools, 7(2):60–??, February 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Gordon:1982:ANA

[Gor82b] H. T. Gordon. Approximation numerical algorithms — rough approximation. Dr. Dobb’s Journal of Software Tools, 7(7):54–56, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Gordon:1982:RTI

[Gor82c] H. T. Gordon. Review of Threaded Interpretive Languages, by Ronald G. Loeliger. Dr. Dobb’s Journal of Software Tools, 7(4):60–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Gordon:1983:TC

[Gor83a] H. T. Gordon. 6809 threaded code. Dr. Dobb’s Journal of Software Tools, 8(3):62–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Gordon:1983:FDA

[Gor83b] H. T. Gordon. Fast divisibility algorithms. Dr. Dobb’s Journal of Software Tools, 8(6):14–??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Grimes:19xx:ATS

[GPxx] Jack Grimes and Mike Potel. The architecture of the Taligent system. Dr. Dobb’s Journal of Software Tools, ??(??):36–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Gehani:1989:CCR

[GR89a] Narain H. Gehani and William D.

Roome. Concurrent C for real-time programming. Dr. Dobb’s Journal of Software Tools, 14(11):38, 40, 43–45, 100, November 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Gehani:1989:DES

[GR89b] Narain H. Gehani and William D.

Roome. Discrete event simulation in Concurrent C. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):24–28, 30, 71–72, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Grappel:1987:OIM

[Gra87] Robert D. Grappel. Optimizing integer multiplications by constant multipliers. Dr. Dobb’s Journal of Software Tools, 12(3):34–35, 37, March 1987. CODEN DDJOEB. ISSN 0888-3076.

Gregory:1983:XHF

[Gre83] Roger Gregory. Xanadu — hypertext from the future. Dr. Dobb’s Journal of Software Tools, 8(1):28– 31, 34–35, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Green:1989:PMM

[Gre89a] Thomas Green. 80386 protected mode and multitasking. Dr. Dobb’s Journal of Software Tools, 14(9):64–72, 109–113, September 1989. CODEN DDJOEB. ISSN 0888-3076.

Green:1989:CMK

[Gre89b] Thomas Green. A C++ multitasking kernel. Dr. Dobb’s Journal of Software Tools, 14(2):45–46, 48, 51, 84, 86, 88, 90, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Grigonis:1982:GCG

[Gri82] Richard Grigonis. 5th generation computers — guest essay. Dr. Dobb’s Journal of Software Tools, 7(12):60–??, December 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Grigonis:1983:SMF

[Gri83] Richard Grigonis. And still more fifth generation computers! Dr. Dobb’s Journal of Software Tools, 8(8):54–??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Grigonis:1984:SGC

[Gri84a] Richard Grigonis. Sixth generation computers. Dr. Dobb’s Journal of Software Tools, 9(5):37–46, 48– 49, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Griswold:1984:CMB

[Gri84b] Michael Griswold. CP/M 2.2 BIOS RSX in CPME. Dr. Dobb’s Journal of Software Tools, 9(7):22–??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Grigonis:1986:SGM

[Gri86] Richard Grigonis. Sixth generation minds. Dr. Dobb’s Journal of Software Tools, 11(2):12–??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Grigonis:1987:MLE

[Gri87a] Richard Grigonis. MYCIN-Like expert systems. Dr. Dobb’s Journal of Software Tools, 12(4):42–??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Grigonis:1987:MES

[Gri87b] Richard W. Grigonis. Mycin-Like expert systems. Dr. Dobb’s Journal of Software Tools, 112(4):42– ??, April 1987. CODEN DDJOEB. ISSN 0888-3076.

Grinzo:19xx:R

[Grixx] Lou Grinzo. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Grosberg:1984:SIP

[Gro84] John A. Grosberg. Some ideas on program structure [letter] in CUPN. Dr. Dobb’s Journal of Software Tools, 9(12):97–??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Grosberg:1988:ODU

[Gro88a] J. A. Grosberg. Object-oriented dimensional units. Dr. Dobb’s Journal of Software Tools, 13(9): 52–62, 94–99, September 1988. CODEN DDJOEB. ISSN 0888-3076.

Grossman:1988:D

[Gro88b] Franklin Grossman. Debugging with the 80386. Dr. Dobb’s Journal of Software Tools, 13(2): ??, February 1988. CODEN DDJOEB. ISSN 0888-3076.

Garwood:1989:US

[GS89] Michael W. Garwood and Andrew E. Schweig. UNIX streams. Dr. Dobb’s Journal of Software Tools, 14(1):50, 53, 55, 56, 58, 60, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Glenn:1982:ILP

[GT82] George Glenn and Norman E. Thagard. An improved LET processor for LLL Basic. Dr. Dobb’s Journal of Software Tools, 7(6):26– ??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Guthery:1989:ENC

[Gut89] Scott B. Guthery. Are the Emperor’s new clothes object oriented? Dr. Dobb’s Journal of Software Tools, 14(12):80, 82–84, 86, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Halfant:1982:SC

[Hal82] Matthew Halfant. Small-C for the 9900. Dr. Dobb’s Journal of Software Tools, 7(7):66–??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Halcomb:1984:RAR

[Hal84] Jay Halcomb. Review of automation of reasoning ed. Jorg Siekmann and Graham Wilson. Dr. Dobb’s Journal of Software Tools, 9(2):82–??, February 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Hamlin:1982:Pa

[Ham82a] Chris Hamlin. PDP-1802. Dr. Dobb’s Journal of Software Tools, 7(6):??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Hamlin:1982:Pb

[Ham82b] Chris Hamlin. PDP-1802. Dr. Dobb’s Journal of Software Tools, 7(7):??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Hammond:1983:CMF

[Ham83] Nicholas Hammond. CP/M file directory expansion of 8in SS-SD in DDC. Dr. Dobb’s Journal of Software Tools, 8(6):9–??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Ham:1986:FFS

[Ham86a] Michael Ham. Factoring in Forth in SP. Dr. Dobb’s Journal of Software Tools, 11(10):104–??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Ham:1986:FPS

[Ham86b] Michael Ham. Forth: Philosophy standards and practical advice in SP. Dr. Dobb’s Journal of Software Tools, 11(7):112–??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Ham:1986:SPa

[Ham86c] Michael Ham. Structured programming. Dr. Dobb’s Journal of Software Tools, 11(4):10, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Ham:1986:SPb

[Ham86d] Michael Ham. Structured programming. Dr. Dobb’s Journal of Software Tools, 11(7):??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Hamilton:1986:FIP

[Ham86e] Dennis E. Hamilton. Fast integer powers for Pascal. Dr. Dobb’s Journal of Software Tools, 11(2): 36–??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Ham:1987:NNF

[Ham87a] Michael Ham. Naming names [in Forth] in SP. Dr. Dobb’s Journal of Software Tools, 12(1):110–??, January 1987. CODEN DDJOEB. ISSN 1044-789X.

Ham:1987:NIF

[Ham87b] Michael Ham. Numeric input in Forth in SP. Dr. Dobb’s Journal of Software Tools, 12(4):140– ??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Ham:1987:SPa

[Ham87c] Michael Ham. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(1):??, January 1987. CODEN DDJOEB. ISSN 1044-789X.

Ham:1987:SPb

[Ham87d] Michael Ham. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(4):??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Ham:1987:SPc

[Ham87e] Michael Ham. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(7):??, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Ham:1987:SPS

[Ham87f] Michael Ham. Structured programming: Software design rules. Dr. Dobb’s Journal of Software Tools, 12(7):112–116, July 1987. CODEN DDJOEB. ISSN 0888-3076.

Harrison:1983:DEM

[Har83a] Hank Harrison. Dobb’s ex machina. Dr. Dobb’s Journal of Software Tools, 8(2):??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Harrison:1983:E

[Har83b] Hank Harrison. Editorial. Dr.

Dobb’s Journal of Software Tools, 8(1):??, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Harrison:1983:SAC

[Har83c] Hank Harrison. Stone age computers — 6 000 years of computing science. Dr. Dobb’s Journal of Software Tools, 8(1):36–??, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Harry:1983:DP

[Har83d] David Harry. A DML parser. Dr.

Dobb’s Journal of Software Tools, 8(11):46–??, November 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1986:IC

[Har86] Hal W. Hardenbergh. Inefficient C. Dr. Dobb’s Journal of Software Tools, 11(1):16–??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Harrington:1987:MBA

[Har87a] Jan L. Harrington. Macintosh buttons and Amiga gadgets. Dr. Dobb’s Journal of Software Tools, 12(1):40–42, 44, 47–49, January 1987. CODEN DDJOEB. ISSN 0888-3076.

Harrington:1987:UBU

[Har87b] Jan L. Harrington. Unix BBS using shell scripts. Dr. Dobb’s Journal of Software Tools, 12(6):54–??, June 1987. CODEN DDJOEB. ISSN 0888-3076.

Harris:1987:XCR

[Har87c] Bennette R. Harris. The xor chain revisited. Dr. Dobb’s Journal of Software Tools, 12(9):36–40, 62, 65–66, September 1987. CODEN DDJOEB. ISSN 0888-3076.

Harrison:19xx:R

[Harxx] Hank Harrison. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Head:1982:CMEa

[Hea82a] Gene Head. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 7(7):??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Head:1982:CMEb

[Hea82b] Gene Head. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 7(10):??, October 1982. CODEN DDJOEB. ISSN 1044-789X.

Head:1982:CMEc

[Hea82c] Gene Head. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 7(11):??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Head:1982:NUC

[Hea82d] Gene Head. New USER for CP/M. Dr. Dobb’s Journal of Software Tools, 7(5):42–??, May 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Head:1983:CMEa

[Hea83a] Gene Head. CP/M exchange. Dr.

Dobb’s Journal of Software Tools, 8(1):??, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Head:1983:CMEb

[Hea83b] Gene Head. CP/M exchange. Dr.

Dobb’s Journal of Software Tools, 8(2):??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Head:1983:CMEc

[Hea83c] Gene Head. CP/M exchange. Dr.

Dobb’s Journal of Software Tools, 8(3):??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Head:1983:CMEd

[Hea83d] Gene Head. CP/M exchange.

Dr. Dobb’s Journal of Software Tools, 8(4):??, April 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Head:1984:DIP

[Hea84a] Gene Head. dBASE II programming techniques. Dr. Dobb’s Journal of Software Tools, 9(6):28–31, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Head:1984:MDI

[Hea84b] Gene Head. More dBASE II programming techniques. Dr. Dobb’s Journal of Software Tools, 9(10): 26–28, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Head:1984:SQD

[Hea84c] Gene Head. Solutions to quirks in dBASE II. Dr. Dobb’s Journal of Software Tools, 9(5):98–??, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Head:1985:FMa

[Hea85a] A. K. Head. Filter for MS-DOS [TEE] in 16BST. Dr. Dobb’s Journal of Software Tools, 10(4):102– ??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Head:1985:FMb

[Hea85b] A. K. Head. Filter for MS-DOS [TRIM] in 16BST. Dr. Dobb’s Journal of Software Tools, 10(6):95–??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Head:1985:MDT

[Hea85c] Gene Head. More dBASE tips and techniques. Dr. Dobb’s Journal of Software Tools, 10(2):71– ??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Head:19xx:R

[Heaxx] Gene Head. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Heckel:1985:ZDN

[Hec85] Paul Heckel. Zoomracks: Designing a new software metaphor in TSD. Dr. Dobb’s Journal of Software Tools, 10(11):48–??, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Hejlsberg:1989:COT

[Hej89] Anders Hejlsberg. Container object types in Turbo Pascal. Dr. Dobb’s Journal of Software Tools, 14(11):56, 58, 60, 62, 64, 108– 112, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Heller:1988:DCM

[Hel88] Steve Heller. A double cross for MASM. Dr. Dobb’s Journal of Software Tools, 13(10):46–48, 50, 52, 87–88, 90, 92, 96, 99, 148, October 1988. CODEN DDJOEB. ISSN 0888-3076.

Heller:1989:EH

[Hel89] Steve Heller. Extensible hashing.

Dr. Dobb’s Journal of Software Tools, 14(11):66, 69–70, 116, 118, 120–121, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Hendrix:1982:SCC

[Hen82a] J. E. Hendrix. Small-C compiler v.2. Dr. Dobb’s Journal of Software Tools, 7(12):16–??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Hendrix:1982:SSP

[Hen82b] J. E. Hendrix. Small shell: Part 2 of a north* VOS. Dr. Dobb’s Journal of Software Tools, 7(1):27–??, January 1982. CODEN DDJOEB. ISSN 1044-789X.

Hendrix:1985:SCU

[Hen85] J. E. Hendrix. Small-C update.

Dr. Dobb’s Journal of Software Tools, 10(8):84–??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Henderson-Gragg:1982:RHS

[HG82] Richard Henderson-Gragg. Review of Hayes Stack Smartmodem. Dr. Dobb’s Journal of Software Tools, 7(5):58–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Hildebrand:1988:MPO

[Hil88]  Daniel Hildebrand.  Message-

passing operating systems. Dr. Dobb’s Journal of Software Tools, 13(6):34–??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Hinson:1984:RMM

[Hin84] Elbert F. Hinson. Review of micro-SUB:MATH from Foehn Consulting. Dr. Dobb’s Journal of Software Tools, 9(4):76–??, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Hinckley:1989:OWS

[Hin89] Kee Hinckley. The OSF windowing system. Dr. Dobb’s Journal of Software Tools, 14(3):78–80, March 1989. CODEN DDJOEB. ISSN 0888-3076.

Hogan:1985:CCW

[Hog85a] Thom Hogan. Cursor control in word processors in TSD. Dr. Dobb’s Journal of Software Tools, 10(4):86–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Hogan:1985:UDV

[Hog85b] Thom Hogan. Using decision variables in graphics primitives. Dr. Dobb’s Journal of Software Tools, 10(5):40–42, 44–46, 48, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Holub:1984:GCU

[Hol84] Allen I. Holub. Grep C — a Unix-Like, generalized, regular expression parser in C. Dr. Dobb’s Journal of Software Tools, 9(10):50–??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCa

[Hol85a] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(3):??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCb

[Hol85b] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(4):??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCc

[Hol85c] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(7):??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCd

[Hol85d] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(8):??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCe

[Hol85e] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(9):??, September 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCf

[Hol85f] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(10):??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Holub:1985:CCg

[Hol85g] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 10(12):??, December 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCa

[Hol86a] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(1):??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCb

[Hol86b] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(2):??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCc

[Hol86c] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(3):??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCd

[Hol86d] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(4):??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCe

[Hol86e] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(5):??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCf

[Hol86f] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(6):??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCg

[Hol86g] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 11(7):??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCh

[Hol86h] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 11(8):??, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCi

[Hol86i] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 11(9):??, September 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCj

[Hol86j] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 11(10):??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:CCk

[Hol86k] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 11(11):??, November 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1986:EPV

[Hol86l] Allen I. Holub. Ethics in programming in VP. Dr. Dobb’s Journal of Software Tools, 11(12):14–??, December 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1986:DCV

[Hol86m] Allen I. Holub. In defense of C in VP. Dr. Dobb’s Journal of Software Tools, 11(8):16–??, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:NCI

[Hol87a] Allen Holub. NR: a C implementation of Nroff. Dr. Dobb’s Journal of Software Tools, 12(3):48–??, March 1987. CODEN DDJOEB. ISSN 0888-3076.

Holub:1987:NHE

[Hol87b] Allen Holub. Nroff: Hashing, expressions, and roman numerals. Dr. Dobb’s Journal of Software Tools, 12(2):90–??, February 1987. CODEN DDJOEB. ISSN 0888-3076.

Holub:1987:PQ

[Hol87c] Allen Holub. Priority queues. Dr. Dobb’s Journal of Software Tools, 12(6):92–??, June 1987. CODEN DDJOEB. ISSN 0888-3076.

Holub:1987:SAD

[Hol87d] Allen Holub. Statistical applications of digital low-pass filters, exec bug in Microsoft C. Dr. Dobb’s Journal of Software Tools, 12(5):102–??, May 1987. CODEN DDJOEB. ISSN 0888-3076.

Holub:1987:CCa

[Hol87e] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 12(1):??, January 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCb

[Hol87f] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 12(2):??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCc

[Hol87g] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 12(3):??, March 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCd

[Hol87h] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 12(4):??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCe

[Hol87i] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(5):??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCf

[Hol87j] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(6):??, June 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCg

[Hol87k] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(7):??, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCh

[Hol87l] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(8):??, August 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCi

[Hol87m] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(9):??, September 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCj

[Hol87n] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(10):??, October 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCk

[Hol87o] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 12(11):??, November 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1987:CCl

[Hol87p] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 12(12):??, December 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1987:CUW

[Hol87q] Allen I. Holub. Curses: UNIXcompatible windowing output functions. Dr. Dobb’s Journal of Software Tools, 12(7):94–104, 74– 93, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:EPV

[Hol87r] Allen I. Holub. Education and programming in VP. Dr. Dobb’s Journal of Software Tools, 12(4):14–??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:RL

[Hol87s] Allen I. Holub. Running light. Dr. Dobb’s Journal of Software Tools, 12(5):??, May 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Holub:1988:CCa

[Hol88a] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 13(1):??, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Holub:1988:CCb

[Hol88b] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 13(2):??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Holub:1988:CCc

[Hol88c] Allen I. Holub. C Chest. Dr. Dobb’s Journal of Software Tools, 13(4):??, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Holub:1988:CCd

[Hol88d] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 13(5):??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Holub:1988:CCe

[Hol88e] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 13(6):??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Holub:1988:CCf

[Hol88f] Allen I. Holub. C Chest. Dr.

Dobb’s Journal of Software Tools, 13(7):??, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Horn:1985:RNV

[Hor85] Bruce Horn. Review of Neon version 1.0 from Kriya Systems Inc. Dr. Dobb’s Journal of Software Tools, 10(10):96–??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Howard:1983:ECS

[How83] Alan D. Howard. Enhancing the C screen editor. Dr. Dobb’s Journal of Software Tools, 8(5):38–??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Howell:1987:AS

[How87] James Howell. An alternative to Soundex. Dr. Dobb’s Journal of Software Tools, 12(11):62– ??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Hendrix:1984:NLS

[HP84] James E. Hendrix and Ernest Payne. A new library for Small-C. Dr. Dobb’s Journal of Software Tools, 9(5):50–??, May 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Huang:1984:FCF

[Hua84] Timothy Huang. First Chinese Forth — A double-headed approach. Dr. Dobb’s Journal of Software Tools, 9(6):32–34, 36–39, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Hughes:1982:SRF

[Hug82] Lawrence Hughes. Send/receive — file transfer utilities (via RS-232 link). Dr. Dobb’s Journal of Software Tools, 7(8):41–??, August 1982. CODEN DDJOEB. ISSN 1044-789X.

Hummel:1984:GBL

[Hum84] Gwen Hummel. GENERATE.BAS — linking machine language routines to IBM BASIC. Dr. Dobb’s Journal of Software Tools, 9(1): 83–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Hunt:1983:JPA

[Hun83] Daniel S. Hunt. JRT Pascal: Another look. Dr. Dobb’s Journal of Software Tools, 8(1):44–??, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Hutchison:19xx:X0791

[Hutxx] William G. Hutchison, Jr. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Hymowech:1988:FFF

[Hym88a] M. Hymowech. Find that function! (function finder utilities).

Dr. Dobb’s Journal of Software Tools, 13(8):30–31, 33–38, 70–78, August 1988. CODEN DDJOEB. ISSN 0888-3076.

Hymowech:1988:FFI

[Hym88b] Marvin Hymowech. Find that function — from inside Brief! Dr. Dobb’s Journal of Software Tools, 13(12):70–72, December 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Ilg:1989:TEN

[IB89] Gregory R. S. Ilg and R. J. Brown.

A timed event network scheduler in Forth. Dr. Dobb’s Journal of Software Tools, 14(2):52, 54, 56, 58, 92–94, 96, February 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Ingerman:1982:MMY

[Ing82] Peter Zilahy Ingerman. Make the most or your check digits. Dr. Dobb’s Journal of Software Tools, 7(4):36–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Irving:1983:FSB

[Irv83] Robert Irving. Faster smaller BA-SIC. A common-sense guide to. Dr. Dobb’s Journal of Software Tools, 8(3):66–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Irving:19xx:R

[Irvxx] Robert Irving. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

James:1982:PMF

[Jam82] John S. James. Program for a modem: Forth code from the conference tree. Dr. Dobb’s Journal of Software Tools, 7(5):18–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

James:1989:U

[Jam89] Rahner James. Undocumented DOS. Dr. Dobb’s Journal of Software Tools, 14(6):26–28, 30, 34, 84, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Jennings:1985:DTT

[Jen85] Earle Jennings. Diagnostic tools in Turbo Pascal [letter]. Dr. Dobb’s Journal of Software Tools, 10(2): 10–??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Johansson:1982:SCA

[Joh82] Jan-Henrik Johansson. Small-C argc and argv for under CP/M in DDC. Dr. Dobb’s Journal of Software Tools, 7(12):62–??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1984:FBP

[Joh84] John R. Johnson. A file browser program. Dr. Dobb’s Journal of Software Tools, 9(12):60–??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1989:PDP

[Joh89] Margaret Johnson. The portability dream (portable code between windowing environments). Dr. Dobb’s Journal of Software Tools, 14(3):70–72, 74–75, 77, 102–104, March 1989. CODEN DDJOEB. ISSN 0888-3076.

Johnson:19xx:R

[Johxx] John R. Johnson. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:19xx:Ra

[Jolxxa] Lynne Greer Jolitz. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:19xx:Rb

[Jolxxb] William Frederick Jolitz. Reviews.

Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CO-DEN DDJOEB. ISSN 1044-789X.

Jolly:19xx:R

[Jolxxc] George W. Jolly. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Jones:1983:YYC

[Jon83] Do-While Jones. Yes you can trace through BDOS. Dr. Dobb’s Journal of Software Tools, 8(6):60–??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Jones:1986:LAM

[Jon86] Do-While Jones. Learning Ada on a micro. Dr. Dobb’s Journal of Software Tools, 11(2):42– ??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Jones:1987:DDT

[Jon87] Do-While Jones. Dimensional data types. Dr. Dobb’s Journal of Software Tools, 12(5):50–??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Jones:1988:DTR

[Jon88] R. L. Jones. Do they really pay for every line? (parallel programming). Dr. Dobb’s Journal of Software Tools, 13(8):130, 132, 134– 135, 138, 140, August 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Johnson:1989:DLL

[JS89] Margaret Johnson and Mark Solinski. Dynamic link libraries under Microsoft Windows. Dr. Dobb’s Journal of Software Tools, 14(3): 28, 29, 32, 35, 37, 82, 84, 86–91, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Kachigan:19xx:R

[Kacxx] Karl R. Kachigan. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Kan:1982:ASC

[Kan82a] Michael K. Kan. Add a system clock. Dr. Dobb’s Journal of Software Tools, 7(5):30–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Kan:1982:PES

[Kan82b] Michael K. Kan. Performance evaluation in small computers. Dr. Dobb’s Journal of Software Tools, 7(8):34–??, August 1982. CODEN DDJOEB. ISSN 1044-789X.

Kaplon:1984:DCA

[Kap84] Morton F. Kaplon. DO-COM.ASM — load and execute COM programs in 16BST. Dr. Dobb’s Journal of Software Tools, 9(4):88–??, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Kaplon:19xx:R

[Kapxx] Morton F. Kaplon. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Kar:1989:DM

[Kar89] Rabindra P. Kar. Data-flow multitasking. Dr. Dobb’s Journal of Software Tools, 14(11):16–18, 20, 22, 24, 84, November 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Katz:1986:MPM

[Kat86] Howard Katz. A Mandelbrot program for the Macintosh. Dr. Dobb’s Journal of Software Tools, 11(11):42–45, November 1986. CODEN DDJOEB. ISSN 0888-3076.

Kawabata:1984:SFP

[Kaw84] Hugh M. Kawabata. Savage floating point benchmark in Fortran in 16BST. Dr. Dobb’s Journal of Software Tools, 9(1):83–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Kienle:1989:NGO

[Kie89] Steven Kienle. Network graphs in object Pascal. Dr. Dobb’s Journal of Software Tools, 14(12):17– 18, 20, 22, 24–25, 104–106, 108, 111, December 1989. CODEN DDJOEB. ISSN 1044-789X.

King:1983:JDM

[Kin83] Gordon King. Julian dates for microcomputers. Dr. Dobb’s Journal of Software Tools, 8(6):66–??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

King:1985:RSP

[Kin85] Stephen King. Review of SAVVY PC version 4.0 from Excalibur technologies corp. Dr. Dobb’s Journal of Software Tools, 10(3): 116–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

King:1988:APP

[Kin88a] K. N. King. Ada for Pascal programmers. Dr. Dobb’s Journal of Software Tools, 13(9):34– 51, September 1988. CODEN DDJOEB. ISSN 0888-3076.

King:1988:DRS

[Kin88b] T. King. Dynamic run-time structures. Dr. Dobb’s Journal of Software Tools, 13(11):50–54, 92– 93, November 1988. CODEN DDJOEB. ISSN 0888-3076.

King:1989:FS

[Kin89a] Todd King. Force-based simulations. Dr. Dobb’s Journal of Software Tools, 14(9):40–42, 94– 96, September 1989. CODEN DDJOEB. ISSN 0888-3076.

King:1989:UNN

[Kin89b] Todd King. Using neural networks for pattern recognition. Dr. Dobb’s Journal of Software Tools, 14(1): 14–28, 90–95, January 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Klein:1987:WMD

[Kle87] Andy Klein. Writing MS-DOS device drivers in C. Dr. Dobb’s Journal of Software Tools, 12(9): 44–??, September 1987. CODEN DDJOEB. ISSN 0888-3076.

Klimasauskas:1989:NNN

[Kli89] Casimir C. Klimasauskas. Neural nets and noise filtering. Dr. Dobb’s Journal of Software Tools, 14(1):

32–48, 96–100, January 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Kokkonen:1989:MME

[Kok89] Kim Kokkonen. More memory for DOS Exec. Dr. Dobb’s Journal of Software Tools, 14(4):14–15, 18, 22, 66, 68–71, April 1989. CODEN DDJOEB. ISSN 0888-3076.

Kossow:1982:CAa

[Kos82a] Al Kossow. 68000 cross assembler.

Dr. Dobb’s Journal of Software Tools, 7(10):??, October 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Kossow:1982:CAb

[Kos82b] Al Kossow. 68000 cross assembler. Dr. Dobb’s Journal of Software Tools, 7(11):??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Kossow:1982:MPS

[Kos82c] Al Kossow. Multi-68000s in a personal system. Dr. Dobb’s Journal of Software Tools, 7(6):22–??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Kar:1989:RRB

[KP89] Rabindra P. Kar and Kent Porter.

RHEALSTONE a real-time benchmarking proposal. Dr. Dobb’s Journal of Software Tools, 14(2): 14–16, 18, 22, 24, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Kalinsky:1988:ESD

[KR88] David Kalinsky and James Ready.

Embedded systems design — A special CASE? Dr. Dobb’s Journal of Software Tools, 13(??):37– ??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Krantz:1984:R

[Kra84] Donald G. Krantz. RESORT. Dr. Dobb’s Journal of Software Tools, 9(11):88–93, November 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Krantz:1985:CPC

[Kra85] Donald G. Krantz. Christensen protocols in C. Dr. Dobb’s Journal of Software Tools, 10(6):66–??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Krantz:1987:WD

[Kra87] Donald G. Krantz. What’s the Diff? Dr. Dobb’s Journal of Software Tools, 12(8):30–??, August 1987. CODEN DDJOEB. ISSN 1044-789X.

Kronman:19xx:R

[Kroxx] James Kronman. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Krute:1988:Ma

[Kru88a] Stan Krute. To the Macs. Dr. Dobb’s Journal of Software Tools, 13(1):??, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Krute:1988:Mb

[Kru88b] Stan Krute. To the Macs. Dr. Dobb’s Journal of Software Tools, 13(4):??, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Krute:1988:Mc

[Kru88c] Stan Krute. To the Macs. Dr. Dobb’s Journal of Software Tools, 13(5):??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Krute:1988:XXH

[Kru88d] Stan Krute. XCMD and XFCN:

HyperCard’s software slots. Dr. Dobb’s Journal of Software Tools, 13(9):90–91, 108–109, September 1988. CODEN DDJOEB. ISSN 0888-3076.

Kvaleberg:1982:SCD

[Kva82] Egil Kvaleberg. Small-C DISK-DOC: A repair and maintenance utility. Dr. Dobb’s Journal of Software Tools, 7(4):26–??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Kindel:19xx:COM

[KWxx] Charles Kindel and Sara Williams.

The component object model. Dr. Dobb’s Journal of Software Tools, ??(??):14–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Lacey:1983:HCB

[Lac83] David S. Lacey. Highfalutin’ Computin’ with Bob Orrfelt On Your Timex Sinclair 1000 Computer, by Bob Orrfelt. Dr. Dobb’s Journal of Software Tools, 8(5):72–??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Ladd:1988:SMD

[Lad88] Scott Robert Ladd. Spelunking MS-DOS: documenting the undocumented. Dr. Dobb’s Journal of Software Tools, 13(12):56–60, December 1988. CODEN DDJOEB. ISSN 0888-3076.

Ladd:1989:BTC

[Lad89a] Scott Robert Ladd. Benchmarking Turbo C and quick C. Dr. Dobb’s Journal of Software Tools, 14(8):89–??, August 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Ladd:1989:CSC

[Lad89b] Scott Robert Ladd. C++ string classes. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):18–21, 68–70, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Ladd:1989:CMC

[Lad89c] Scott Robert Ladd. Comparing Modula-2 and C++. Dr. Dobb’s Journal of Software Tools, 14(1): 62–68, 102–104, January 1989. CODEN DDJOEB. ISSN 0888-3076.

Ladd:1989:GKA

[Lad89d] Scott Robert Ladd. Going from K and R to ANSI C. Dr. Dobb’s Journal of Software Tools, 14(8): 74, 76–80, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Ladd:1989:GE

[Lad89e] Scott Robert Ladd. Guest editorial. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):??, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Ladd:1989:QCV

[Lad89f] Scott Robert Ladd. Quick C versus Turbo C. Dr. Dobb’s Journal of Software Tools, 14(5):68–73, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Lane:1989:GPP

[Lan89] Alex Lane. Generating parsers with PCYACC. Dr. Dobb’s Journal of Software Tools, 14(6):76–77, 79, 81, 110–112, June 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Lane:19xx:R

[Lanxxa] Alex Lane. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Langevin:19xx:R

[Lanxxb] R. A. Langevin. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Lazarev:1988:PPS

[Laz88] G. L. Lazarev. Prolog/V: Prolog in the Smalltalk environment. Dr. Dobb’s Journal of Software Tools, 13(11):68–80, 98– 102, November 1988. CODEN DDJOEB. ISSN 0888-3076. Describes briefly (mainly by examples) a commercial Prolog system developed and integrated in Smalltalk/V.

Lazarev:1989:ESP

[Laz89] Gregory L. Lazarev. Executable specifications with prolog. Dr. Dobb’s Journal of Software Tools, 14(10):61–62, 64, 66–68, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Leblang:1988:AWT

[Leb88] David B. Leblang. Applying workstation technology to CASE. Dr. Dobb’s Journal of Software Tools, 13(??):50–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Lee:1983:FCR

[Lee83] Daniel L. Lee. A fast circle routine. Dr. Dobb’s Journal of Software Tools, 8(5):32, 35–37, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Levitt:1987:PSE

[Lev87] David Levitt. Pushing the sound envelope. Dr. Dobb’s Journal of Software Tools, 12(5):16–19, May 1987. CODEN DDJOEB. ISSN 0888-3076.

Lewis:1988:SSP

[Lew88] Ted G. Lewis. Software and the single programmer. Dr. Dobb’s Journal of Software Tools, 13(??): 18–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1986:FWI

[Lin86] Craig A. Lindley. Forth Windows for the IBM PC. Dr. Dobb’s Journal of Software Tools, 11(7):46– ??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1987:MTP

[Lin87] Craig A. Lindley. Multitasking with Turbo Pascal. Dr. Dobb’s Journal of Software Tools, 12 (7):42–??, July 1987. CODEN DDJOEB. ISSN 0888-3076.

Lindley:1988:SPA

[Lin88] Craig A. Lindley. A serial protocol analyzer program. Dr. Dobb’s Journal of Software Tools, 13(2): 30–??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Louis:1984:RSW

[Lou84] Gregory Louis. Review of SMAL/ 80 from Westico. Dr. Dobb’s Journal of Software Tools, 9(1):98–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Lozes:19xx:R

[Lozxx] Richard L. Lozes. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Lafleur:1985:FYM

[LR85] Thomas Lafleur and Susan Raab.

Fatten your Mac. Dr. Dobb’s Journal of Software Tools, 10(1):18–??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Lunsford:1983:RVF

[Lun83] Daniel R. Lunsford. Review of VERSA-FILE from Analytic Management Systems. Dr. Dobb’s Journal of Software Tools, 8(11): 97–??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Lyons:1983:RNR

[Lyo83] Virginia Lyons. Review of The Network Revolution: Confessions of a Computer Scientist, by Jacque Vallee. Dr. Dobb’s Journal of Software Tools, 8(2):65–??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Macpherson:1983:SCC

[Mac83] Andrew Macpherson. Small-C compiler v.2 bug fix [letter]. Dr. Dobb’s Journal of Software Tools, 8(2):10–??, February 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Mak:1989:SUP

[Mak89] Nico Mak. SWAP (utility program). Dr. Dobb’s Journal of Software Tools, 14(4):44–48, 72– 75, 77–78, 80, April 1989. CODEN DDJOEB. ISSN 0888-3076.

Malpas:1985:PL

[Mal85] John Malpas. Programming in logic. Dr. Dobb’s Journal of Software Tools, 10(3):36–38, 40–41, March 1985. CODEN DDJOEB. ISSN 1044-789X. TK 7885 A1 D6.

Manning:1983:FS

[Man83] Mark Manning. The Forth sort. Dr. Dobb’s Journal of Software Tools, 8(9):102–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Marasco:1986:HFL

[Mar86a] Joseph Marasco. How to fix line glitches. Dr. Dobb’s Journal of Software Tools, 11(6):32–??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Marasco:1986:VMM

[Mar86b] Joseph Marasco. A variable metric minimizer. Dr. Dobb’s Journal of Software Tools, 11(3):24–??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Marasco:1987:RNR

[Mar87a] Joseph Marasco. Review of Numerical Recipes: The Art of Scientific Computing, by Press et al. Dr. Dobb’s Journal of Software Tools, 12(5):146–??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Mariella:1987:IDF

[Mar87b] Ray Mariella. Integers don’t float. Dr. Dobb’s Journal of Software Tools, 12(12):48–??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Margulis:1988:PMI

[Mar88] Neil Margulis. 80386 protected mode initialization. Dr. Dobb’s Journal of Software Tools, 13(10): 36–37, 39, 84, 86, October 1988. CODEN DDJOEB. ISSN 0888-3076.

Margulis:1989:AMM

[Mar89] Neil Margulis. Advanced 80386 memory management. Dr. Dobb’s Journal of Software Tools, 14(4): 24, 28–30, April 1989. CODEN DDJOEB. ISSN 0888-3076.

Mathews:1988:TBT

[Mat88] James Mathews. Threaded binary trees. Dr. Dobb’s Journal of Software Tools, 13(3):42–??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Maurer:1982:IUD

[Mau82a] W. D. Maurer. Improvement upon a division program by Leventhal. Dr. Dobb’s Journal of Software Tools, 7(3):20–??, March 1982. CODEN DDJOEB. ISSN 1044-789X.

Maurer:1982:SM

[Mau82b] W. D. Maurer. Simplified 68000 mnemonics. Dr. Dobb’s Journal of Software Tools, 7(10):36–??, October 1982. CODEN DDJOEB. ISSN 1044-789X.

May:1987:DIC

[May87] William D. May. 3-D images from contour maps. Dr. Dobb’s Journal of Software Tools, 12(11): 18–??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

May:1988:SD

[May88] William D. May. A simple decompiler. Dr. Dobb’s Journal of Software Tools, 13(6):50–??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Mayer:19xx:QMI

[Mayxx] Thom Mayer. QuickDraw meets ImageWriter. Dr. Dobb’s Journal of Software Tools, ??(1):26–??, January 19xx. CODEN DDJOEB. ISSN 1044-789X.

McCabe:1983:SFV

[McC83] C. Kevin McCabe. Say “Forth Votrax driver”. Dr. Dobb’s Journal of Software Tools, 8(4):55–??, April 1983. CODEN DDJOEB. ISSN 1044-789X.

McDermott:1984:OSC

[McD84] Edward McDermott. Optimizing strings in C. Dr. Dobb’s Journal of Software Tools, 9(4):18–20, 22– 23, April 1984. CODEN DDJOEB. ISSN 1044-789X.

McIvor:1986:GFO

[McI86] Robert A. McIvor. A general first-order sorting algorithm. Dr. Dobb’s Journal of Software Tools, 11(9):32–??, September 1986. CO-DEN DDJOEB. ISSN 1044-789X.

McKeon:1983:SCO

[McK83] Brian McKeon. A Small-C operating system. Dr. Dobb’s Journal of Software Tools, 8(3):36–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

McKillip:1984:RUR

[McK84] Michael McKillip. Review of U.S.

Robotics’ S-100 modem. Dr. Dobb’s Journal of Software Tools, 9(2):70, 72–74, February 1984. CODEN DDJOEB. ISSN 1044-789X.

McLaughlin:1989:SA

[McL89] Michael P. McLaughlin. Simulated annealing. Dr. Dobb’s Journal of Software Tools, 14(9):26–37, 88–93, September 1989. CODEN DDJOEB. ISSN 0888-3076.

McMahon:1987:DMO

[McM87] Steve McMahon. Dynamic memory overlays for Turbo Pascal. Dr. Dobb’s Journal of Software Tools, 12(6):50–52, 78–79, June 1987. CODEN DDJOEB. ISSN 0888-3076.

McManis:1989:MGC

[McM89] Charles McManis. Multitasking OS and graphics coprocessors. Dr. Dobb’s Journal of Software Tools, 14(7):36–42, 90–91, July 1989. CODEN DDJOEB. ISSN 0888-3076.

McNierney:1986:NIP

[McN86] Ed McNierney. New issues in PC graphics. Dr. Dobb’s Journal of Software Tools, 11(11):30–32, 35– 36, 38, November 1986. CODEN DDJOEB. ISSN 0888-3076.

McWorter:1983:FCP

[McW83a] William A. McWorter, Jr. Fast circle plotter. Dr. Dobb’s Journal of Software Tools, 8(12):18–??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

McWorter:1983:MAT

[McW83b] William A. McWorter, Jr. Mc-WORDER — A tiny text editor. Dr. Dobb’s Journal of Software Tools, 8(8):36, 38–40, 42, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Meadow:1988:AUU

[Mea88a] Anthony Meadow. Apple’s A/UX: UNIX for the rest of us? Dr. Dobb’s Journal of Software Tools, 13(??):44–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Meadow:1988:MPR

[Mea88b] Anthony Meadow. A Mac programmer’s resource kit. Dr. Dobb’s Journal of Software Tools, 13(??): 8–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Melnikof:1989:NNS

[Mel89] Steve Melnikof. Neural networks for signal processing: A case study. Dr. Dobb’s Journal of Software Tools, 14(1):36–??, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Menico:1989:DTP

[Men89a] Costas Menico. Debugging TSR programs. Dr. Dobb’s Journal of Software Tools, 14(2):67–68, 70, 104–106, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Menico:1989:FSS

[Men89b] Costas Menico. Faster string searches (Boyer–Moore algorithm). Dr. Dobb’s Journal of Software Tools, 14(7):74–75, 98–99, July 1989. CODEN DDJOEB. ISSN 0888-3076.

Menico:1989:HFT

[Men89c] Costas Menico. High-speed file transfers with NetBIOS. Dr. Dobb’s Journal of Software Tools, 14(10):38, 40, 42–43, 88–91, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Merilatt:1989:CDM

[Mer89] Randall Merilatt. C dynamic memory use. Dr. Dobb’s Journal of Software Tools, 14(8):62, 64, 66– 67, 125–127, August 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Mesirov:1984:SRS

[Mes84] Richard Mesirov. Subroutine to read stock history files. Dr. Dobb’s Journal of Software Tools, 9(3): 10–??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Meyer:1983:SCA

[Mey83] Michael Meyer. Some comments on Augusta short circuits [letter]. Dr. Dobb’s Journal of Software Tools, 8(3):6–??, March 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Meyer:1989:WCS

[Mey89] Bertrand Meyer. Writing correct software. Dr. Dobb’s Journal of Software Tools, 14(12):48, 52, 54, 56, 58, 60, 63, 125, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Michalski:1982:SVG

[Mic82] Marek W. Michalski. A simple vector generation algorithm. Dr. Dobb’s Journal of Software Tools, 7(12):58–??, December 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Michener:1988:TSK

[Mic88] John R. Michener. A tool for secret key cryptography. Dr. Dobb’s Journal of Software Tools, 13(8): 50–52, 55, 96, August 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Mills:1982:CAP

[Mil82] Jonathan Mills. 8080/8086 cross assembler part II. Dr. Dobb’s Journal of Software Tools, 7(4):16– ??, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Miller:1984:IIP

[Mil84] Craig Miller. Implementing OUT: and INP: in PIP.COM in CPME. Dr. Dobb’s Journal of Software Tools, 9(5):18–??, May 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Miller:1987:AIH

[Mil87] Ron Miller. Automated interrupt handling in C. Dr. Dobb’s Journal of Software Tools, 12(11): 54–??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Mischel:1989:GVD

[Mis89a] James Mischel. A global variable device driver for MS-DOS. Dr. Dobb’s Journal of Software Tools, 14(10):70, 73, 104–111, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Mischel:1989:WAE

[Mis89b] James Mischel. Writing AWK-like extensions to C. Dr. Dobb’s Journal of Software Tools, 14(6):64–66, 68, 70, 94, 96, 98, 103–104, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Mitchell:1983:AAP

[Mit83a] Edward Mitchell. Augusta — 2.

the Augusta P-Code interpreter. Dr. Dobb’s Journal of Software Tools, 8(3):20–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASa

[Mit83b] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(1): 14, 16–23, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASb

[Mit83c] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(3):14–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASc

[Mit83d] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(5):14– ??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASd

[Mit83e] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(7):14– ??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:APIa

[Mit83f] Edward Mitchell. Augusta, part III — recursive descent compilers. Dr. Dobb’s Journal of Software Tools, 8(5):13–17, 20–30, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:APIb

[Mit83g] Edward Mitchell. Augusta, part IV the Augusta compiler (continued). Dr. Dobb’s Journal of Software Tools, 8(7):12–??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1984:SWP

[Mit84] Edward Mitchell. A simple window package. Dr. Dobb’s Journal of Software Tools, 9(1):36–43, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1985:DCS

[Mit85a] Bridger Mitchell. A debugging console for SID ZSID and DDT in CPME. Dr. Dobb’s Journal of Software Tools, 10(6):106–??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1985:MDD

[Mit85b] Stan Mitchell. MS-DOS devicedriver chain display/dump in 16BST. Dr. Dobb’s Journal of Software Tools, 10(5):119–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:19xx:R

[Mitxx] Phil Mitchell. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

McKay:1984:CMG

[MM84a] Derek McKay and Bridger Mitchell. CP/M 2.2 goes PUBlic. Dr. Dobb’s Journal of Software Tools, 9(11):48–??, November 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Mitchell:1984:CGP

[MM84b] Bridger Mitchell and Derek McKay. CP/M 2.2 goes public. Dr. Dobb’s Journal of Software Tools, 9(11):48–??, November 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Moir:1989:MSS

[Moi89] Dale Moir. Maintaining system security. Dr. Dobb’s Journal of Software Tools, 14(6):75, 156–157, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Monroe:1982:FFP

[Mon82] Alfred J. Monroe. Forth floating point package. Dr. Dobb’s Journal of Software Tools, 7(9): 16–??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Monroe:1983:FDF

[Mon83] Alfred J. Monroe. An 8080 Fig-Forth directory and file system. Dr. Dobb’s Journal of Software Tools, 8(4):38–40, April 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Moon:1988:AAV

[Moo88a] Raymond Moon. Arguments and automatic variables in assembly language. Dr. Dobb’s Journal of Software Tools, 13(9):82–88, 110– 111, September 1988. CODEN DDJOEB. ISSN 0888-3076.

Moore:1988:MMA

[Moo88b] Robert J. Moore. Mapping DOS memory allocation. Dr. Dobb’s Journal of Software Tools, 13(11): 56–64, 94–96, November 1988. CODEN DDJOEB. ISSN 0888-3076.

Morgenstern:1983:PLC

[Mor83] L. Morgenstern. Program linkage by coroutines — Forth to Basic. Dr. Dobb’s Journal of Software Tools, 8(2):26–31, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Moran:1984:SST

[Mor84] Thomas W. Moran. Some spreadsheet tests of the Savage floatingpoint benchmark in 16BST. Dr. Dobb’s Journal of Software Tools, 9(8):106–??, August 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Morton:1986:DDB

[Mor86] Michael Morton. A digital dissolve for bit-mapped graphics screens. Dr. Dobb’s Journal of Software Tools, 11(11):48–50, 52, 54–56, November 1986. CODEN DDJOEB. ISSN 0888-3076.

Moran:19xx:R

[Morxx] James Moran. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Mott:1984:STP

[Mot84] James Mott. A simple token parsing filter for DOS 2.0. Dr. Dobb’s Journal of Software Tools, 9(11): 99–??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Murray:1989:LOP

[MP89] William H. Murray and Chris H.

Pappas. Line-of-best-fit (OS/2 Presentation Manager). Dr. Dobb’s Journal of Software Tools, 14(7):14–20, 78–82, July 1989. CODEN DDJOEB. ISSN 0888-3076.

McLarty:1984:IMP

[MS84] Hugh McLarty and David W.

Smith. An introduction to Modula-2 for Pascal programmers. Dr. Dobb’s Journal of Software Tools, 9(5):22–24, 26–27, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Murphy:1983:GLR

[Mur83a] James Murphy. Graphics low-res color for IBM PC setup. Dr. Dobb’s Journal of Software Tools, 8(3):17–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Murphy:1983:RFR

[Mur83b] Walter V. Murphy. RECLAIM — A file reclamation utility for destroyed [CP/M] directories. Dr. Dobb’s Journal of Software Tools, 8(3):82–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Musselman:1987:TBA

[Mus87] John Musselman. Two-bit analogto-digital conversion. Dr. Dobb’s Journal of Software Tools, 12 (6):22–24, June 1987. CODEN DDJOEB. ISSN 0888-3076.

Mouhanna:1989:QQ

[MV89] Joseph Mouhanna and Michael Vose. The QuickPascal in Quick-Pascal. Dr. Dobb’s Journal of Software Tools, 14(12):64–66, 68, 70, 126–128, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Naro:1987:PRC

[Nar87] Richard Naro. Putting ROM code in its place. Dr. Dobb’s Journal of Software Tools, 12(12):38–??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1985:SSP

[Nel85] Amer W. Nelson. Shortcut to SCISTAR: For Prowriter users. Dr. Dobb’s Journal of Software Tools, 10(4):24–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1986:P

[Nel86] Ross Nelson. Programming on the 80386. Dr. Dobb’s Journal of Software Tools, 11(10):28–??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1987:AWM

[Nel87] Russell Nelson. ARC wars: MS-DOS archiving utilities. Dr. Dobb’s Journal of Software Tools, 12(3):26–28, 30, March 1987. CO-DEN DDJOEB. ISSN 0888-3076.

Nelson:1989:LDC

[Nel89] Mark R. Nelson. LZW data compression. Dr. Dobb’s Journal of Software Tools, 14(10):29, 32, 34, 36, 86–87, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Nelson:19xx:R

[Nelxx] Ross Nelson. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):

??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Nevin:1989:AAA

[Nev89] Randall Nevin. Autorouting with the A* algorithm. Dr. Dobb’s Journal of Software Tools, 14(9): 16–23, 82–86, September 1989. CODEN DDJOEB. ISSN 0888-3076.

Newberry:1983:RAC

[New83] Steve Newberry. Review of AC-Txx cross assemblers from SOR-CIM. Dr. Dobb’s Journal of Software Tools, 8(8):108–??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Newberry:1984:HAF

[New84] Steve Newberry. History of ackermann’s function [letter]. Dr. Dobb’s Journal of Software Tools, 9(4):9–??, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Nicholson:1982:BIT

[Nic82] Leonard K. Nicholson. Background information on text editors. Dr. Dobb’s Journal of Software Tools, 7(2):55–??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Nutter:1988:ADC

[Nut88] Stewart Nutter. An aid to documenting C. Dr. Dobb’s Journal of Software Tools, 13(8):40–44, 48, 84–87, 90–94, August 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Odette:1982:ZF

[Ode82] Lou L. Odette. Z8000 Forth. Dr.

Dobb’s Journal of Software Tools, 7(9):48–??, September 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Odette:1983:NCW

[Ode83] Lou L. Odette. Nondeterministic control words in Forth. Dr. Dobb’s Journal of Software Tools, 8(9):44, 46, 51–??, September 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Odette:1984:CS

[Ode84] Lou L. Odette. Computing with streams. Dr. Dobb’s Journal of Software Tools, 9(9):50– ??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Ordman:1982:RPL

[Ord82] Edward T. Ordman. Review of Programming Language Translation, by R. E. Berry. Dr. Dobb’s Journal of Software Tools, 7(9): 67–??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Ordman:1984:RIP

[Ord84] Eunice B. Ordman. Review of The IBM PC Connection, by James W. Coffron. Dr. Dobb’s Journal of Software Tools, 9(10):120–??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Ouverson:1982:UCB

[Ouv82] Marlin Ouverson. Using computer bulletin boards. Dr. Dobb’s Journal of Software Tools, 7(5):50–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Ouverson:1984:RSR

[Ouv84] Marlin Ouverson. Review of SuperScripsit from radio shack for TRS-80 models I and III. Dr.

Dobb’s Journal of Software Tools, 9(1):90–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Ouverson:19xx:E

[Ouvxx] Marlin Ouverson. Editorial. Dr.

Dobb’s Journal of Software Tools, ??(??):1982–??, ???? 19xx. CO-DEN DDJOEB. ISSN 1044-789X.

Palmer:1985:MME

[Pal85] Gary B. Palmer. A magic mushroom for the Epson Alice: A practical program for printing phonetic characters. Dr. Dobb’s Journal of Software Tools, 10(4):58–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Park:1984:ESW

[Par84] Jack Park. Expert systems and the weather. Dr. Dobb’s Journal of Software Tools, 9(4):24–28, 30– 31, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Park:1986:CAW

[Par86] Jack Park. A cellular automaton written in expert-2. Dr. Dobb’s Journal of Software Tools, 11(4):42–??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Parker:1988:UAD

[Par88] Jeffrey M. Parker. Using an API as a developer platform. Dr. Dobb’s Journal of Software Tools, 13(??):28–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Parsons:19xx:R

[Parxx] Ronald G. Parsons. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Passe:1983:BAH

[Pas83a] Steve Passe. Binary to ASCII hex conversions. Dr. Dobb’s Journal of Software Tools, 8(7):94–??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Passe:1983:KM

[Pas83b] Steve Passe. Kernel for the MC68000 A. Dr. Dobb’s Journal of Software Tools, 8(11):20– ??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Staff:1982:RCV

[PCC82] PCC Staff. Review of COMM-WHIZ from VolksMicro Computer Systems Inc. Dr. Dobb’s Journal of Software Tools, 7(5):53–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Perry:1983:FA

[Per83] Michael A. Perry. 68000 Forth assembler. Dr. Dobb’s Journal of Software Tools, 8(9):28– ??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Peterson:1989:SPP

[Pet89a] Mark C. Peterson. Setting precedence (Ptree search algorithm). Dr. Dobb’s Journal of Software Tools, 14(9):44–50, 98– 101, September 1989. CODEN DDJOEB. ISSN 0888-3076.

Peterson:1989:MAC

[Pet89b] Steve Peterson. A memory allocation compaction system. Dr. Dobb’s Journal of Software Tools, 14(4):50–56, 82–90, April 1989. CODEN DDJOEB. ISSN 0888-3076.

Pfefferkorn:19xx:R

[Pfexx] Charles Pfefferkorn. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Pierson:1987:FPM

[Pie87] Dan L. Pierson. Four PRO-LOGs for the Macintosh. Dr. Dobb’s Journal of Software Tools, 12(4):30–??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Piotrowski:1984:CMC

[Pio84] Walt Piotrowski. CP/M on the Commodore 64. Dr. Dobb’s Journal of Software Tools, 9(6):14–??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Piotrowski:1985:MMC

[Pio85] Walt Piotrowski. MBOOT and MODEM7 for the C-64’s CP/M. Dr. Dobb’s Journal of Software Tools, 10(1):62–??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Pirko:1983:IBA

[Pir83] Robert Pirko. IBM BASIC: Analysis of code with tables in DDC. Dr. Dobb’s Journal of Software Tools, 8(3):82–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Pittman:1985:MHF

[Pit85] Tom Pittman. Mac the hack in Festschrift for Doctor Dobb. Dr. Dobb’s Journal of Software Tools, 10(2):28–??, February 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Poole:1989:CPV

[Poo89] Kevin E. Poole. C printer for VMS and Unix. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):44–??, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1987:GTTa

[Por87a] Kent Porter. A graphics toolbox for Turbo C. Dr. Dobb’s Journal of Software Tools, 12(11):??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Porter:1987:GTTb

[Por87b] Kent Porter. A graphics toolbox for Turbo C. Dr. Dobb’s Journal of Software Tools, 12(12):??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:IDS

[Por88a] K. Porter. Improving on DIR (structured programming). Dr. Dobb’s Journal of Software Tools, 13(8):120, 122–123, 126, 128, 142– 144, 147–148, August 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Porter:1988:SAM

[Por88b] K. Porter. State of the art in Modula-2 [review of four compilers]. Dr. Dobb’s Journal of Software Tools, 13(9):64–81, 100– 106, September 1988. CODEN DDJOEB. ISSN 0888-3076.

Porter:1988:SPk

[Por88c] K. Porter. Structured programming: huge arrays revisited. Dr. Dobb’s Journal of Software Tools, 13(10):100–102, 105–106, 136–137, October 1988. CODEN DDJOEB. ISSN 0888-3076.

Porter:1988:NG

[Por88d] Kent Porter. The Norton Guides.

Dr. Dobb’s Journal of Software Tools, 13(2):116–121, February 1988. CODEN DDJOEB. ISSN 1044-789X. Examining Room: Reviews Department.

Porter:1988:SPa

[Por88e] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(2):??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPb

[Por88f] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(3):??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPd

[Por88g] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(5):??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPe

[Por88h] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(6):??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPf

[Por88i] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(7):??, July 1988.

CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPg

[Por88j] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(8):??, August 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPh

[Por88k] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(9):??, September 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPj

[Por88l] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPl

[Por88m] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(11):??, November 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPm

[Por88n] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 13(12):??, December 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:SPI

[Por88o] Kent Porter. Structured programming: Implementing Wirth’s LineDrawing module. Dr. Dobb’s Journal of Software Tools, 13(4): 118, 121–122, 124, 127, 128, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:WIM

[Por88p] Kent Porter. What is Modula-2? Dr. Dobb’s Journal of Software Tools, 13(??):??, ?? 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPa

[Por89a] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(2):??, February 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPb

[Por89b] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(3):??, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPc

[Por89c] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(4):??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPd

[Por89d] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(5):??, May 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPe

[Por89e] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(6):??, June 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPf

[Por89f] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(7):??, July 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:GPg

[Por89g] Kent Porter. Graphics programming. Dr. Dobb’s Journal of Software Tools, 14(8):??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:1989:SP

[Por89h] Kent Porter. Structured programming. Dr. Dobb’s Journal of Software Tools, 14(1):??, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Porter:19xx:R

[Porxx] Kent Porter. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Pottebaum:1989:CTPb

[Pot89a] K. L. Pottebaum. Creating TSR programs. II. Dr. Dobb’s Journal of Software Tools, 14(6):72–74, 106–107, 109, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Pottebaum:1989:CTPa

[Pot89b] Ken L. Pottebaum. Creating TSR programs with Turbo Pascal. I. Dr. Dobb’s Journal of Software Tools, 14(5):14–16, 20, 74, 76, 81– 82, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Parsons:1985:LMA

[PP85] Ronald G. Parsons and Vincent Parsons. Let’s mouse around: A Turbo Pascal/Microsoft mouse sketching program for IBM PCs. Dr. Dobb’s Journal of Software Tools, 10(4):74–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Prince:1984:SFP

[Pri84] Thomas Prince. Savage floating point benchmark in logo in 16BST. Dr. Dobb’s Journal of Software Tools, 9(6):82–??, June 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Provenzano:1984:RIP

[Pro84] Thomas Provenzano. Review of Interactive Programming Environments, by Barstow, Shrobe, and Sandewall. Dr. Dobb’s Journal of Software Tools, 9(11):122– ??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Pulier:1983:FCA

[Pul83] Myron L. Pulier. Faster circles for Apples. Dr. Dobb’s Journal of Software Tools, 8(12):18–??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Quirk:1989:TPF

[Qui89] Kent Quirk. Translating PCX files.

Dr. Dobb’s Journal of Software Tools, 14(8):30–34, 36, 105–108, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Ragen:1985:Ia

[Rag85a] Alex Ragen. Of interest. Dr.

Dobb’s Journal of Software Tools, 10(6):??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Ragen:1985:Ib

[Rag85b] Alex Ragen. Of interest. Dr. Dobb’s Journal of Software Tools, 10(9):??, September 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Ragen:1985:Ic

[Rag85c] Alex Ragen. Of interest. Dr. Dobb’s Journal of Software Tools, 10(11):??, November 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Ralston:1989:DI

[Ral89] Kathleen Evans Ralston. 1988 DDJ index. Dr. Dobb’s Journal of Software Tools, 14(3):93–??, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Ramsey:1983:SSU

[Ram83] John C. Ramsey. SKIPIF — A submit utility for conditional command skipping in CPME. Dr. Dobb’s Journal of Software Tools, 8(8):100–??, August 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Raskin:19xx:GE

[Rasxx] Jef Raskin. Guest editorial. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Rasp:1984:HAA

[RB84] John Rasp and L. Brooks. How accurate is accurate? Dr. Dobb’s Journal of Software Tools, 9(2): 27–??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Relph:1985:CCM

[RCS85] Richard A. Relph, C SIG of PicoNet, and Silicon Valley Computer Society. C compilers for MS-DOS. Dr. Dobb’s Journal of Software Tools, 10(8):30–??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Ream:1982:PSO

[Rea82] Edward K. Ream. A portable screen-oriented editor. Dr. Dobb’s Journal of Software Tools, 7(1): 18–??, January 1982. CODEN DDJOEB. ISSN 1044-789X.

Ream:1983:RBCa

[Rea83a] Edward K. Ream. RED - A better C screen editor. Dr. Dobb’s Journal of Software Tools, 8(7):??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Ream:1983:RBCb

[Rea83b] Edward K. Ream. RED - A better C screen editor. Dr. Dobb’s Journal of Software Tools, 8(8):??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Ream:1986:PCB

[Rea86] Edward K. Ream. PL/68K: C becomes 68000 assembly language. Dr. Dobb’s Journal of Software Tools, 11(1):26–??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Reis:1988:TPU

[Rei88] Anthony J. Dos Reis. Theorem proving using semantic resolution. Dr. Dobb’s Journal of Software Tools, 13(4):50–??, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Reilly:19xx:R

[Reixx] Douglas Reilly. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Relph:1987:OCC

[Rel87a] Richard Relph. Optimizing compilers for C. Dr. Dobb’s Journal of Software Tools, 12(8):42–47, August 1987. CODEN DDJOEB. ISSN 0888-3076.

Relph:1987:PAC

[Rel87b] Richard Relph. Preparing for ANSI C. Dr. Dobb’s Journal of Software Tools, 12(8):16–18, 20, 22–23, August 1987. CODEN DDJOEB. ISSN 0888-3076.

Relph:1987:DAL

[Rel87c] Richard A. Relph. Developing 80386 applications … today. Dr. Dobb’s Journal of Software Tools, 12(7):16–??, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Relph:1988:SIV

[Rel88a] Richard A. Relph. Soft-ICE ver. 1.01 from Nu-Mega Technologies. Dr. Dobb’s Journal of Software Tools, 13(5):128–??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Relph:1988:STF

[Rel88b] Richard A. Relph. Speed trials: five Cs compared (compiler reviews). Dr. Dobb’s Journal of Software Tools, 13(8):18–20, 22, 24, 26–28, August 1988. CODEN DDJOEB. ISSN 0888-3076.

Relph:19xx:R

[Relxx] Richard A. Relph. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Relph:1986:BCC

[RHV86] Richard Relph, Steve Hahn, and Fred Viles. Benchmarking C compilers. Dr. Dobb’s Journal of Software Tools, 11(8):30–32, 34– 38, 40–44, 46–52, 58, 60, 62, 66, 104–109, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Rindsberg:1985:UPP

[Rin85] Donald Rindsberg. The ultimate parallel print spooler. Dr. Dobb’s Journal of Software Tools, 10(7):46–??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Ritter:1986:GCM

[Rit86] Terry Ritter. The great CRC mystery. Dr. Dobb’s Journal of Software Tools, 11(2):26–??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Ratcliff:1988:PMG

[RM88] John W. Ratcliff and David E.

Metzener. Pattern matching: The gestalt approach. Dr. Dobb’s Journal of Software Tools, 13(7):46, 47, 59–51, 68–72, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Robinson:1985:DFT

[RMB85] Howard H. Robinson, Philip D.

Morse, II, and Sidney A. Bowhill. Design of a Forth target compiler. Dr. Dobb’s Journal of Software Tools, 10(10):52–58, 60, 61, 63, 66– 74, 76, 78, 81–92, October 1985.

CODEN DDJOEB. ISSN 1044-789X.

Robichaux:1982:RDD

[Rob82] Joey Robichaux. Review of Devil’s DP Dictionary, by Stan Kelly-Bootle. Dr. Dobb’s Journal of Software Tools, 7(2):60–??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Robinson:1988:ESE

[Rob88] Phillip Robinson. Editorial: Software engineering. Dr. Dobb’s Journal of Software Tools, 13(??):4–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Rodriguez:1985:BDS

[Rod85] Suzanne Rodriguez. The Bob and Dennis show in Festschrift for Doctor Dobb. Dr. Dobb’s Journal of Software Tools, 10(2): 26–??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Rodman:1986:SCA

[Rod86] Richard Rodman. Series 32000 cross assembler. Dr. Dobb’s Journal of Software Tools, 11(12):48– 49, December 1986. CODEN DDJOEB. ISSN 0888-3076.

Rogers:1984:UIB

[Rog84] John Rogers. A Unix internals bibliography. Dr. Dobb’s Journal of Software Tools, 9(12):50–??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Rohling:1984:PPC

[Roh84] Theodore Rohling. PABX and the personal computer. Dr. Dobb’s Journal of Software Tools, 9(2):

60–62, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Rollins:1983:LGD

[Rol83] Daniel Rollins. 8088 line generator [with demo in BASIC]. Dr. Dobb’s Journal of Software Tools, 8(6):73– ??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Ross:1985:BRD

[Ros85] David Ross. BASIC riddle in DDC. Dr. Dobb’s Journal of Software Tools, 10(6):16–??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Rosford:1989:AIC

[Ros89] John Rosford. Avoiding INIT collisions at boot time. Dr. Dobb’s Journal of Software Tools, 14(??): 15–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Rothstein:19xx:R

[Rotxx] Joseph B. Rothstein. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Rovira:1989:PO

[Rov89] Charles A. Rovira. Persistent objects. Dr. Dobb’s Journal of Software Tools, 14(??):41–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Rowe:1989:OT

[Row89] Keith Rowe. Optimization technology. Dr. Dobb’s Journal of Software Tools, 14(6):56, 58–62, June 1989. CODEN DDJOEB. ISSN 0888-3076.

Rowe:19xx:R

[Rowxx] Robin Rowe. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Ruddy:1988:E

[Rud88] Sara Noah Ruddy. Editorial. Dr. Dobb’s Journal of Software Tools, 13(1):??, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Russell:1985:CMP

[Rus85] Stephen Russell. CP/M plus RSX as fix for bug in random disk read errors in CPME. Dr. Dobb’s Journal of Software Tools, 10(3): 108–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Russell:1989:GHA

[Rus89] Stephen Russell. A generic Heapsort algorithm in C. Dr. Dobb’s Journal of Software Tools, 14(8): 81–82, 84, 86, 88, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Rush:19xx:O

[Rusxx] Jeff Rush. OpenDoc. Dr. Dobb’s Journal of Software Tools, ??(??): 30–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Ruzinsky:1983:CPP

[Ruz83a] Steven A. Ruzinsky. Cosine procedure in Polyforth using the 8087 [letter]. Dr. Dobb’s Journal of Software Tools, 8(3):10–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Ruzinsky:1983:FMOa

[Ruz83b] Steven A. Ruzinsky. Fast matrix operations in Forth. Dr. Dobb’s Journal of Software Tools, 8(6):??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Ruzinsky:1983:FMOb

[Ruz83c] Steven A. Ruzinsky. Fast matrix operations in Forth. Dr. Dobb’s Journal of Software Tools, 8(7): ??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Ruzinsky:1983:FMOc

[Ruz83d] Steven A. Ruzinsky. Fast matrix operations in Forth. Dr. Dobb’s Journal of Software Tools, 8(8):??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Ruzinsky:1983:FMOd

[Ruz83e] Steven A. Ruzinsky. Fast matrix operations in Forth, part III. Dr. Dobb’s Journal of Software Tools, 8(8):44–48, 50–53, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Ruzinsky:1984:SMA

[Ruz84] Steven A. Ruzinsky. A simple minimax algorithm. Dr. Dobb’s Journal of Software Tools, 9(7):84–??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Raeside:1982:PCP

[RW82] D. E. Raeside and K. F. Wegner. A pattern classification program. Dr. Dobb’s Journal of Software Tools, 7(3):22–??, March 1982. CODEN DDJOEB. ISSN 1044-789X.

Rylander:1985:SSD

[Ryl85] Richard Rylander. Solid shape drawing on the Commodore 64. Dr. Dobb’s Journal of Software Tools, 10(5):50–??, May 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Saltsman:19xx:R

[Salxx] Tom Saltsman. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Scarpelli:1983:TZT

[Sca83] Anthony T. Scarpelli. TRS-80 8080 to Z80 translator. Dr. Dobb’s Journal of Software Tools, 8(4):60– 62, April 1983. CODEN DDJOEB. ISSN 1044-789X.

Scacchitti:1986:CT

[Sca86] Fred A. Scacchitti. The cryptographer’s toolbox. Dr. Dobb’s Journal of Software Tools, 11(5):58– ??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Schreiner:1984:CDS

[Sch84a] Axel Schreiner. cc — A driver for A Small-C programming system. Dr. Dobb’s Journal of Software Tools, 9(6):40–??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Schreiner:1984:PSC

[Sch84b] Axel Schreiner. p: A Small-C preprocessor. Dr. Dobb’s Journal of Software Tools, 9(7):46–??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Schlobohm:1985:TAP

[Sch85a] Dean Schlobohm. Tax advisor — A Prolog program analyzing income tax issues. Dr. Dobb’s Journal of Software Tools, 10(3):64–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Schmidt:1985:RMF

[Sch85b] Michael Schmidt. Review of Modula-Feb86 version 1.04 from Logitech. Dr. Dobb’s Journal of Software Tools, 10(2):74– ??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Schreiner:1985:UEa

[Sch85c] Axel Schreiner. UNIX exchange.

Dr. Dobb’s Journal of Software Tools, 10(6):??, June 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Schreiner:1985:UEb

[Sch85d] Axel Schreiner. UNIX exchange.

Dr. Dobb’s Journal of Software Tools, 10(9):??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Schildt:1989:BYO

[Sch89a] Herbert Schildt. Building your own C interpreter. Dr. Dobb’s Journal of Software Tools, 14(8): 38, 42, 45–46, 48–49, 110–122, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Schildt:1989:PMA

[Sch89b] Herbert Schildt. A Presentation Manager application template. Dr. Dobb’s Journal of Software Tools, 14(3):16–18, 21–22, 26–27, 81, March 1989. CODEN DDJOEB. ISSN 0888-3076.

Schmit:1989:MAC

[Sch89c] Mike Schmit. MS-DOS assemblers compared. Dr. Dobb’s Journal of Software Tools, 14(1):70–88, January 1989. CODEN DDJOEB. ISSN 0888-3076.

Schulman:1989:LWP

[Sch89d] Andrew Schulman. Linking while the program is running. Dr. Dobb’s Journal of Software Tools, 14(11): 46, 48, 50–52, 54–55, 102, 104, 106, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Schulman:19xx:R

[Schxx] Andrew Schulman. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Seal:1982:GLR

[Sea82] John J. Seal. Graphics low-res color for IBM PC BASIC setup program for in 16BST. Dr. Dobb’s Journal of Software Tools, 7(12): 12–??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Servello:1989:IMC

[Ser89] Mark Servello. Implementing multiple computer communications links. Dr. Dobb’s Journal of Software Tools, 14(10):18–20, 22, 24, 26, 81–85, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Seymour:1982:MMB

[Sey82] Henry A. Seymour. Memory management in BASIC. Dr. Dobb’s Journal of Software Tools, 7(11): 44–??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Seymour:1984:IP

[Sey84] Henry A. Seymour. An introduction to parsing. Dr. Dobb’s Journal of Software Tools, 9(12):78–80, 82–84, 86, December 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Shaw:1983:FVW

[Sha83] George W. Shaw, II. Forth-83 and vocabularies: When a page is not a page. Dr. Dobb’s Journal of Software Tools, 8(9):90–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1984:NSP

[Sha84] Namir Clement Shammas. NBA-SIC: A structured preprocessor for MBASIC. Dr. Dobb’s Journal of Software Tools, 9(1):24–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1986:MCI

[Sha86a] Namir Clement Shammas. Modula-2 compilers for the IBM PC. Dr. Dobb’s Journal of Software Tools, 11(10):48–??, October 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Shammas:1986:SPa

[Sha86b] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 11(6): ??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1986:SPb

[Sha86c] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 11 (8):??, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1986:SPc

[Sha86d] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 11(11): ??, November 1986. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1986:SPd

[Sha86e] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 11(12): ??, December 1986. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1986:TPP

[Sha86f] Namir Clement Shammas. Turbo Pascal procedural parameters, local modules in Modula-2. Dr. Dobb’s Journal of Software Tools, 11(12):96–??, December 1986. CO-DEN DDJOEB. ISSN 0888-3076.

Shaw:1986:FSP

[Sha86g] George W. Shaw, II. A Forth standards proposal: Extended control structures. Dr. Dobb’s Journal of Software Tools, 11(7):30–??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPa

[Sha87a] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(2): ??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPb

[Sha87b] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12 (3):??, March 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPc

[Sha87c] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(5): ??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPd

[Sha87d] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12 (8):??, August 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPe

[Sha87e] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(9): ??, September 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPf

[Sha87f] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12 (10):??, October 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:SPg

[Sha87g] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 12(11): ??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1987:TBC

[Sha87h] Namir Clement Shammas. True Basic challenges Modula-2. Dr. Dobb’s Journal of Software Tools, 12(5):124–126, 128–130, May 1987. CODEN DDJOEB. ISSN 0888-3076.

Shammas:1988:SP

[Sha88] Namir Clement Shammas. Structured programming. Dr. Dobb’s Journal of Software Tools, 13 (1):??, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Shiffman:1983:ADV

[Shi83] Richard Shiffman. Anatomy of a digital vector and curve generator. Dr. Dobb’s Journal of Software Tools, 8(7):66–68, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Shoemaker:1982:PHF

[Sho82a] D. C. Shoemaker. Put the H-8 front panel to work. Dr. Dobb’s Journal of Software Tools, 7(3): 17–??, March 1982. CODEN DDJOEB. ISSN 1044-789X.

Shoer:1982:NSE

[Sho82b] Larry Shoer. North star-to-epson connection. Dr. Dobb’s Journal of Software Tools, 7(5):26–??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Shoemaker:19xx:R

[Shoxx] D. C. Shoemaker. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Silvey:1984:RSE

[Sil84] Garry M. Silvey. Resident system extensions — RSX under CP/M plus. Dr. Dobb’s Journal of Software Tools, 9(7):36–??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Simon:1983:BPF

[Sim83] Michael L. Simon. Building a programmable frequency synthesizer. Dr. Dobb’s Journal of Software Tools, 8(12):80–84, 88, 90–91, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Simon:1986:REZ

[Sim86] Morris Simon. Review of echelon’s Z-System. Dr. Dobb’s Journal of Software Tools, 11(12):36– 38, 42–47, December 1986. CO-DEN DDJOEB. ISSN 0888-3076.

Siyan:1989:CCP

[Siy89] Karanjit S. Siyan. Coping with complex programs. Dr. Dobb’s Journal of Software Tools, 14(3): 60, 62, 65–66, 69, March 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Skjellum:1982:ACC

[Skj82a] Anthony Skjellum. Argum — A “C” command line processor. Dr. Dobb’s Journal of Software Tools, 7(8):10–??, August 1982. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1982:SFS

[Skj82b] Anthony Skjellum. Shell of five subroutines to expand wildcards under UNIX. Dr. Dobb’s Journal of Software Tools, 7(11):12– ??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1983:CUPa

[Skj83a] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 8(10): ??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1983:CPN

[Skj83b] Anthony Skjellum. C/Unix programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 8(10): 16–18, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1983:CUPb

[Skj83c] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 8(12): ??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPa

[Skj84a] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(2): ??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPb

[Skj84b] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(4):??, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPc

[Skj84c] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(6):??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPd

[Skj84d] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(8):??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPe

[Skj84e] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(9): ??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CPNa

[Skj84f] Anthony Skjellum. C/Unix programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9 (9):116–118, 120–122, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPf

[Skj84g] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(10): ??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CPNb

[Skj84h] Anthony Skjellum. C/Unix programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(10): 92–??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Smith:1983:FPA

[Smi83a] Burks Smith. Floating point arithmetic and numeric representation in computers [letter] in DDC. Dr. Dobb’s Journal of Software Tools, 8(2):55–??, February 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Smith:1983:SID

[Smi83b] Robert L. Smith. Signed integer division. Dr. Dobb’s Journal of Software Tools, 8(9):86– ??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Smith:1984:CMB

[Smi84] William Smith. CP/M 2.2 BIOS function: SELDSK. Dr. Dobb’s Journal of Software Tools, 9(3): 12–??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Smith:1985:LDP

[Smi85] Bruce A. Smith. Line drawing and point plotting routine for the IBM PC/PCjr in 16BST. Dr. Dobb’s Journal of Software Tools, 10(1): 109–??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Smith:1989:FSM

[Smi89] Donald W. Smith. Finite state machines for XModem. Dr. Dobb’s Journal of Software Tools, 14(10): 45–46, 48, 50, 92–103, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Softky:1986:MSP

[Sof86] Sheldon D. Softky. Modeling a system in PROLOG. Dr. Dobb’s Journal of Software Tools, 11 (4):46–??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Solntseff:1982:FPS

[Sol82] N. Solntseff. Forth programming style. Dr. Dobb’s Journal of Software Tools, 7(9):30–??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Solntseff:1983:PFM

[Sol83] N. Solntseff. Precompiled Forth modules. Dr. Dobb’s Journal of Software Tools, 8(9):82– ??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Sola:1985:RSP

[Sol85] Steven Anthony Sola. Review of SPSS/PC from SPSS, Inc. Dr. Dobb’s Journal of Software Tools, 10(4):84–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Sorens:1986:APB

[Sor86] Michael J. Sorens. Automatic porting between Pascal dialects. Dr. Dobb’s Journal of Software Tools, 11(3):50–??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Speiser:1983:SFP

[Spe83] Jeffrey Speiser. Savage floating point benchmark in PASCAL in 16BST. Dr. Dobb’s Journal of Software Tools, 8(11):112– ??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1987:Ea

[Spe87a] Tyler Sperry. Editorial. Dr. Dobb’s Journal of Software Tools, 12(9): ??, September 1987. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1987:Eb

[Spe87b] Tyler Sperry. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(10):??, October 1987. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1987:Ec

[Spe87c] Tyler Sperry. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(12):??, December 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Sperry:1987:RLa

[Spe87d] Tyler Sperry. Running light. Dr.

Dobb’s Journal of Software Tools, 12(9):??, September 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Sperry:1987:RLb

[Spe87e] Tyler Sperry. Running light. Dr. Dobb’s Journal of Software Tools, 12(10):??, October 1987. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1987:RLc

[Spe87f] Tyler Sperry. Running light. Dr. Dobb’s Journal of Software Tools, 12(12):??, December 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Sperry:1988:VCF

[Spe88a] Tyler Sperry. 386 vs. 030: The crowded fast lane. Dr. Dobb’s Journal of Software Tools, 13(1): 16–??, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1988:Ea

[Spe88b] Tyler Sperry. Editorial. Dr. Dobb’s Journal of Software Tools, 13(2):??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1988:Eb

[Spe88c] Tyler Sperry. Editorial. Dr. Dobb’s Journal of Software Tools, 13(3):??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1988:Ec

[Spe88d] Tyler Sperry. Editorial. Dr. Dobb’s Journal of Software Tools, 13(4): ??, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1988:RLa

[Spe88e] Tyler Sperry. Running light. Dr. Dobb’s Journal of Software Tools, 13(2):??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1988:RLb

[Spe88f] Tyler Sperry. Running light. Dr.

Dobb’s Journal of Software Tools, 13(3):??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Sperry:1988:RLc

[Spe88g] Tyler Sperry. Running light. Dr.

Dobb’s Journal of Software Tools, 13(4):??, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Staneff:1983:SHF

[Sta83] John Staneff. Small-C help facility. Dr. Dobb’s Journal of Software Tools, 8(10):40–??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Staneff:1984:SCG

[Sta84] John Staneff. Small-C concordance generator. Dr. Dobb’s Journal of Software Tools, 9(8):86–90, 92–96, 98–105, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Stallman:1985:GM

[Sta85] Richard Stallman. The GNU manifesto. Dr. Dobb’s Journal of Software Tools, 10(3):30–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Starr:1989:CLM

[Sta89] Robert F. Starr. C list manager. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):48–54, 90–91, 93, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Stallings:19xx:R

[Staxx] William Stallings. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Stetson:1983:RUG

[Ste83] Eunice B. Stetson. Review of User’s Guidebook to Digital CMOS Integrated Circuits, by Eugene R. Hnatek. Dr. Dobb’s Journal of Software Tools, 8(5):70–??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Steele:1985:PRD

[Ste85] Alfred Steele. ProDOS and the RAM disk. Dr. Dobb’s Journal of Software Tools, 10(12):84–??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Steinman:1986:WMT

[Ste86] Jan W. Steinman. The worm memory test in TRTA. Dr. Dobb’s Journal of Software Tools, 11(9): 114–??, September 1986. CODEN DDJOEB. ISSN 1044-789X.

Steppe:1987:FCA

[Ste87] Tom Steppe. File comparison algorithms. Dr. Dobb’s Journal of Software Tools, 12(9):28–??, September 1987. CODEN DDJOEB. ISSN 0888-3076.

Stein:1988:OOP

[Ste88a] Jacob Stein. Object-oriented programming and databases. Dr. Dobb’s Journal of Software Tools, 13(3):18–??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1988:CPR

[Ste88b] A. Stevens. C programming: off and running …. Dr. Dobb’s Journal of Software Tools, 13(8):

98, 101–102, 104, 106–107, 109– 110, 113, August 1988. CODEN DDJOEB. ISSN 0888-3076.

Stevens:1988:CPS

[Ste88c] A. Stevens. C programming:

screen control, programming as art, and C++. Dr. Dobb’s Journal of Software Tools, 13(10):108, 111–115, 128–130, 133–134, October 1988. CODEN DDJOEB. ISSN 0888-3076.

Stevens:1988:CPa

[Ste88d] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 13(8):??, August 1988. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1988:CPb

[Ste88e] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 13(9):??, September 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1988:CPc

[Ste88f] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1988:CPd

[Ste88g] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 13(11):??, November 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1988:CPe

[Ste88h] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 13(12):??, December 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Stewart:1988:AVD

[Ste88i] James Stewart. Ada validation: The diary of a vendor. Dr. Dobb’s Journal of Software Tools, 13(9): 45–??, September 1988. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CLL

[Ste89a] A. Stevens. C++ and linked lists. Dr. Dobb’s Journal of Software Tools, 14(11):133–134, 137–140, 152–153, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Stevens:1989:CPG

[Ste89b] Al Stevens. C programmer’s guide to C++. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):32–41, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPa

[Ste89c] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 14(1):??, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPb

[Ste89d] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 14(2):??, February 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPc

[Ste89e] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 14(3):??, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPd

[Ste89f] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 14(4):??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPe

[Ste89g] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(5):??, May 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPf

[Ste89h] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(6):??, June 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPh

[Ste89i] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(7):??, July 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPi

[Ste89j] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(8):??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPj

[Ste89k] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(9):??, September 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPk

[Ste89l] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(10):??, October 1989. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPl

[Ste89m] Al Stevens. C programming. Dr.

Dobb’s Journal of Software Tools, 14(11):??, November 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1989:CPm

[Ste89n] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 14(12):??, December 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1989:IDR

[Ste89o] Al Stevens. Interview with Dennis Ritchie and Bjarne Stroustrup. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):8–??, Winter 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:19xx:R

[Stexx] Al Stevens. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??): ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Stith:1984:BAU

[Sti84] John E. Stith. Balancing act: The ultimate checkbook balancing program. Dr. Dobb’s Journal of Software Tools, 9(7):110–??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Stitt:1988:UAC

[Sti88] Martin Stitt. Using action charts for software development. Dr. Dobb’s Journal of Software Tools, 13(9):18–33, 92–93, September 1988. CODEN DDJOEB. ISSN 0888-3076.

Streich:1989:PMD

[Str89] Mark Streich. A parallel make with DESQview. Dr. Dobb’s Journal of Software Tools, 14(11):28, 30, 32, 34, 36, 86, 92–94, 97–99, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Suman:1987:WWH

[Sum87] Michael Suman. What’s wrong with high-level languages in VP. Dr. Dobb’s Journal of Software Tools, 12(2):14–??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1984:Ia

[Sut84a] R. P. Sutherland. Of interest. Dr.

Dobb’s Journal of Software Tools, 9(10):??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1984:Ib

[Sut84b] R. P. Sutherland. Of interest. Dr.

Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1984:Ic

[Sut84c] R. P. Sutherland. Of interest. Dr.

Dobb’s Journal of Software Tools, 9(12):??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1985:Ia

[Sut85a] R. P. Sutherland. Of interest. Dr.

Dobb’s Journal of Software Tools, 10(1):??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1985:Ib

[Sut85b] R. P. Sutherland. Of interest. Dr.

Dobb’s Journal of Software Tools, 10(2):??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1985:Ic

[Sut85c] R. P. Sutherland. Of interest. Dr.

Dobb’s Journal of Software Tools, 10(3):??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1985:Id

[Sut85d] R. P. Sutherland. Of interest. Dr. Dobb’s Journal of Software Tools, 10(4):??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1985:RTM

[Sut85e] R. P. Sutherland. Review of Turing’s Man: Western Culture in the Computer Age, by J. David Bolter. Dr. Dobb’s Journal of Software Tools, 10(3):122–??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Sutherland:1988:E

[Sut88] R. P. Sutherland. Editorial. Dr. Dobb’s Journal of Software Tools, 13(??):??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Ea

[Swa84a] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 9(5):??, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Eb

[Swa84b] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 9(6):??, June 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Ec

[Swa84c] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 9(7):??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Ed

[Swa84d] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 9(8):??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Ee

[Swa84e] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(9):??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Ef

[Swa84f] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(10):??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Eg

[Swa84g] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:Eh

[Swa84h] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(12):??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:SDa

[Swa84i] Michael Swaine. The software designer. Dr. Dobb’s Journal of Software Tools, 9(10):??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1984:SDb

[Swa84j] Michael Swaine. The software designer. Dr. Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ea

[Swa85a] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 10(1):??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Eb

[Swa85b] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(2):??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ec

[Swa85c] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(3):??, March 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ed

[Swa85d] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(4):??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ee

[Swa85e] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(5):??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ef

[Swa85f] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(6):??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Eg

[Swa85g] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(7):??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Eh

[Swa85h] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 10(8):??, August 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ei

[Swa85i] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 10(9):??, September 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ej

[Swa85j] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 10(10):??, October 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:Ek

[Swa85k] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 10(11):??, November 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1985:El

[Swa85l] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 10(12):??, December 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1985:I

[Swa85m] Michael Swaine. Of interest. Dr.

Dobb’s Journal of Software Tools, 10(7):??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:SDa

[Swa85n] Michael Swaine. The software designer. Dr. Dobb’s Journal of Software Tools, 10(1):??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:SDb

[Swa85o] Michael Swaine. The software designer. Dr. Dobb’s Journal of Software Tools, 10(2):??, February 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:SDc

[Swa85p] Michael Swaine. The software designer. Dr. Dobb’s Journal of Software Tools, 10(5):??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:SDd

[Swa85q] Michael Swaine. The software designer. Dr. Dobb’s Journal of Software Tools, 10(9):??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1985:WOE

[Swa85r] Michael Swaine. Windowing operating environments: TopView GEM and Windows. Dr. Dobb’s Journal of Software Tools, 10(12): 28–33, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Ea

[Swa86a] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 11(1):??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Eb

[Swa86b] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 11(2):??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Ec

[Swa86c] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 11(3):??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Ed

[Swa86d] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 11(4):??, April 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Ee

[Swa86e] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(8):??, August 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Ef

[Swa86f] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(9):??, September 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Eg

[Swa86g] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(10):??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Eh

[Swa86h] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(11):??, November 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1986:Ei

[Swa86i] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(12):??, December 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFa

[Swa86j] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 11(6):??, June 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFb

[Swa86k] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 11(7):??, July 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFc

[Swa86l] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 11(8):??, August 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFd

[Swa86m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 11(9):??, September 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFe

[Swa86n] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 11(10):??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFf

[Swa86o] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 11(11):??, November 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1986:SFg

[Swa86p] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 11(12):??, December 1986. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Ea

[Swa87a] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 12(1):??, January 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Eb

[Swa87b] Michael Swaine. Editorial. Dr. Dobb’s Journal of Software Tools, 12(2):??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Ec

[Swa87c] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(3):??, March 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Ed

[Swa87d] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(4):??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Ee

[Swa87e] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(5):??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Ef

[Swa87f] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(6):??, June 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Eg

[Swa87g] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(7):??, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Eh

[Swa87h] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(8):??, August 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:Ei

[Swa87i] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 12(11):??, November 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFa

[Swa87j] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(1):??, January 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFb

[Swa87k] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(2):??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFc

[Swa87l] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(3):??, March 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFd

[Swa87m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(4):??, April 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFe

[Swa87n] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(5):??, May 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFf

[Swa87o] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(6):??, June 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFg

[Swa87p] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 12(7):??, July 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFh

[Swa87q] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 12(8):??, August 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFi

[Swa87r] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 12(9):??, September 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFj

[Swa87s] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 12(10):??, October 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFk

[Swa87t] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 12(11):??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1987:SFl

[Swa87u] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 12(12):??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SPT

[Swa88a] M. Swaine. SD 88: Prolog tools and transputers (editorial). Dr. Dobb’s Journal of Software Tools, 13(6):110–??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:E

[Swa88b] Michael Swaine. Editorial. Dr.

Dobb’s Journal of Software Tools, 13(11):80–??, November 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPa

[Swa88c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(5):??, May 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPb

[Swa88d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(6):??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPc

[Swa88e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(7):??, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPd

[Swa88f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(8):??, August 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPe

[Swa88g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(9):??, September 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPf

[Swa88h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPg

[Swa88i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(11):??, November 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:PPh

[Swa88j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 13(12):??, December 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:RLa

[Swa88k] Michael Swaine. Running light.

Dr. Dobb’s Journal of Software Tools, 13(1):??, January 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:RLb

[Swa88l] Michael Swaine. Running light.

Dr. Dobb’s Journal of Software Tools, 13(9):??, September 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFa

[Swa88m] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 13(1):??, January 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFb

[Swa88n] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 13(2):??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFc

[Swa88o] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 13(3):??, March 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFd

[Swa88p] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(4):??, April 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFe

[Swa88q] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(5):??, May 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFf

[Swa88r] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(6):??, June 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFg

[Swa88s] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(7):??, July 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFh

[Swa88t] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(8):??, August 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFi

[Swa88u] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(9):??, September 1988. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1988:SFj

[Swa88v] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 13(12):??, December 1988.

CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:TEN

[Swa89a] M. Swaine. Two early neural net implementations. Dr. Dobb’s Journal of Software Tools, 14(11): 124, 128–129, 131, November 1989. CODEN DDJOEB. ISSN 0888-3076.

Swaine:1989:PPa

[Swa89b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(1):??, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPb

[Swa89c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(2):??, February 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPc

[Swa89d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(3):??, March 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPd

[Swa89e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(4):??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPe

[Swa89f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(5):??, May 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPf

[Swa89g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(6):??, June 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPg

[Swa89h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(7):??, July 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPh

[Swa89i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(8):??, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPi

[Swa89j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(9):??, September 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPj

[Swa89k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(10):??, October 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPk

[Swa89l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(11):??, November 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:PPl

[Swa89m] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 14(12):??, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFa

[Swa89n] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 14(1):??, January 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFb

[Swa89o] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 14(2):??, February 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFc

[Swa89p] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 14(3):??, March 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFd

[Swa89q] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 14(4):??, April 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFe

[Swa89r] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 14(5):??, May 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFf

[Swa89s] Michael Swaine. Swaine’s flames.

Dr. Dobb’s Journal of Software Tools, 14(6):??, June 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFg

[Swa89t] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 14(7):??, July 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFh

[Swa89u] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 14(8):??, August 1989. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFi

[Swa89v] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 14(9):??, September 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFj

[Swa89w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 14(10):??, October 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFk

[Swa89x] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 14(11):??, November 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1989:SFl

[Swa89y] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 14(12):??, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Swaine:19xx:RN

[Swaxx] Michael Swaine. Reviews and notes. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx.

CODEN DDJOEB. ISSN 1044-789X.

Takara:1984:PDU

[Tak84] Ken Takara. Program design using pseudocode. Dr. Dobb’s Journal of Software Tools, 9(3):70–77, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Tan:1983:IST

[Tan83] B. T. G. Tan. Instructions of the 6800 and 6502 SBC TSX and TXS. Dr. Dobb’s Journal of Software Tools, 8(5):67–??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Taylor:1982:CRG

[Tay82] Jeffrey L. Taylor. Cross-reference generator in C — A program conversion aid. Dr. Dobb’s Journal of Software Tools, 7(6):50–??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Taylor:1983:SRF

[Tay83] Robert Taylor. SEND and RCV:

A Forth implementation of the XMODEM protocol. Dr. Dobb’s Journal of Software Tools, 8(9): 66–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1989:SCP

[TB89] David Thomas and Randolph Best. Smalltalk + C: the power of two. Dr. Dobb’s Journal of Software Tools, 14(8):16–18, 94– 98, 100, August 1989. CODEN DDJOEB. ISSN 0888-3076.

Tello:1987:AIa

[Tel87a] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(2):??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIb

[Tel87b] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(3):??, March 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIc

[Tel87c] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(4):??, April 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AId

[Tel87d] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(5):??, May 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIe

[Tel87e] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(6):??, June 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIf

[Tel87f] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(7):??, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIg

[Tel87g] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(8):??, August 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIh

[Tel87h] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(9):??, September 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIi

[Tel87i] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 12(11):??, November 1987. CODEN DDJOEB. ISSN 1044-789X.

Tello:1987:AIX

[Tel87j] Ernest R. Tello. Artificial intelligence: the Xerox 1186 Lisp Machine. Dr. Dobb’s Journal of Software Tools, 12(7):118, 120– 122, 124–125, July 1987. CODEN DDJOEB. ISSN 0888-3076.

Tello:1987:OLP

[Tel87k] Ernest R. Tello. Object-oriented Lisp on PCs. Dr. Dobb’s Journal of Software Tools, 12(5):132– 134, 136, May 1987. CODEN DDJOEB. ISSN 0888-3076.

Tello:1987:OP

[Tel87l] Ernest R. Tello. Object-oriented programming. Dr. Dobb’s Journal of Software Tools, 12(3):126, 128– 129, 132–134, March 1987. CO-DEN DDJOEB. ISSN 0888-3076.

Tello:1987:OPS

[Tel87m] Ernest R. Tello. Object-oriented programming in SCOOPS. Dr. Dobb’s Journal of Software Tools, 12(6):116–117, 98–99, June 1987. CODEN DDJOEB. ISSN 0888-

Tello:1988:TP

[Tel88a] E. Tello. Turbo Prolog 2.0. Dr. Dobb’s Journal of Software Tools, 13(10):54–55, October 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Tello:1988:CC

[Tel88b] E. R. Tello. A class-ier C. Dr. Dobb’s Journal of Software Tools, 13(12):74–78, December 1988. CO-DEN DDJOEB. ISSN 0888-3076.

Tello:1988:OOL

[Tel88c] E. R. Tello. Objective-C (objectoriented language). Dr. Dobb’s Journal of Software Tools, 13(8): 56–58, 60, 63, 65–66, 69, August 1988. CODEN DDJOEB. ISSN 0888-3076.

Tello:1988:AI

[Tel88d] Ernest R. Tello. Artificial intelligence. Dr. Dobb’s Journal of Software Tools, 13(1):??, January 1988. CODEN DDJOEB. ISSN 1044-789X.

Tello:19xx:R

[Telxx] Ernest R. Tello. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1984:SGM

[TF84] John Thomas and Guy Fletcher. Scistar: Greek and math symbols with Wordstar. Dr. Dobb’s Journal of Software Tools, 9(8):26–29, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1986:SDE

[Tho86] John A. Thomas. Survey of data encryption in DOL. Dr.

Dobb’s Journal of Software Tools, 11(6):16–??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1987:RL

[Tho87] Levi Thomas. Running light. Dr.

Dobb’s Journal of Software Tools, 12(6):??, June 1987. CODEN DDJOEB. ISSN 1044-789X.

Thomson:1989:UEM

[Tho89] Paul Thomson. Using extended memory on the PC AT. Dr. Dobb’s Journal of Software Tools, 14(1): 106, 108, 119, January 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Tichenor:1988:VAC

[Tic88] Mark Tichenor. Virtual arrays in C. Dr. Dobb’s Journal of Software Tools, 13(5):46–??, May 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Tilton:1985:ZAL

[Til85] David S. Tilton. Z80 assembly language riddle in DDC. Dr. Dobb’s Journal of Software Tools, 10(1): 17–??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Ting:1982:FDF

[Tin82] C. H. Ting. Formal definition of FORTH. Dr. Dobb’s Journal of Software Tools, 7(2):19–??, February 1982. CODEN DDJOEB. ISSN 1044-789X.

Ting:1983:GF

[Tin83] C. H. Ting. GO in Forth.

Dr. Dobb’s Journal of Software Tools, 8(9):54–56, 59–60, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Tonkin:1988:GB

[Ton88a] Bruce W. Tonkin. Getting down to BASICS. Dr. Dobb’s Journal of Software Tools, 13(7):52–??, July 1988. CODEN DDJOEB. ISSN 1044-789X.

Tonkin:1988:IEB

[Ton88b] Bruce W. Tonkin. Inserting elements into a Basic integer array. Dr. Dobb’s Journal of Software Tools, 13(11):66–67, November 1988. CODEN DDJOEB. ISSN 0888-3076.

Tonkin:1989:LDP

[Ton89a] Bruce W. Tonkin. Languageindependent dynamic pseudostructures. Dr. Dobb’s Journal of Software Tools, 14(5):39–48, 94– 96, May 1989. CODEN DDJOEB. ISSN 0888-3076.

Tonkin:1989:PLB

[Ton89b] Bruce W. Tonkin. PDQ: Less baggage faster code [review]. Dr. Dobb’s Journal of Software Tools, 14(12):88, 90, 92–94, December 1989. CODEN DDJOEB. ISSN 1044-789X.

Tonkin:1989:QPC

[Ton89c] Bruce W. Tonkin. QuickPak professional from crescent software [review] in ER. Dr. Dobb’s Journal of Software Tools, 14(4):96–??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Trask:1985:RBV

[Tra85] Matthew Trask. Review of Better-BASIC version 1.1 from Summit Software. Dr. Dobb’s Journal of Software Tools, 10(5):108–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Tracy:1987:FCa

[Tra87a] Martin Tracy. The Forth column.

Dr. Dobb’s Journal of Software Tools, 12(10):??, October 1987. CODEN DDJOEB. ISSN 1044-789X.

Tracy:1987:FCb

[Tra87b] Martin Tracy. The Forth column.

Dr. Dobb’s Journal of Software Tools, 12(12):??, December 1987. CODEN DDJOEB. ISSN 1044-789X.

Tracy:1987:FSP

[Tra87c] Martin Tracy. A Forth standard prelude. Dr. Dobb’s Journal of Software Tools, 12(10):40–??, October 1987. CODEN DDJOEB. ISSN 0888-3076.

Tracy:1988:MRC

[Tra88a] M. Tracy. More real-time connections (Forth). Dr. Dobb’s Journal of Software Tools, 13(8):114, 116–118, August 1988. CODEN DDJOEB. ISSN 0888-3076.

Tracy:1988:FCa

[Tra88b] Martin Tracy. The Forth column.

Dr. Dobb’s Journal of Software Tools, 13(2):??, February 1988. CODEN DDJOEB. ISSN 1044-789X.

Tracy:1988:FCb

[Tra88c] Martin Tracy. The Forth column.

Dr. Dobb’s Journal of Software Tools, 13(4):??, April 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Tracy:1988:FCc

[Tra88d] Martin Tracy. The Forth column. Dr. Dobb’s Journal of Software Tools, 13(6):??, June 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Tracy:1988:FCd

[Tra88e] Martin Tracy. The Forth column. Dr. Dobb’s Journal of Software Tools, 13(8):??, August 1988. CO-DEN DDJOEB. ISSN 1044-789X.

Tracy:1988:FCe

[Tra88f] Martin Tracy. The Forth column. Dr. Dobb’s Journal of Software Tools, 13(10):??, October 1988. CODEN DDJOEB. ISSN 1044-789X.

Thersites:1984:IKE

[TT84a] Joan Thersites and John A. Thomas. An infinite key encryption system. Dr. Dobb’s Journal of Software Tools, 9(8):44–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1984:IKE

[TT84b] John A. Thomas and Joan Thersites. Infinite key encryption system. Dr. Dobb’s Journal of Software Tools, 9(8):44–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1986:CBP

[TT86] Levi Thomas and Nicholas Turner. The CompuServe B protocol: A better way to send files. Dr. Dobb’s Journal of Software Tools, 11(6): 38–40, 90, 92, 94–99, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Thomas:1987:TEM

[TT87] Levi Thomas and Nicholas Turner.

Text editors: In matters of taste …. Dr. Dobb’s Journal of Software Tools, 12(2):16–??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Thompson:1988:TKB

[TT88] Bill Thompson and Beverly Thompson. Topics in knowledgebased languages. Dr. Dobb’s Journal of Software Tools, 13(4):40–42, 43, 45, 47, 49, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:CCM

[Tur86a] Nicholas Turner. Code compression with mini-interpreters in TRTA. Dr. Dobb’s Journal of Software Tools, 11(5):110–??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:Ea

[Tur86b] Nicholas Turner. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(5):??, May 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:Eb

[Tur86c] Nicholas Turner. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(6):??, June 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:Ec

[Tur86d] Nicholas Turner. Editorial. Dr.

Dobb’s Journal of Software Tools, 11(7):??, July 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:RLa

[Tur86e] Nicholas Turner. Running light. Dr. Dobb’s Journal of Software Tools, 11(8):??, August 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Turner:1986:RLb

[Tur86f] Nicholas Turner. Running light. Dr. Dobb’s Journal of Software Tools, 11(9):??, September 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:RLc

[Tur86g] Nicholas Turner. Running light. Dr. Dobb’s Journal of Software Tools, 11(10):??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:RLd

[Tur86h] Nicholas Turner. Running light. Dr. Dobb’s Journal of Software Tools, 11(11):??, November 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:RLe

[Tur86i] Nicholas Turner. Running light. Dr. Dobb’s Journal of Software Tools, 11(12):??, December 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1986:SMO

[Tur86j] Nicholas Turner. A simple multitasking operating system for realtime applications. Dr. Dobb’s Journal of Software Tools, 11(1): 44–??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Turner:1987:CWT

[Tur87a] Nicholas Turner. 680XX computers: Where are they going? Dr. Dobb’s Journal of Software Tools, 12(1):16–18, January 1987. CO-DEN DDJOEB. ISSN 0888-3076.

Turner:1987:EIL

[Tur87b] Nicholas Turner. 680xx experimental interpreted language proposal in TRTA. Dr. Dobb’s Journal of Software Tools, 12(1): 126–??, January 1987. CODEN DDJOEB. ISSN 1044-789X.

Turner:1987:RLa

[Tur87c] Nicholas Turner. Running light.

Dr. Dobb’s Journal of Software Tools, 12(1):??, January 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Turner:1987:RLb

[Tur87d] Nicholas Turner. Running light.

Dr. Dobb’s Journal of Software Tools, 12(2):??, February 1987. CODEN DDJOEB. ISSN 1044-789X.

Turner:1987:RLc

[Tur87e] Nicholas Turner. Running light.

Dr. Dobb’s Journal of Software Tools, 12(3):??, March 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Turner:1987:RLd

[Tur87f] Nicholas Turner. Running light.

Dr. Dobb’s Journal of Software Tools, 12(4):??, April 1987. CO-DEN DDJOEB. ISSN 1044-789X.

Turner:1989:MWH

[Tur89] Kenneth Turner. Modifying WizardCopy for hard disks [sidebar] in “WizardCopy for fast Backups” (D. Gaspar). Dr. Dobb’s Journal of Software Tools, 14(??):45– ??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Ullman:1984:FHT

[Ull84] Ronald Ullman. Fast Hartley transform [letter]. Dr. Dobb’s Journal of Software Tools, 9(12): 8–??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Unger:1985:RAP

[Ung85] Stefan H. Unger. Review of APL*PLUS/PC version 3.1. Dr. Dobb’s Journal of Software Tools, 10(1):88–??, January 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Upstill:1988:PCG

[Ups88] Steve Upstill. Photorealism in computer graphics. Dr. Dobb’s Journal of Software Tools, 13(11): 18–30, November 1988. CODEN DDJOEB. ISSN 0888-3076.

Valdes:19xx:IIOb

[Valxxa] Ray Valdés. Implementing interoperable objects. Dr. Dobb’s Journal of Software Tools, ??(??):62– ??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Valdes:19xx:IIOa

[Valxxb] Ray Valdés. Introducing interoperable objects. Dr. Dobb’s Journal of Software Tools, ??(??):4–??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Valdes:19xx:R

[Valxxc] Ray Valdés. Reviews. Dr. Dobb’s Journal of Software Tools, ??(??):

??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

VanZandt:1982:MAS

[Van82] James R. Van Zandt. Modifications and additions to small C compiler [letter]. Dr. Dobb’s Journal of Software Tools, 7(10):6–??, October 1982. CODEN DDJOEB. ISSN 1044-789X.

Varhol:19xx:R

[Varxx] Peter D. Varhol. Reviews. Dr.

Dobb’s Journal of Software Tools, ??(??):??, ???? 19xx. CODEN DDJOEB. ISSN 1044-789X.

Vasile:1984:AF

[Vas84] Eddy Vasile. The accent finder.

Dr. Dobb’s Journal of Software Tools, 9(5):82–86, 88–96, May 1984. CODEN DDJOEB. ISSN 1044-789X.

Vose:1988:RTA

[Vos88] G. Michael Vose. OS/2’s a realtime alternative. Dr. Dobb’s Journal of Software Tools, 13(6):42–??, June 1988. CODEN DDJOEB. ISSN 1044-789X.

Vose:1989:BA

[VW89] G. Michael Vose and David Weil.

A benchmark apologia. Dr. Dobb’s Journal of Software Tools, 14(2): 36–38, 40, 43, February 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Walker:1986:DAM

[WA86] Bill Walker and Stephen Alexander. Data abstraction with Modula-2. Dr. Dobb’s Journal of Software Tools, 11(2):62–64, 94, 95, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Walworth:1984:VUI

[Wal84] Alan Walworth. Varieties of Unix — an introductory guide to microcomputer Unixes. Dr. Dobb’s Journal of Software Tools, 9(12): 24–32, 34–37, December 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Walsh:1985:MBS

[Wal85] Dale Walsh. Modems: 2400 bit/sec and beyond. Dr. Dobb’s Journal of Software Tools, 10(6):56–??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Ward:1984:GRC

[War84] Terry A. Ward. A guide to resources for the C programmer. Dr. Dobb’s Journal of Software Tools, 9(11):74–78, 80–82, 84–86, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Ward:1985:MPS

[War85] Terry A. Ward. Monty plays Scrabble: Issues in human factors engineering. Dr. Dobb’s Journal of Software Tools, 10(4):18–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Watson:1983:RCS

[Wat83] Trina Watson. Review of Crunchers-21 — Simple Games for the Timex/Sinclair 1000 2x, by Yin Chiu and Henry Mullish. Dr. Dobb’s Journal of Software Tools, 8(2):67–??, February 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Waters:1989:OCM

[Wat89a] Bryan Waters. Object C and the Macintosh control panel. Dr.

Dobb’s Journal of Software Tools, 14(??):51–??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Waters:1989:WMD

[Wat89b] Bryan Waters. Writing Macintosh device drivers. Dr. Dobb’s Journal of Software Tools, 14(??):37– ??, ???? 1989. CODEN DDJOEB. ISSN 1044-789X.

Weiger:1983:LCF

[Wei83] Wendy Weiger. Linking CP/M functions to your high-level program. Dr. Dobb’s Journal of Software Tools, 8(2):32–42, 44– 54, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Weissman:1986:SME

[Wei86] Gregg Weissman. Speeding MS DOS execution. Dr. Dobb’s Journal of Software Tools, 11(3):44–??, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Weiskamp:1989:RCE

[Wei89] Keith Weiskamp. Review of “C”erious toolkit from TSR systems ltd. in ER. Dr. Dobb’s Journal of Software Tools, 14(4):92–??, April 1989. CODEN DDJOEB. ISSN 1044-789X.

Weiskamp:1989:IE

[WH89] Keith Weiskamp and Loren Heiny.

An icon editor. Dr. Dobb’s Journal of Software Tools, 14(7):24–33, 84– 88, July 1989. CODEN DDJOEB. ISSN 0888-3076.

Wharton:1983:DCM

[Wha83] John H. Wharton. In defense of CP/M plus [letter]. Dr. Dobb’s Journal of Software Tools, 8(9): 10–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

White:1987:CID

[Whi87] Ronald G. White. Compressing image data with quadtrees. Dr. Dobb’s Journal of Software Tools, 12(3):16–??, March 1987. CODEN DDJOEB. ISSN 0888-3076.

White:1989:HD

[Whi89] Benjamin White. Hamming-Code decoding. Dr. Dobb’s Journal of Software Tools, 14(10):52, 54, 56, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Wiesenberg:1982:Ia

[Wie82a] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 7(5):??, May 1982. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1982:Ib

[Wie82b] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 7(6):??, June 1982. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1982:Ic

[Wie82c] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 7(7):??, July 1982. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1982:Id

[Wie82d] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 7(9):??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1982:Ie

[Wie82e] Michael Wiesenberg. Of interest.

Dr. Dobb’s Journal of Software Tools, 7(10):??, October 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1982:If

[Wie82f] Michael Wiesenberg. Of interest.

Dr. Dobb’s Journal of Software Tools, 7(11):??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1982:Ig

[Wie82g] Michael Wiesenberg. Of interest.

Dr. Dobb’s Journal of Software Tools, 7(12):??, December 1982. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ia

[Wie83a] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(1):??, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ib

[Wie83b] Michael Wiesenberg. Of interest.

Dr. Dobb’s Journal of Software Tools, 8(2):??, February 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ic

[Wie83c] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(3):??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Id

[Wie83d] Michael Wiesenberg. Of interest.

Dr. Dobb’s Journal of Software Tools, 8(4):??, April 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ie

[Wie83e] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(5):??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:If

[Wie83f] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(6):??, June 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ig

[Wie83g] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(7):??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ih

[Wie83h] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(8):??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ii

[Wie83i] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(9):??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ij

[Wie83j] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(10):??, October 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Ik

[Wie83k] Michael Wiesenberg. Of interest. Dr. Dobb’s Journal of Software Tools, 8(11):??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1983:Il

[Wie83l] Michael Wiesenberg. Of interest.

Dr. Dobb’s Journal of Software Tools, 8(12):??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1984:CCa

[Wie84a] Michael Wiesenberg. Computer calisthenics. Dr. Dobb’s Journal of Software Tools, 9(7):??, July 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1984:CCb

[Wie84b] Michael Wiesenberg. Computer calisthenics. Dr. Dobb’s Journal of Software Tools, 9(11):??, November 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiesenberg:1984:CCc

[Wie84c] Michael Wiesenberg. Computer calisthenics. Dr. Dobb’s Journal of Software Tools, 9(12):??, December 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Ea

[Wig83a] Reynold Wiggins. Editorial. Dr.

Dobb’s Journal of Software Tools, 8(2):??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Eb

[Wig83b] Reynold Wiggins. Editorial. Dr.

Dobb’s Journal of Software Tools, 8(4):??, April 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Ec

[Wig83c] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(5):??, May 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Ed

[Wig83d] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(7):??, July 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Ee

[Wig83e] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(8):??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Ef

[Wig83f] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(9):??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Eg

[Wig83g] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(10):??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Eh

[Wig83h] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(11):??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1983:Ei

[Wig83i] Reynold Wiggins. Editorial. Dr. Dobb’s Journal of Software Tools, 8(12):??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1984:Ea

[Wig84a] Reynold Wiggins. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(1):??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1984:Eb

[Wig84b] Reynold Wiggins. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(2):??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1984:Ec

[Wig84c] Reynold Wiggins. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(3):??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1984:Ed

[Wig84d] Reynold Wiggins. Editorial. Dr.

Dobb’s Journal of Software Tools, 9(4):??, April 1984. CODEN DDJOEB. ISSN 1044-789X.

Wiggins:1984:ROC

[Wig84e] W. James Wiggins. Review of Oubliette [game] from Centaur software. Dr. Dobb’s Journal of Software Tools, 9(2):80–??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Wilcox:1983:SFU

[Wil83a] Alan D. Wilcox. Serial-to-parallel — a flexible utility box. Dr. Dobb’s Journal of Software Tools, 8(8): 28, 31–35, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Willman:1983:DOC

[Wil83b] Bryan M. Willman. Descriptive operators in C [letter]. Dr. Dobb’s Journal of Software Tools, 8(8):9–??, August 1983. CODEN DDJOEB. ISSN 1044-789X.

Willoughby:1983:RRD

[Wil83c] Steve Willoughby. Review of REVAS disassembler from REVASCO. Dr. Dobb’s Journal of Software Tools, 8(2):62–??, February 1983. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1983:UCF

[Wil83d] William E. Wilson. Unix to CP/M — floppy disk file conversion. Dr. Dobb’s Journal of Software Tools, 8(10):20–??, October 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Wilton:1983:BLC

[Wil83e] Richard Wilton. BIGLOAD — load .COM file larger than 64K in 16BST. Dr. Dobb’s Journal of Software Tools, 8(11):110– ??, November 1983. CODEN DDJOEB. ISSN 1044-789X.

Wilton:1983:OMI

[Wil83f] Richard Wilton. Optical mouse interface in 16BST. Dr. Dobb’s Journal of Software Tools, 8(9): 120–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1984:MBM

[Wil84] Dale Wilson. More on binary magic numbers. Dr. Dobb’s Journal of Software Tools, 9(3):78–??, March 1984. CODEN DDJOEB. ISSN 1044-789X.

Wilcox:1985:DRC

[Wil85a] Alan D. Wilcox. Designing a realtime clock for the S-100 bus. Dr. Dobb’s Journal of Software Tools, 10(7):56–??, July 1985. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1985:BPI

[Wil85b] William E. Wilson. Batch procedures improved in CP/M ver. 2.2 in CPME. Dr. Dobb’s Journal of Software Tools, 10(4):112– ??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1985:CME

[Wil85c] William E. Wilson. CP/M exchange. Dr. Dobb’s Journal of Software Tools, 10(12):114–??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1985:PPC

[Wil85d] William E. Wilson. PIP patch CP/M V2.2 in CPME. Dr. Dobb’s Journal of Software Tools, 10(4): 112–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Wilton:1985:UFP

[Wil85e] Richard Wilton. Unstructured Forth programming — an introduction. Dr. Dobb’s Journal of Software Tools, 10(1):86–??, January 1985. CODEN DDJOEB. ISSN 1044-789X.

Wilcox:1986:BFS

[Wil86] Alan D. Wilcox. Bringing up the 68000 — A first step. Dr. Dobb’s Journal of Software Tools, 11(1): 60–??, January 1986. CODEN DDJOEB. ISSN 1044-789X.

Winer:1985:SNS

[Win85] David Winer. Some notes on software design in TSD. Dr. Dobb’s Journal of Software Tools, 10(4):86–??, April 1985. CODEN DDJOEB. ISSN 1044-789X.

Winter:1989:AMC

[Win89] Ron Winter. Automatic module control revisited. Dr. Dobb’s Journal of Software Tools, 14(13 (C Sourcebook for the 1990s)):42– 46, 73–85, Winter 1989. CODEN DDJOEB. ISSN 1044-789X.

Wischmeyer:1983:SFC

[Wis83] Edward Wischmeyer. Some Forth coding standards. Dr. Dobb’s Journal of Software Tools, 8(9): 94–??, September 1983. CODEN DDJOEB. ISSN 1044-789X.

Woodhull:1982:FSE

[Woo82] Albert S. Woodhull. Forth screen editor for the H-19. Dr. Dobb’s Journal of Software Tools, 7(9): 42–??, September 1982. CODEN DDJOEB. ISSN 1044-789X.

Woodhull:1985:BBM

[Woo85] Albert S. Woodhull. BANKSWAP: A banked memory debugging tool for CP/M plus. Dr. Dobb’s Journal of Software Tools, 10(12):34– ??, December 1985. CODEN DDJOEB. ISSN 1044-789X.

Zigon:1989:RLE

[Zig89] Robert Zigon. Run length encoding. Dr. Dobb’s Journal of Software Tools, 14(2):126–??, February 1989. CODEN DDJOEB. ISSN 0888-3076.

Zimniewicz:1987:EIP

[Zim87] Thomas A. Zimniewicz. Extended IBM PC COM port driver. Dr. Dobb’s Journal of Software Tools, REFERENCES 116

12(6):42–??, June 1987. CODEN DDJOEB. ISSN 0888-3076.